### v2

# A Myhill-Nerode Theorem for Generalized Automata, with Applications to Pattern Matching and Compression

Abstract
The model of generalized automata, introduced by Eilenberg in 1974, allows representing a regular language more concisely than conventional automata by allowing edges to be labeled not only with characters, but also strings. Giammaresi and Montalbano introduced a notion of determinism for generalized automata [STACS 1995]. While generalized deterministic automata retain many properties of conventional deterministic automata, the uniqueness of a minimal generalized deterministic automaton is lost.
In the first part of the paper, we show that the lack of uniqueness can be explained by introducing a set $\mathcal{W(A)}$ associated with a generalized automaton $\mathcal{A}$. The set $\mathcal{W(A)}$ is always trivially equal to the set of all prefixes of the language recognized by the automaton, if $\mathcal{A}$ is a conventional automaton, but this need not be true for generalized automata. By fixing $\mathcal{W(A)}$, we are able to derive for the first time a full Myhill-Nerode theorem for generalized automata, which contains the textbook Myhill-Nerode theorem for conventional automata as a degenerate case.
In the second part of the paper, we show that the set $\mathcal{W(A)}$ leads to applications for pattern matching and data compression. Wheeler automata [TCS 2017, SODA 2020] are a popular class of automata that can be compactly stored using $e\log\sigma(1+o(1))+O(e)$ bits ($e$ being the number of edges, $\sigma$ being the size of the alphabet) in such a way that pattern matching queries can be solved in $\tilde{O}(m)$ time ($m$ being the length of the pattern). In the paper, we show how to extend these results to generalized automata. More precisely, a Wheeler generalized automata can be stored using $\mathfrak{e}\log\sigma(1+o(1))+O(e+rn)$ bits so that pattern matching queries can be solved in $\tilde{O}(rm)$ time, where $\mathfrak{e}$ is the total length of all edge labels, $r$ is the maximum length of an edge label and $n$ is the number of states.
keywords:
Generalized Automata, Myhill-Nerode Theorem, Regular Languages, Wheeler Graphs, FM-index, Burrows-Wheeler Transform
category:
[…]
1 Introduction
The class of regular languages can be defined starting from non-deterministic finite automata (NFAs). In his monumental work [17] on automata theory (which dates back to 1974), Eilenberg proposed a natural generalization of NFAs where edges can be labeled not only with characters but with (possibly empty) finite strings, the so-called generalized non-deterministic finite automata (GNFAs). While classical automata are only a special case of generalized automata, it is immediate to realize that generalized automata can only recognize regular languages, because it is well-known that epsilon transitions do not add expressive power [31], and a string-labeled edge can be decomposed into a path of edges labeled only with characters. However, generalized automata can represent regular languages more concisely than classical automata. A standard measure of the complexity of a regular language is the the minimum number of states of some automaton recognizing the language, and generalized automata may have fewer states than conventional automata. In generalized automata, we assume that both the number of states and the number of edges are finite, but the number of edges cannot be bounded by some function of the number of states and the size of the finite alphabet (and so edge labels may be arbitrarily long). As a consequence, in principle it is not clear whether the problem of determining the minimum number states of some generalized automaton recognizing a given language is decidable. In [29], Hashiguchi showed that the problem is decidable by proving that there must exist a state-minimal generalized automaton for which the lengths of edge labels can be bounded by a function that only depends on the size of the syntactic monoid recognizing the language.
An NFA is a deterministic finite automaton (DFA) if no state has two distinct outgoing edges with the same label. This local notion of determinism extends to global determinism, that is, given a string $\alpha$, there exists at most one path labeled $\alpha$ that can be followed starting from the initial state. However, this is not true for generalized automata (see Figure 1). When considering generalized automata, we must add the additional requirement that no state has two distinct outgoing edges such that one edge label is a prefix of the other edge label. By adding this requirement, we retrieve global determinism, thus obtaining generalized deterministic finite automata (GDFAs).
[…]
Figure 1: No state has two distinct outgoing edges with the same label, but there are two distinct paths labeled $abc$ from the initial state.
For every regular language, there exists a unique deterministic automaton recognizing the language and having the minimum number of states among all deterministic automata recognizing the language, the minimal DFA of the language. More generally, a classical textbook result in automata theory is the Myhill-Nerode theorem. Let $\operatorname{Pref}(\mathcal{L})$ be the set of all strings prefixing at least one string in the language $\mathcal{L}$. We have the following result.
Theorem 1.1 (Myhill-Nerode theorem).
Let $\mathcal{L}\subseteq\Sigma^{*}$. The following are equivalent:
  1. 1.
    $\mathcal{L}$ is recognized by an NFA.
  2. 2.
    The Myhill-Nerode equivalence $\equiv_{\mathcal{L}}$ has finite index.
  3. 3.
    There exists a right-invariant equivalence relation $\sim$ on $\operatorname{Pref}(\mathcal{L})$ of finite index such that $\mathcal{L}$ is the union of some $\sim$-classes.
  4. 4.
    $\mathcal{L}$ is recognized by a DFA.
Moreover, if one of the above statements is true (and so all the above statements are true), then there exists a unique minimal DFA recognizing $\mathcal{L}$ (that is, two DFAs having the minimum number of states among all DFAs that recognize $\mathcal{L}$ must be isomorphic).
The problem of studying the notion of determinism in the setting of generalized automata was approached by Giammaresi and Montalbano [27, 26]. The notion of isomorphism can be naturally extended to GDFAs (intuitively, two GDFAs are isomorphic if they are the same GDFA up to renaming the states), and the natural question is whether one can analogously define the minimal GDFA of a regular language. This is not possible: in general, there can exist two or more non-isomorphic state-minimal GDFAs recognizing a given regular language. Consider the two distinct GDFAs in Figure 2. It is immediate to check that the two (non-isomorphic) GDFAs recognize the same language, and it can be shown that the no GDFA with less than three states can recognize the same language [27].
[…]
Figure 2: Two state-minimal GDFAs recognizing the same regular language.
The non-uniqueness of a state-minimal GDFAs seems to imply a major difference in the behavior of generalized automata compared to conventional automata, so it looks like there is no hope to derive a structural result like the Myhill-Nerode theorem in the model of the generalized automata. It is natural to wonder whether the lack of uniqueness should be interpreted as a weakness of the model of generalized automata, or rather as a consequence of some deeper property. Consider a conventional automaton $\mathcal{A}$ that recognizes a language $\mathcal{L}$. As typical in automata theory, we can assume that all states are reachable from the initial state, and all states are either final or they allow reaching a final state. Then, the set $\mathcal{W(A)}$ of all strings that can be read starting from the initial state and reaching some state is exactly equal to $\operatorname{Pref}(\mathcal{L})$. However this is no longer true in the model of generalized automata: typically, we do not have $\mathcal{W(A)}=\operatorname{Pref}(\mathcal{L})$, but only $\mathcal{W(A)}\subseteq\operatorname{Pref}(\mathcal{L})$. For example, consider Figure 2. In both automata we have $a^{3}\in\operatorname{Pref}(\mathcal{L})$, but we have $a^{3}\in\mathcal{W(A)}$ only for the GDFA on the left.
Given $\mathcal{W}\subseteq\operatorname{Pref}(\mathcal{L})$, we say that a GNFA $\mathcal{A}$ recognizing $\mathcal{L}$ is a $\mathcal{W}$-GNFA if $\mathcal{W(A)}=\mathcal{W}$. We will show that, if $\mathcal{L}$ is recognized by a $\mathcal{W}$-GDFA, then there exists a unique state-minimal $\mathcal{W}$-GDFA recognizing $\mathcal{L}$. In particular, our result will imply the uniqueness of the minimal automaton for standard DFAs, because for DFAs it must necessarily be $\mathcal{W}=\operatorname{Pref}(\mathcal{L})$.
We will actually prove much more. We will show that, once we fix $\mathcal{W}$, then nondeterminism and determinism still have the same expressive power, and it is possible to derive a characterization in terms of equivalence relations. In other words, we will prove a Myhill-Nerode theorem for generalized automata. To this end, we will introduce the notion of locally bounded set (Definition 3.2), which we can use to prove the following result.
Theorem 1.2 (Myhill-Nerode theorem for generalized automata).
Let $\mathcal{L}\subseteq\Sigma^{*}$ and let $\mathcal{W}\subseteq\Sigma^{*}$ be a locally bounded set such that $\mathcal{L}\cup\{\epsilon\}\subseteq\mathcal{W}\subseteq\operatorname{Pref}(\mathcal{L})$. The following are equivalent:
  1. 1.
    $\mathcal{L}$ is recognized by a $\mathcal{W}$-GNFA.
  2. 2.
    The Myhill-Nerode equivalence $\equiv_{\mathcal{L},\mathcal{W}}$ has finite index.
  3. 3.
    There exists a right-invariant equivalence relation $\sim$ on $\mathcal{W}$ of finite index such that $\mathcal{L}$ is the union of some $\sim$-classes.
  4. 4.
    $\mathcal{L}$ is recognized by a $\mathcal{W}$-GDFA.
Moreover, if one of the above statements is true (and so all the above statements are true), then there exists a unique minimal $\mathcal{W}$-GDFA recognizing $\mathcal{L}$ (that is, two $\mathcal{W}$-GDFAs having the minimum number of states among all $\mathcal{W}$-GDFAs that recognize $\mathcal{L}$ must be isomorphic).
In particular, we will show that there is no loss of generality in assuming that $\mathcal{W}\subseteq\Sigma^{*}$ is a locally bounded set such that $\mathcal{L}\cup\{\epsilon\}\subseteq\mathcal{W}\subseteq\operatorname{Pref}(\mathcal{L})$, because these are necessary conditions for the existence of a $\mathcal{W}$-GNFA. We conclude that our Myhill-Nerode theorem for GNFAs provides the first structural result for the model of generalized automata.
In the second part of the paper, we show that the set $\mathcal{W(A)}$ sheds new light on the String Matching in Labeled Graphs (SMLG) problem. The SMLG problem has a fascinating history that dates back to more than 30 years ago. Loosely speaking, the SMLG problem can be defined as follows: given a directed graph whose nodes are labeled with nonempty strings and given a pattern string, decide whether the pattern can be read by following a path on the graph and concatenating the node labels. The SMLG problem is a natural generalization of the classical pattern matching problem on texts (which requires determining whether a pattern occurs in a text) because texts can be seen as graphs consisting of a single path. The pattern matching problem on text can be efficiently solved in $O(n+m)$ time ($n$ being the length of the text, $m$ being the length of the pattern) by using the Knuth-Morris-Pratt algorithm [34]. The SMLG problem is more challenging, and the complexity can be affected by the specific variant of the pattern matching problem under consideration or the class of graphs to which the problem is restricted. For example, in the (approximate) variant where one allows errors in the graph the problem becomes NP-hard [5], so generally errors are only allowed in the pattern. The SMLG problem was studied extensively during the nineties [35, 1, 39, 5, 41, 37]; Amir et al. showed how to solve the (exact) SMLG problem on arbitrary graphs in $O(\mathfrak{e}+me)$ time [5], where $e$ is the number of edges in the graph, $m$ is the length of the pattern, and $\mathfrak{e}$ is the total length of all labels in the graph. Recently, the SMLG problem has been back in the spotlight. Equi et al. [20, 18, 19] showed that, on arbitrary graphs, for every $\epsilon>0$ the SMLG cannot be solved in $O(me^{1-\epsilon})$ or $O(m^{1-\epsilon}e)$ time, unless the Orthogonal Vectors hypothesis fails. In applications (especially in bioinformatics) we often need faster algorithms, so the SMLG problem has been restricted to class of graphs on which it can be solved more efficiently. For example, Elastic Founder graphs can be used to represent multiple sequence alignments (MSA), a central model of biological evolution, and on Elastic Founder graphs the SMLG problem can be solved in linear time under a number of assumptions which only have a limited impact on the generality of the model [22, 21, 42].
The pattern matching problem on texts has been revolutionized by the invention of the Burrows-Wheeler Transform [10] and the FM-index [23, 24], which allow solving pattern matching queries efficiently on compressed text, thus establishing a new paradigm in bioinformatics (where the huge increase of genomic data requires the development of space-efficient algorithms) [43]. Recently, these ideas were extended to NFAs. In particular, Wheeler NFAs are a popular class of automata on which the SMLG problem can be solved in linear time, while only storing a compact representation of the Wheeler NFA [25, 3]. A special case of Wheeler NFAs are de Bruijn graphs [9], which are used to perform Eulerian sequence assembly [32, 40, 7]. Wheeler NFAs are also of relevant theoretical interest: for example, the powerset construction applied to a Wheeler NFA leads to a linear blow-up in the number of states of the equivalent DFA, and the equivalent DFA is Wheeler [4]; on arbitrary NFAs, the blow-up can be exponential.
The missing step is to determine whether it is possible to generalize the Burrows-Wheeler Transform and the FM-index to GNFAs, so that the resulting data structures can also be applied to Elastic Founder graphs and other classes of graphs where labels can be arbitrary strings. Indeed, in data compression it is common to consider edge-labeled graphs where one compresses unary paths in the graph to save space and the path is replaced by a single edge labeled with the concatenation of all labels. For example, some common data structures that are stored using this mechanism are Patricia trees, suffix trees and pangenomes [38, 6, 36]. In the following, we only consider GNFAs without $\epsilon$-transitions (that is, GNFAs where no edge is labeled with the empty string $\epsilon$) because in general a GNFA may contain arbitrarily long paths consisting of adjacent $\epsilon$-transitions, so the SMLG problem becomes more difficult and our mechanism, based on the FM-index, fails. We say that a GNFA is an $r$-GNFA if all edge labels have length at most $r$ (so a GNFA is a conventional NFA if an only if it is a $1$-GNFA). Let $m$, $e$ and $\mathfrak{e}$ as above, let $n$ be the number of states, and let $\sigma=|\Sigma|$. In Section 4, we will extend the notion of Wheelerness to GNFAs. The key ingredient will be the same set $\mathcal{W(A)}$ that we use in our Myhill-Nerode theorem: we will consider a partial order $\preceq_{\mathcal{A}}$ which sorts the set of all states with respect to the co-lexicographica order of the strings in $\mathcal{W(A)}$ (See Definition 4.1). We will then prove the following result.
Theorem 1.3 (FM-index of generalized automata).
Let $\mathcal{A}$ be a Wheeler $r$-GNFA. Then, we can encode $\mathcal{A}$ by using $\mathfrak{e}\log\sigma(1+o(1))+O(e+rn)$ bits so that later on, given a pattern $\alpha\in\Sigma^{*}$ of length $m$, we can solve the SMLG problem on $\mathcal{A}$ in $O(rm(\log r+\log\log\sigma))$ time. Within the same time bound we can also decide whether $\alpha\in\mathcal{L(A)}$.
If $r=1$ (that is, if $\mathcal{A}$ is a conventional Wheeler NFA), we conclude that we can encode $\mathcal{A}$ by using $e\log\sigma(1+o(1))+O(e)$ bits (we assume that every state is reachable from the initial state, thus $e\geq n-1)$ so that we can solve the SMLG problem in $O(m\log\log\sigma)$ time, that is, we retrieve the time and space bound which were already known for Wheeler automata [25, 14]. If $r=O(1)$, we can still solve pattern matching queries in linear time (for constant alphabets), thus breaking the lower bound by Equi et al., while only storing a compact representation of $\mathcal{A}$.
Due to space constraints, some proofs and some auxiliary results can be found in the appendix.
2 Preliminary Definitions
Let $\Sigma$ be a finite alphabet. We denote by $\Sigma^{*}$ the set of all finite strings on $\Sigma$. We denote by $\epsilon$ the empty string, and $\Sigma^{+}=\Sigma^{*}\setminus\{\epsilon\}$ is the set of all nonempty finite strings on $\Sigma$. For $i\geq 0$, let $\Sigma^{i}\subseteq\Sigma^{*}$ be the set of all strings of length $i$ (we will often interpret $\Sigma^{i}$ as a new alphabet). If $\mathcal{L}\subseteq\Sigma^{*}$, let $\operatorname{Pref}(\mathcal{L})$ be the set of all prefixes of some string in $\mathcal{L}$. Note that if $\mathcal{L}\not=\emptyset$, then $\epsilon\in\operatorname{Pref}(\mathcal{L)}$. We say that $\mathcal{L}\subseteq\Sigma^{*}$ is prefix-free if no string in $\mathcal{L}$ is a strict prefix of another string in $\mathcal{L}$. Note that if $\mathcal{L}$ is prefix-free and $\epsilon\in\mathcal{L}$, then =$\mathcal{L}=\{\epsilon\}$. If $\mathcal{L}\subseteq\Sigma^{*}$, the prefix-free kernel of $\mathcal{L}$ is the set $\mathcal{K}(\mathcal{L})$ of all strings in $\mathcal{L}$ whose strict prefixes are all not in $\mathcal{L}$. Note that $\mathcal{K}(\mathcal{L})$ is always prefix-free, and $\mathcal{L}$ is prefix-free if and only if $\mathcal{L}=\mathcal{K}(\mathcal{L})$.
Let us recall the definition of generalized deterministic finite automaton (GDFA) [26, 27].
Definition 2.1.
A generalized non-deterministic finite automaton (GNFA) is a 4-tuple $\mathcal{A}=(Q,E,s,F)$, where $Q$ is a finite set of states, $E\subseteq Q\times Q\times\Sigma^{*}$ is a finite set of string-labeled edges, $s\in Q$ is the initial state and $F\subseteq Q$ is a set of final states. Moreover, we assume that each $u\in Q$ is reachable from the initial state and is co-reachable, that is, it is either final or allows reaching a final state.
A generalized deterministic finite automaton (GDFA) is a GNFA such that for every $u\in Q$ no edge leaving $u$ is labeled with $\epsilon$, distinct edges leaving $u$ have distinct labels, and the set of all strings labeling some edge leaving $u$ is prefix-free.
The assumption that every state is reachable and co-reachable is standard in automata theory because all states that do not satisfy this requirement can be removed without changing the recognized language. A conventional NFA (DFA, respectively) is a GNFA (GDFA, respectively) where all edges are labeled with characters from $\Sigma$. Note that we explicitly require a GNFA to have finitely many edges (in conventional NFAs, the finiteness of the number of states automatically implies the finiteness of the number of edges, the alphabet being finite). If we allowed a GNFA to have infinitely many edges, then any nonempty (possibly non-regular) language would be recognized by a GNFA with two states, where the first state is initial, the second state is final, all edges go from the first state to the second state, and a string labels an edge if and only if it is in the language. By requiring a GNFA to have finitely many edges, the class of recognized languages is exactly the class of regular languages, because it is easy to transform a GNFA into a NFA with $\epsilon$-transitions that recognizes the same language by proceeding as follows: for every edge $(u^{\prime},u,\rho)\in E$, with $\rho=r_{1},\dots,r_{|\rho|}\in\Sigma^{+}$, where $r_{1},\dots,r_{|\rho|}\in\Sigma$ and $|\rho|\geq 2$, we delete the edge $(u^{\prime},u,\rho)$, we add $|\rho|-1$ new states $z_{1},\dots,z_{|\rho|-1}$ and then we add the edges $(u^{\prime},z_{1},r_{1})$, $(z_{1},z_{2},r_{2})$, $\dots$, $(z_{|\rho|-1},u,r_{|\rho|})$ (none of the new states is made initial or final).
Let us introduce some notation that will be helpful in the paper.
Definition 2.2.
Let $\mathcal{A}=(Q,E,s,F)$ be a GNFA.
  1. 1.
    For every $\alpha\in\Sigma^{*}$, let $I_{\alpha}$ be the set of all states that can be reached from the initial state by following edges whose labels, when concatenated, yield $\alpha$. In other words, for some $t\geq 0$ there exist edges $(s,u_{1},\alpha_{1}),(u_{1},u_{2},\alpha_{2}),(u_{2},u_{3},\alpha_{3}),\dots,(u_{t-1},u_{t},\alpha_{t})$ such that $\alpha=\alpha_{1}\alpha_{2}\alpha_{3}\dots\alpha_{t}$.
  2. 2.
    Let $\mathcal{L(A)}$ be the language recognized by $\mathcal{A}$, that is, $\mathcal{L(A)}=\{\alpha\in\Sigma^{*}\;|\;I_{\alpha}\cap F\not=\emptyset\}$.
  3. 3.
    For every $u\in Q$, let $I_{u}$ be the set of all strings that can be read from the initial state to $u$ by concatenating edge labels, that is, $I_{u}=\{\alpha\in\Sigma^{*}\;|\;u\in I_{\alpha}\}$. Note that for every $u\in Q$ we have $\emptyset\subsetneqq I_{u}\subseteq\operatorname{Pref}(\mathcal{L(A)})$ because every state is reachable and co-reachable.
When $\mathcal{A}$ is not clear from the context, we write $I_{\alpha}^{\mathcal{A}}$ and $I_{u}^{\mathcal{A}}$.
Lastly, following the introduction of the paper, we can naturally define the SMLG problem for GNFAs.
Definition 2.3.
Let $\mathcal{A}$ be a GNFA. The String Matching in Labeled Graphs (SMLG) problem for GNFAs is defined as follows: build a data structure that encodes $\mathcal{A}$ such that, given a string $\alpha$, we can efficiently compute the set of all states reached by a path suffixed by $\alpha$.
3 The Myhill-Nerode Theorem for Generalized Automata
The Myhill-Nerode theorem for conventional automata (Theorem 1.1) provides some algebraic properties that $\operatorname{Pref}(\mathcal{L})$ must satisfy for $\mathcal{L}\subseteq\Sigma^{*}$ to be a regular language. Intuitively, the reason why $\operatorname{Pref}(\mathcal{L})$ captures the regularity of a regular language (that is, the link between the algebraic characterization and the automata characterization of regular languages) is that, given an NFA $\mathcal{A}=(Q,E,s,F)$ that recognizes $\mathcal{L}$, we have $\bigcup_{u\in Q}I_{u}=\operatorname{Pref}(\mathcal{L)}$ because if $\alpha\in\operatorname{Pref}(\mathcal{L})$, one can read $\alpha$ on $\mathcal{A}$ starting from the initial state. However, if more generally $\mathcal{A}=(Q,E,s,F)$ is a GNFA that recognizes $\mathcal{L}$, then we only have $\bigcup_{u\in Q}I_{u}\subseteq\operatorname{Pref}(\mathcal{L)}$, because if $\alpha\in\operatorname{Pref}(\mathcal{L})$, then one can read $\alpha$ on $\mathcal{A}$ starting from the initial state, but it may happen that we have read only a strict prefix of the label of the last edge, if the label is a string of length at least two.
Let us give the following definition.
Definition 3.1.
Let $\mathcal{A}=(Q,E,s,F)$ be a GNFA. Define:
$\mathcal{W(A)}=\bigcup_{u\in Q}I_{u}.$
We say that $\mathcal{A}$ is a $\mathcal{W(A)}$-GNFA.
Note that for every $\alpha\in\Sigma^{*}$ we have $I_{\alpha}\not=\emptyset$ if and only if $\alpha\in\mathcal{W(A)}$. Moreover, $\mathcal{L(A)}\cup\{\epsilon\}\subseteq\mathcal{W(A)}\subseteq\operatorname{Pref}(\mathcal{L(A)})$, because (i) if $\alpha\in\mathcal{L(A)}$, then $I_{\alpha}\cap F\not=\emptyset$ and in particular $\alpha\in\mathcal{W(A)}$, (ii) $\epsilon\in I_{s}$, (iii) $I_{u}\subseteq\operatorname{Pref}(\mathcal{L(A)})$ for every $u\in Q$. Let us prove an additional property of $\mathcal{W(A)}$. Pick $\alpha\in\mathcal{W(A)}$, and consider the set $T_{\alpha}=\{\rho\in\Sigma^{+}\;|\;\alpha\rho\in\mathcal{W(A)}\}$. Consider the prefix-free kernel $\mathcal{K}(T_{\alpha})$. If $\rho\in\mathcal{K}(T_{\alpha})$, then $I_{\alpha\rho}\not=\emptyset$, but for every $\rho^{\prime}\in\Sigma^{+}$ being a strict nonempty prefix of $\rho$ we have $I_{\alpha\rho^{\prime}}=\emptyset$. This implies that $|\rho|\leq r$, where $r$ is the maximum of the lengths of edge labels. We conclude that $\mathcal{K}(T_{\alpha})$ must be finite, because $\Sigma$ is finite. This leads to the following definition.
Definition 3.2.
  1. 1.
    Let $\mathcal{W}\subseteq\Sigma^{*}$. We say that $\mathcal{W}$ is locally bounded if for every $\alpha\in\mathcal{W}$ we have that $\mathcal{K}(T_{\alpha})$ is finite, where $T_{\alpha}=\{\rho\in\Sigma^{+}\;|\;\alpha\rho\in\mathcal{W}\}$.
  2. 2.
    Let $\mathcal{A}=(Q,E,s,F)$ be a GNFA, and let $\mathcal{W}\subseteq\Sigma^{*}$ be a locally bounded set such that $\mathcal{L(A)}\cup\{\epsilon\}\subseteq\mathcal{W}\subseteq\operatorname{Pref}(\mathcal{L(A)})$. We say that $\mathcal{A}$ is a $\mathcal{W}$-GNFA if $\mathcal{W(A)}=\mathcal{W}$.
Remark 3.3.
Let $\mathcal{A}=(Q,E,s,F)$ be a GDFA. Let $\alpha\in\mathcal{W(A)}$. If $\alpha=\epsilon$, then $I_{\epsilon}=\{s\}$ because no edge is labeled with $\epsilon$. If $|\alpha|>1$, then there exist a prefix $\alpha_{1}\in\Sigma^{*}$ of $\alpha$ and $u_{1}\in Q$ such that $(s,u_{1},\alpha_{1})\in E$, and since $\mathcal{A}$ is a GDFA, we have $\alpha_{1}\in\Sigma^{+}$, and both $\alpha_{1}$ and $u_{1}$ are unique. In particular, $\alpha_{1}\in\mathcal{W(A)}$. If $\alpha_{1}$ is a strict prefix of $\alpha$, we can repeat the argument starting from $u_{1}$. We conclude that for every $\alpha\in\mathcal{W(A)}$ we have $|I_{\alpha}|=1$. As a consequence, if $u,v\in Q$ are distinct, then $I_{u}\cap I_{v}=\emptyset$. In the following, if $\mathcal{A}$ is a GDFA and $\alpha\in\mathcal{W(A)}$, we will identify $I_{\alpha}$ and the state being its unique element.
Moreover, our argument shows that, if $\mathcal{A}$ is a GDFA, then for every $\alpha\in\mathcal{W(A)}$ such that $|\alpha|>0$, the longest strict prefix of $\alpha$ in $\mathcal{W(A)}$ is the unique strict prefix $\alpha^{\prime}$ of $\alpha$ in $\mathcal{W(A)}$ such that, letting $\rho\in\Sigma^{+}$ be the string for which $\alpha=\alpha^{\prime}\rho$, we have $(I_{\alpha^{\prime}},I_{\alpha},\rho)\in E$. This implies that if $\mathcal{A}$ is a GDFA and $\alpha\in\mathcal{W(A)}$, then $\mathcal{K}(T_{\alpha})=\{\rho\in\Sigma^{+}\;|\;\alpha\rho\in\mathcal{W(A)},(I_{\alpha},I_{\alpha\rho},\rho)\in E\}$.
In the classical Myhill-Nerode we consider equivalence relations defined on $\operatorname{Pref}(\mathcal{L})$. In our setting, we will need to define equivalence relations on subsets of $\operatorname{Pref}(\mathcal{L})$. This leads to the following general definition.
Definition 3.4.
Let $\mathcal{W}\subseteq\Sigma^{*}$ and let $\sim$ be an equivalence relation on $\mathcal{W}$. We say that $\sim$ is right-invariant if:
$(\forall\alpha,\beta\in\mathcal{W})(\forall\phi\in\Sigma^{*})(\alpha\sim\beta\to((\alpha\phi\in\mathcal{W}\Longleftrightarrow\beta\phi\in\mathcal{W})\land(\alpha\phi\in\mathcal{W}\Longrightarrow\alpha\phi\sim\beta\phi)).$
Remark 3.5.
Notice that the property defining a right-invariant equivalence relation is trivially true if $\phi$ is the empty string, so it can be rephrased as follows:
$(\forall\alpha,\beta\in\mathcal{W})(\forall\phi\in\Sigma^{+})(\alpha\sim\beta\to((\phi\in T_{\alpha}\Longleftrightarrow\phi\in T_{\beta})\land(\phi\in T_{\alpha}\Longrightarrow\alpha\phi\sim\beta\phi)).$
Let us prove that $\sim$ is right-invariant if and only if:
$(\forall\alpha,\beta\in\mathcal{W})(\forall\phi\in\Sigma^{+})(\alpha\sim\beta\to((\phi\in\mathcal{K}(T_{\alpha})\Longleftrightarrow\phi\in\mathcal{K}(T_{\beta}))\land(\phi\in\mathcal{K}(T_{\alpha})\Longrightarrow\alpha\phi\sim\beta\phi)).$
(normal-$\Rightarrow$) Let $\alpha,\beta\in\mathcal{W}$ such that $\alpha\sim\beta$, and let $\phi\in\mathcal{K}(T_{\alpha})$. We must prove that $\phi\in\mathcal{K}(T_{\beta})$ and $\alpha\phi\sim\beta\phi$. Since $\phi\in\mathcal{K}(T_{\alpha})\subseteq T_{\alpha}$, we immediately obtain $\phi\in T_{\beta}$ and $\alpha\phi\sim\beta\phi$, so we only have to prove that $\phi\in\mathcal{K}(T_{\beta})$. Since for every $\phi^{\prime}\in\Sigma^{+}$ we have $\phi^{\prime}\in T_{\alpha}$ if and only if $\phi^{\prime}\in T_{\beta}$, then $T_{\alpha}=T_{\beta}$, and so $\mathcal{K}(T_{\alpha})=\mathcal{K}(T_{\beta})$. As a consequence, from $\phi\in\mathcal{K}(T_{\alpha})$ we conclude $\phi\in\mathcal{K}(T_{\beta})$.
(normal-$\Leftarrow$) Let $\alpha,\beta\in\mathcal{W}$ such that $\alpha\sim\beta$, and let $\phi\in T_{\alpha}$. We must prove that $\phi\in T_{\beta}$ and $\alpha\phi\sim\beta\phi$. Let $\phi_{1},\dots,\phi_{s}$ be all prefixes of $\phi$ such that $\alpha\phi_{i}\in\mathcal{W}$ for every $1\leq i\leq s$, where $\phi_{i}$ is a strict prefix of $\phi_{i+1}$ for every $11\leq i\leq s-$ 1. Note that $s\geq 2$, $\phi_{1}=\epsilon$ and $\phi_{s}=\phi$. For every $11\leq i\leq s-$ 1, let $\psi_{i}\in\Sigma^{+}$ be such that $\phi_{i+1}=\phi_{i}\psi_{i}$. Notice that by definition we have $\psi_{i}\in\mathcal{K}(T_{\alpha\phi_{i}})$ for every $11\leq i\leq s-$ 1. Since $\alpha,\beta\in\mathcal{W}$, $\alpha\sim\beta$ and $\psi_{1}\in\mathcal{K}(T_{\alpha\phi_{1}})=\mathcal{K}(T_{\alpha})$, we obtain $\psi_{1}\in\mathcal{K}(T_{\beta})$ and $\alpha\phi_{2}=\alpha\psi_{1}\sim\beta\psi_{1}=\beta\phi_{2}$. Since $\alpha\phi_{2},\beta\phi_{2}\in\mathcal{W}$, $\alpha\phi_{2}\sim\beta\phi_{2}$ and $\psi_{2}\in\mathcal{K}(T_{\alpha\phi_{2}})$, we obtain $\psi_{2}\in\mathcal{K}(T_{\beta\phi_{2}})$ and $\alpha\phi_{3}=\alpha\phi_{2}\psi_{2}\sim\beta\phi_{2}\psi_{2}=\beta\phi_{3}$. By continuing like that, we conclude that $\phi\in T_{\beta}$ and $\alpha\phi=\alpha\phi_{s}\sim\beta\phi_{s}=\beta\phi$.
In general, an equivalence relation is not right-invariant. Let us show how to define a canonical right-invariant equivalence relation starting from any equivalence relation.
Lemma 3.6.
Let $\mathcal{W}\subseteq\Sigma^{*}$ and let $\sim$ be an equivalence relation on $\mathcal{W}$. For every $\alpha,\beta\in\mathcal{W}$, let:
$\alpha\sim_{r}\beta\Longleftrightarrow(\forall\phi\in\Sigma^{*})((\alpha\phi\in\mathcal{W}\Longleftrightarrow\beta\phi\in\mathcal{W})\land(\alpha\phi\in\mathcal{W}\Longrightarrow\alpha\phi\sim\beta\phi)).$
Then $\sim_{r}$ is an equivalence relation on $\mathcal{W}$, it is right-invariant and it is the coarsest right-invariant equivalence relation on $\mathcal{W}$ refining $\sim$. We say that $\sim_{r}$ is the right-invariant refinement of $\sim$.
The Myhill-Nerode equivalence plays a major role in the classical Myhill-Nerode theorem. Let us show how we can extend it when $\mathcal{W}$ is not necessarily equal to $\operatorname{Pref}(\mathcal{L})$.
Definition 3.7.
Let $\mathcal{L},\mathcal{W}\subseteq\Sigma^{*}$. The Myhill-Nerode equivalence on $\mathcal{L}$ and $\mathcal{W}$ is the equivalence relation $\equiv_{\mathcal{L},\mathcal{W}}$ on $\mathcal{W}$ defined as the right-invariant refinement of $\sim_{\mathcal{L},\mathcal{W}}$, where $\sim_{\mathcal{L},\mathcal{W}}$ is the equivalence relation on $\mathcal{W}$ such that for every $\alpha,\beta\in\mathcal{W}$:
$\alpha\sim_{\mathcal{L},\mathcal{W}}\beta\Longleftrightarrow(\alpha\in\mathcal{L}\Longleftrightarrow\beta\in\mathcal{L}).$
If $\mathcal{W}=\operatorname{Pref}(\mathcal{L})$, then we retrieve the classical Myhill-Nerode equivalence relation for $\mathcal{L}$. Let us describe some elementary properties of $\equiv_{\mathcal{L},\mathcal{W}}$.
Lemma 3.8.
Let $\mathcal{L},\mathcal{W}\subseteq\Sigma^{*}$. Then $\equiv_{\mathcal{L},\mathcal{W}}$ is right-invariant and $\mathcal{L}$ is the union of some $\equiv_{\mathcal{L},\mathcal{W}}$-classes.
Proof 3.9.
First, $\equiv_{\mathcal{L},\mathcal{W}}$ is right-invariant because it is a right-invariant refinement by definition. Moreover, $\mathcal{L}$ is the union of some $\sim_{\mathcal{L},\mathcal{W}}$-classes, and so also of some $\equiv_{\mathcal{L},\mathcal{W}}$-classes because $\equiv_{\mathcal{L},\mathcal{W}}$ refines $\sim_{\mathcal{L},\mathcal{W}}$.
Let $\mathcal{A}$ be a conventional NFA, and define the equivalence relation $\sim_{\mathcal{A}}$ on $\operatorname{Pref}(\mathcal{L(A)})$ as follows: for every $\alpha,\beta\in\operatorname{Pref}(\mathcal{L(A)})$, let $\alpha\sim_{\mathcal{A}}\beta$ if and only if $I_{\alpha}=I_{\beta}$. This equivalence relation is an intermediate tool in the Myhill-Nerode theorem for conventional automata, and it can be also defined for a generalized automata $\mathcal{A}$ by considering the equivalence relation $\sim_{\mathcal{A}}$ on $\mathcal{W(A)}$ such that for every $\alpha,\beta\in\mathcal{W(A)}$ we have $\alpha\sim_{\mathcal{A}}\beta$ if and only if $I_{\alpha}=I_{\beta}$. If $\mathcal{A}$ is an NFA (or an NFA with $\epsilon$-transitions), then $\sim_{\mathcal{A}}$ is right-invariant, because for every $\alpha\in\operatorname{Pref}(\mathcal{L(A)})$ and for every prefix $\alpha^{\prime}$ of $\alpha$, any path from the initial state to a node in $I_{\alpha}$ must go through a node in $I_{\alpha^{\prime}}$. However, in general this property is not true if $\mathcal{A}$ is a GNFA, so $\sim_{\mathcal{A}}$ need not be right-invariant if $\mathcal{A}$ is a GNFA (see Figure 3). Since right-invariance is crucial in the Myhill-Nerode theorem, we will consider the right-invariant refinement of $\sim_{\mathcal{A}}$.
[…]
Figure 3: A GNFA $\mathcal{A}$ such that $\sim_{\mathcal{A}}$ is not right-invariant. Indeed, $a,b,ac,bc\in\mathcal{W(A)}$ and $a\sim_{\mathcal{A}}b$, but $ac\not\sim_{\mathcal{A}}bc$.
Definition 3.10.
Let $\mathcal{A}=(Q,E,s,F)$ be a GNFA. Let $\equiv_{\mathcal{A}}$ be the right-invariant refinement of $\sim_{\mathcal{A}}$, where $\sim_{\mathcal{A}}$ is the equivalence relation on $\mathcal{W(A)}$ such that for every $\alpha,\beta\in\mathcal{W(A)}$:
$\alpha\sim_{\mathcal{A}}\beta\Longleftrightarrow I_{\alpha}=I_{\beta}.$
Remark 3.11.
Let us prove that if $\mathcal{A}$ is GDFA, then $\sim_{\mathcal{A}}$ is right-invariant. Let $\alpha,\beta\in\mathcal{W}$ be such that $I_{\alpha}=I_{\beta}$, and let $\phi\in\mathcal{K}(T_{\alpha})$. By Remark 3.5, we only have to prove that $\phi\in\mathcal{K}(T_{\beta})$ and $I_{\alpha\phi}=I_{\beta\phi}$. By Remark 3.3, we have $\alpha\phi\in\mathcal{W(A)}$ and $(I_{\alpha},I_{\alpha\phi},\phi)\in E$. Hence $(I_{\beta},I_{\alpha\phi},\phi)\in E$, we obtain $I_{\alpha\phi}=I_{\beta\phi}$ and again by Remark 3.3, we conclude $\phi\in\mathcal{K}(T_{\beta})$. Notice that, in fact, the generalized automaton $\mathcal{A}$ in Figure 3 is not a GDFA.
Since $\equiv_{\mathcal{A}}$ is the right-invariant refinement of $\sim_{\mathcal{A}}$, we conclude that, if $\mathcal{A}$ is a GDFA, then $\equiv_{\mathcal{A}}$ and $\sim_{\mathcal{A}}$ are the same equivalence relation.
Let us study the properties of $\equiv_{\mathcal{A}}$.
Lemma 3.12.
Let $\mathcal{A}=(Q,E,s,F)$ be a GNFA. Then, $\equiv_{\mathcal{A}}$ is right-invariant, it refines $\equiv_{\mathcal{L(A)},\mathcal{W(A)}}$, it has finite index and $\mathcal{L(A)}$ is the union of some $\equiv_{\mathcal{A}}$-classes.
The following lemma is crucial to derive our Myhill-Nerode theorem for generalized automata.
Lemma 3.13.
Let $\mathcal{L}\subseteq\Sigma^{*}$ and let $\mathcal{W}\subseteq\Sigma^{*}$ be a locally bounded set such that $\mathcal{L}\cup\{\epsilon\}\subseteq\mathcal{W}\subseteq\operatorname{Pref}(\mathcal{L})$. Assume that $\mathcal{L}$ is the union of some classes of a right-invariant equivalence relation $\sim$ on $\mathcal{W}$ of finite index. Then, $\mathcal{L}$ is recognized by a $\mathcal{W}$-GDFA $\mathcal{A_{\sim}}=(Q_{\sim},E_{\sim},s_{\sim},F_{\sim})$ such that:
  1. 1.
    $|Q_{\sim}|$ is equal to the index of $\sim$.
  2. 2.
    $\equiv_{\mathcal{A_{\sim}}}$ and $\sim$ are the same equivalence relation.
Moreover, if $\mathcal{B}$ is a $\mathcal{W}$-GDFA that recognizes $\mathcal{L}$, then $\mathcal{A_{\equiv_{\mathcal{B}}}}$ is isomorphic to $\mathcal{B}$.
Proof 3.14.
(Sketch) The desired $\mathcal{W}$-GDFA $\mathcal{A_{\sim}}=(Q_{\sim},E_{\sim},s_{\sim},F_{\sim})$ is defined as follows.
  • $Q_{\sim}=\{[\alpha]_{\sim}\;|\;\alpha\in\mathcal{W}\}$.
  • $s_{\sim}=[\epsilon]_{\sim}$.
  • $E_{\sim}=\{([\alpha]_{\sim},[\alpha\rho]_{\sim},\rho)\;|\;\alpha\in\mathcal{W},\rho\in\mathcal{K}(T_{\alpha})\}$.
  • $F_{\sim}=\{[\alpha]_{\sim}\;|\;\alpha\in\mathcal{L}\}$.
Remark 3.15.
In the statement of Lemma 3.13 we cannot remove the assumption that $\mathcal{W}$ is locally bounded, because we have shown that if $\mathcal{A}$ is a GNFA, then $\mathcal{W(A)}$ is locally bounded. However, if $\mathcal{W}$ is not locally bounded, then $\mathcal{A}_{\sim}$ is still a well-defined automaton with finitely many states, but it has infinitely many edges. For example, $\mathcal{W}=\{\epsilon\}\cup a^{*}b$ is not locally bounded because $T_{\epsilon}=a^{*}b$ and $\mathcal{K}(T_{\epsilon})=a^{*}b$ is an infinite set. If $\mathcal{L}=a^{*}b$, then $\mathcal{L}\cup\{\epsilon\}\subseteq\mathcal{W}\subseteq\operatorname{Pref}(\mathcal{L})$. Moreover, $\equiv_{\mathcal{L},\mathcal{W}}$ has finite index (the equivalence classes are $\{\epsilon\}$ and $a^{*}b$), and by Lemma 3.8 we know that $\equiv_{\mathcal{L},\mathcal{W}}$ is right-invariant and $\mathcal{L}$ is the union of some $\equiv_{\mathcal{L},\mathcal{W}}$-classes. We conclude that $\mathcal{A_{\equiv_{\mathcal{L},\mathcal{W}}}}$ is well-defined, but it has infinitely many edges (see Figure 4).
[…]
Figure 4: An example where $\mathcal{W}$ is not locally bounded.
We can now state our Myhill-Nerode theorem for generalized automata.
Theorem 3.16 (Myhill-Nerode theorem for generalized automata).
Let $\mathcal{L}\subseteq\Sigma^{*}$ and let $\mathcal{W}\subseteq\Sigma^{*}$ be a locally bounded set such that $\mathcal{L}\cup\{\epsilon\}\subseteq\mathcal{W}\subseteq\operatorname{Pref}(\mathcal{L})$. The following are equivalent:
  1. 1.
    $\mathcal{L}$ is recognized by a $\mathcal{W}$-GNFA.
  2. 2.
    The Myhill-Nerode equivalence $\equiv_{\mathcal{L},\mathcal{W}}$ has finite index.
  3. 3.
    There exists a right-invariant equivalence relation $\sim$ on $\mathcal{W}$ of finite index such that $\mathcal{L}$ is the union of some $\sim$-classes.
  4. 4.
    $\mathcal{L}$ is recognized by a $\mathcal{W}$-GDFA.
Moreover, if one of the above statements is true (and so all the above statements are true), then there exists a unique minimal $\mathcal{W}$-GDFA recognizing $\mathcal{L}$ (that is, two $\mathcal{W}$-GDFAs having the minimum number of states among all $\mathcal{W}$-GDFAs that recognize $\mathcal{L}$ must be isomorphic).
Proof 3.17.
$(1)\to(2)$ Let $\mathcal{A}$ be a $\mathcal{W}$-GDFA recognizing $\mathcal{L}$. By Lemma 3.12 we have that $\equiv_{\mathcal{A}}$ has finite index and it refines $\equiv_{\mathcal{L},\mathcal{W}}$, so also $\equiv_{\mathcal{L},\mathcal{W}}$ has finite index.
$(2)\to(3)$ By Lemma 3.8 the desired equivalence relation is $\equiv_{\mathcal{L},\mathcal{W}}$.
$(3)\to(4)$ It follows from Lemma 3.13.
$(4)\to(1)$ Every $\mathcal{W}$-GDFA is a $\mathcal{W}$-GNFA.
Now, let us prove that the minimum automaton is $\mathcal{A_{\equiv_{\mathcal{L},\mathcal{W}}}}$ as defined in Lemma 3.13. First, $\mathcal{A_{\equiv_{\mathcal{L},\mathcal{W}}}}$ is well-defined because $\equiv_{\mathcal{L},\mathcal{W}}$ is right-invariant and $\mathcal{L}$ is the union of some $\equiv_{\mathcal{L},\mathcal{W}}$-classes by Lemma 3.8, and $\equiv_{\mathcal{L},\mathcal{W}}$ has finite index by one of the statements that we assume to be true. Now, by Lemma 3.13 the number of states of $\mathcal{A_{\equiv_{\mathcal{L},\mathcal{W}}}}$ is equal to the index of $\equiv_{\mathcal{L},\mathcal{W}}$, or equivalently, of $\equiv_{\mathcal{A_{\equiv_{\mathcal{L},\mathcal{W}}}}}$. On the other hand, let $\mathcal{B}$ be any $\mathcal{W}$-GDFA recognizing $\mathcal{L}$ non-isomorphic to $\mathcal{A_{\equiv_{\mathcal{L},\mathcal{W}}}}$. Then $\equiv_{\mathcal{B}}$ is a refinement of $\equiv_{\mathcal{L},\mathcal{W}}$ by Lemma 3.12, and it must be a strict refinement of $\equiv_{\mathcal{L},\mathcal{W}}$, otherwise $\mathcal{A_{\equiv_{\mathcal{L},\mathcal{W}}}}$ would be equal to $\mathcal{A}_{\equiv_{\mathcal{B}}}$, which by Lemma 3.13 is isomorphic to $\mathcal{B}$, a contradiction. We conclude that the index of $\equiv_{\mathcal{L},\mathcal{W}}$ is smaller than the index of $\equiv_{\mathcal{B}}$, so again by Lemma 3.13 the number of states of $\mathcal{A_{\equiv_{\mathcal{L},\mathcal{W}}}}$ is smaller than the number of states of $\mathcal{A}_{\equiv_{\mathcal{B}}}$ and so of $\mathcal{B}$.
Myhill-Nerode theorem for conventional automata (Theorem 1.1) is a special case of Theorem 3.16, because if $\mathcal{A}$ is an NFA, then $\mathcal{W(A)}=\operatorname{Pref}(\mathcal{L(A)})$. Moreover, Theorem 3.16 is consistent with the example in Figure 2, because, calling $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ the two GDFAs in Figure 2, we have shown that $\mathcal{W}(\mathcal{A}_{1})\not=\mathcal{W}(\mathcal{A}_{2})$.
4 The FM-index of Generalized Automata
In this section, we prove Theorem 1.3. In order to present the main ideas, it will suffice to consider GDFAs. The more general case of GNFAs without $\epsilon$-transitions requires minor technical modifications and will be considered in Appendix C.
Let $V$ be a set. We say that a (binary) relation $\leq$ on $V$ is a partial order if $\leq$ is reflexive, antisymmetric and transitive. A partial order $\leq$ is a total order if for every $u,v\in V$ we have $(u\leq v)\lor(v\leq u)$. We say that $U\subseteq V$ is $\leq$-convex if for every $u,v,z\in V$, if $u\leq v\leq z$ and $u,z\in U$, then $v\in U$. For every $u,v\in V$, we write $u<v$ if $(u\leq v)\land(u\not=v)$.
Let us define Wheeler GDFAs. As customary in the literature on Wheeler automata [3, 16], we assume that there exists a fixed total order $\preceq$ on $\Sigma$ (in our examples, we always assume $a\prec b\prec c\prec\dots$), and $\preceq$ is extended co-lexicographically to $\Sigma^{*}$ (that is, for every $\alpha,\beta\in\Sigma^{*}$ we have $\alpha\prec\beta$ if the reverse string $\alpha^{R}$ is lexicographically smaller than the reverse string $\beta^{R}$). Let $\mathcal{A}=(Q,E,s,F)$ be a GDFA. Let $\preceq_{\mathcal{A}}$ be the reflexive relation on $Q$ such that, for every $u,v\in Q$ with $u\not=v$, we have $u\prec_{\mathcal{A}}v$ if and only if $(\forall\alpha\in I_{u})(\forall\beta\in I_{v})(\alpha\prec\beta)$. Since each $I_{u}$ is nonempty, it is immediate to realize that $\preceq_{\mathcal{A}}$ is a partial order, but in general it is not a total order. We can then give the following definition (see Figure 5 for an example).
Definition 4.1.
Let $\mathcal{A}=(Q,E,s,F)$ be a GDFA. We say that $\mathcal{A}$ is Wheeler if $\preceq_{\mathcal{A}}$ is a total order.
[…]
  • $\mathtt{OUT}_{1}=001011$
  • $\mathtt{OUT}_{2}=001101$
  • $\mathtt{IN}_{1}=100101$
  • $\mathtt{IN}_{2}=101001$
  • $\mathtt{LAB}_{1}=(b)(c)(b)$
  • $\mathtt{LAB}_{2}=(ab)(ac)(bc)$
  • $\mathtt{FIN}=011$
Figure 5: Left: A Wheeler GDFA $\mathcal{A}$. States are numbered following the total order $\preceq_{\mathcal{A}}$. Right: The Burrows-Wheeler Transform (BWT) of $\mathcal{A}$ (see Definition 4.10).
If $\mathcal{A}$ is a conventional DFA, it is not immediately clear that Definition 4.1 is equivalent to the local definition of Wheeler DFA commonly used in the literature [3, 11, 15]. According to the local definition, a DFA $\mathcal{A}=(Q,E,s,F)$ is Wheeler if there exists a total order $\leq$ on $Q$ such that (i) $s$ comes first in the total order, (ii) for every $(u^{\prime},u,a),(v^{\prime},v,b)\in E$, if $u<v$, then $a\preceq b$ and (iii) for every $(u^{\prime},u,a),(v^{\prime},v,a)\in E$, if $u<v$, then $u^{\prime}<v^{\prime}$. Alanko et al. [3, Corollary 3.1] proved that, if such a total order $\leq$ exists, then it is unique and it is equal to $\preceq_{\mathcal{A}}$, so we only have to prove that if $\preceq_{\mathcal{A}}$ is a total order, then it satisfies properties (i), (ii), (iii). This follows from the following lemma.
Lemma 4.2.
Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler GDFA. Then:
  1. 1.
    $s$ comes first in the total order $\preceq_{\mathcal{A}}$.
  2. 2.
    For every $(u^{\prime},u,\rho),(v^{\prime},v,\rho^{\prime})\in E$, if $u\prec_{\mathcal{A}}v$ and $\rho^{\prime}$ is not a strict suffix of $\rho$, then $\rho\preceq\rho^{\prime}$.
  3. 3.
    For every $(u^{\prime},u,\rho),(v^{\prime},v,\rho)\in E$, if $u\prec_{\mathcal{A}}v$, then $u^{\prime}\prec_{\mathcal{A}}v^{\prime}$.
In case 2 of Lemma 4.2 we cannot remove the assumption that $\rho^{\prime}$ is not a strict prefix of $\rho$; as a consequence, we cannot use Lemma 4.2 to provide an equivalent, local definition of Wheeler GDFA (see Figure 6). The local definition of Wheeler DFA easily implies that the problem of deciding whether a given DFA is Wheeler can be solved in polynomial time [3], but since we do not have a local definition of Wheeler GDFA, it is not clear whether the corresponding problem on GDFAs is also solvable in polynomial time (and we saw in the introduction that computational problems on generalized automata are usually hard). However, we can prove that the problem is still tractable by reducing it to computing the partial order $\preceq_{\mathcal{A^{*}}}$ on a conventional DFA $\mathcal{A^{*}}$ equivalent to a given GDFA $\mathcal{A}$.
[…]
Figure 6: Left: A Wheeler GDFA $\mathcal{A}=(Q,E,s,F)$. States are numbered following the total order $\preceq_{\mathcal{A}}$. Note that $(u_{1},u_{2},ba),(u_{4},u_{3},a)\in E$, $u_{2}\prec_{\mathcal{A}}u_{3}$, $a$ is a strict suffix of $ba$ and $a\prec ba$. Right: A GDFA $\mathcal{A}=(Q,E,s,F)$ such that states are numbered following a total order $\leq$ such that (i) $s$ comes first in the total order $\leq$, (ii) for every $(u^{\prime},u,\rho),(v^{\prime},v,\rho^{\prime})\in E$, if $u<v$ and $\rho^{\prime}$ is not a strict suffix of $\rho$, then $\rho\preceq\rho^{\prime}$ and (iii) for every $(u^{\prime},u,\rho),(v^{\prime},v,\rho)\in E$, if $u<v$, then $u^{\prime}<v^{\prime}$. Note that $\mathcal{A}$ is not Wheeler because $u_{2}$ and $u_{3}$ are not $\preceq_{\mathcal{A}}$-comparable, since $b\prec c\prec ac$, $c\in I_{u_{3}}$ and $b,ac\in I_{u_{2}}$.
Lemma 4.3.
Let $\mathcal{A}=(Q,E,s,F)$ be a GDFA, and let $\mathfrak{e}$ be the total length of all edge labels. In $O(\mathfrak{e}\log\mathfrak{e})$ time we can decide whether $\mathcal{A}$ is Wheeler and, if so, we can compute $\preceq_{\mathcal{A}}$.
Proof 4.4.
Let us build a DFA $\mathcal{A}^{*}$ starting from $A$ (see Figure 7 for an example). In general, if $C\subseteq\Sigma^{+}$, we can build a trie such that the set of all strings that can be read following a path starting from the root is equal to the set of all strings prefixing at least one element in $C$. If $C$ is prefix-free, then a string is in $C$ if and only if it can be read from the root to a leaf; we say that such a leaf spells the considered string. For every $u\in Q$, let $C_{u}$ be the prefix-free set of all strings labeling some edge leaving $u$. We build an automaton $\mathcal{A}^{*}=(Q^{*},E^{*},s^{*},F^{*})$ by picking $\mathcal{A}$, removing all edges, and building a trie for every nonempty $C_{u}$ is such a way that (a) the root of the trie is $u$, (b) the leaf that spells $\rho$ is the unique $v\in Q$ such that $(u,v,\rho)\in E$ and (c) every internal state of the trie is a new state. Notice that $Q\subseteq Q^{*}$; we define $s^{*}=s$ and $F^{*}=F$. By construction, $\mathcal{A}^{*}$ is a DFA, and for every $u\in Q$ we have $I_{u}^{\mathcal{A}}=I_{u}^{\mathcal{A^{*}}}$ (so $\mathcal{L(A^{*})}=\mathcal{L(A)}$). As a consequence, $\mathcal{A}$ is Wheeler if and only if the restriction of the partial order $\preceq_{\mathcal{A}^{*}}$ to $Q$ is a total order. Since $\mathcal{A}^{*}$ is a conventional DFA, we can compute the partial order $\preceq_{\mathcal{A}^{*}}$ in polynomial time [16, 33, 13, 8]. For example, the algorithm in [8] runs in $O(|E^{*}|\log|Q^{*}|)$ time, and the claimed time bound follows because $|Q^{*}|-1\leq|E^{*}|=\mathfrak{e}$.
[…]
Figure 7: Left: The GDFA $\mathcal{A}$ in Figure 5. Right: The DFA $A^{*}$ built starting from $\mathcal{A}$ in the proof of Lemma 4.3.
The remaining of the section is devoted to proving that the SMLG problem can be solved efficiently on Wheeler GDFAs. If $\mathcal{A}=(Q,E,s,F)$ is a Wheeler GDFA, we write $Q=\{Q[1],Q[2],\dots,Q[n]\}$, where $Q[1]\prec_{\mathcal{A}}Q[2]\prec_{\mathcal{A}}\dots\prec_{\mathcal{A}}Q[n]$; if $1\leq i\leq j\leq n$, let $Q[i,j]=\{Q[i],Q[i+1],\dots,Q[j-1],Q[j]\}$. If $\alpha,\beta\in\Sigma^{*}$, we write $\alpha\dashv\beta$ if and only if $\alpha$ is a suffix of $\beta$.
Definition 4.5.
Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler GDFA, and let $\alpha\in\Sigma^{*}$. Define:
  • $G^{\prec}(\alpha)=\{u\in Q\;|\;(\forall\beta\in I_{u})(\beta\prec\alpha)\}$;
  • $G_{\dashv}(\alpha)=\{u\in Q\;|\;(\exists\beta\in I_{u})(\alpha\dashv\beta)\}$;
  • $G^{\prec}_{\dashv}(\alpha)=G^{\prec}(\alpha)\cup G_{\dashv}(\alpha)=\{u\in Q\ \;|\;(\forall\beta\in I_{u})(\beta\prec\alpha)\vee(\exists\beta\in I_{u})(\alpha\dashv\beta)\}$.
Intuitively, the set $G_{\dashv}(\alpha)$ is the set of states that the SMLG problem must return on input $\alpha$, and $G^{\prec}(\alpha)$ is the set of all states reached only by strings smaller than $\alpha$. The following lemma shows that, as in the case of conventional Wheeler automata, both $G^{\prec}(\alpha)$ and $G_{\dashv}(\alpha)$ are intervals with respect to the total order $\preceq_{\mathcal{A}}$, and there is no state between $G^{\prec}(\alpha)$ and $G_{\dashv}(\alpha)$.
Lemma 4.6.
Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler GDFA, and let $\alpha\in\Sigma^{*}$. Then:
  1. 1.
    $G^{\prec}(\alpha)\cap G_{\dashv}(\alpha)=\emptyset$.
  2. 2.
    $G_{\dashv}(\alpha)$ is $\preceq_{\mathcal{A}}$-convex.
  3. 3.
    If $u,v\in Q$ are such that $u\prec_{\mathcal{A}}v$ and $v\in G^{\prec}(\alpha)$, then $u\in G^{\prec}(\alpha)$. In other words, $G^{\prec}(\alpha)=Q[1,|G^{\prec}(\alpha)|]$.
  4. 4.
    If $u,v\in Q$ are such that $u\prec_{\mathcal{A}}v$ and $v\in G^{\prec}_{\dashv}(\alpha)$, then $u\in G^{\prec}_{\dashv}(\alpha)$. In other words, $G^{\prec}_{\dashv}(\alpha)=Q[1,|G^{\prec}_{\dashv}(\alpha)|]$.
In order to compute $G_{\dashv}(\alpha)$, it will suffice to compute $G^{\prec}(\alpha)$ and $G^{\prec}_{\dashv}(\alpha)$. Let us show how to compute $G^{\prec}(\alpha)$ and $G^{\prec}_{\dashv}(\alpha)$; to this end, we will constantly use property 3 in Lemma 4.2, which is also crucial for conventional Wheeler automata (it is a generalization of the LF mapping in the FM-index [24]). First, let $G^{*}(\alpha)$ be the set of all states reached by an edge labeled with a string suffixed by $\alpha$. Formally:
$G^{*}(\alpha)=\{v\in Q\;|\;(\exists v^{\prime}\in Q)(\exists\rho\in\Sigma^{*})((v^{\prime},v,\rho)\in E\land(\alpha\dashv\rho))\}.$
Clearly, $G^{*}(\alpha)\subseteq G_{\dashv}(\alpha)$. Now, let us give the following definition.
Definition 4.7.
Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler GDFA. Let $U\subseteq Q$ and $\rho\in\Sigma^{+}$. We denote by $\mathtt{out}(U,\rho)$ the number of edges labeled with $\rho$ that leave states in $U$, and we denote by $\mathtt{in}(U,\rho)$ the number of edges labeled with $\rho$ that enter states in $U$.
If $\alpha\in\Sigma^{*}$ and $0\leq k\leq|\alpha|$, let $p(\alpha,k)$ and $s(\alpha,k)$ be the prefix and the suffix of $\alpha$ of length $k$, respectively. If $\mathcal{A}=(Q,E,s,F)$ is a GDFA and $u\in Q$, let $\lambda(u)$ be the set of all strings in $\Sigma^{+}$ labeling an edge reaching $u$; we denote by $\min_{\lambda(u)}$ and $\max_{\lambda(u)}$ the (co-lexicographically) smallest and largest strings in $\lambda(u)$, respectively.
The next lemma formalizes the following intuition: to compute $G^{\prec}(\alpha)$, we have to consider all states whose incoming edges have a label smaller than $\alpha$; moreover, if the label is $s(\alpha,k)$ (for some $k$), then the start state must be in $G^{\prec}(p(\alpha,|\alpha|-k))$.
Lemma 4.8.
Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler GDFA, and let $\alpha\in\Sigma^{*}$. For $0<k<|\alpha|$, let $f_{k}=\mathtt{out}(Q[1,|G^{\prec}(p(\alpha,|\alpha|-k))|],s(\alpha,k))$. Then, $|G^{\prec}(\alpha)|$ is the largest integer $0\leq j\leq|Q|$ such that (i) $\mathtt{in}(Q[1,j],s(\alpha,k))\leq f_{k}$, for every $0<k<|\alpha|$, and (ii) if $j\geq 1$, then $\max_{\lambda(Q[j])}\prec\alpha$.
The following crucial lemma shows that, in order to compute $|G^{\prec}_{\dashv}(\alpha)|$, we only have to consider $|G^{\prec}(\alpha)|$, the biggest (w.r.t $\preceq_{\mathcal{A}}$) state in $G^{*}(\alpha)$ and the states in $G^{\prec}_{\dashv}(\alpha)\setminus G^{*}(\alpha)$.
Lemma 4.9.
Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler GDFA, and let $\alpha\in\Sigma^{*}$. For $0<k<|\alpha|$, let $f_{k}=\mathtt{out}(Q[1,|G^{\prec}(p(\alpha,|\alpha|-k))|],s(\alpha,k))$ and $g_{k}=\mathtt{out}(Q[1,|G^{\prec}_{\dashv}(p(\alpha,|\alpha|-k))|],s(\alpha,k))$. Then, $g_{k}\geq f_{k}$ for every $0<k<|\alpha|$. Moreover, $|G^{\prec}_{\dashv}(\alpha)|$ is equal to the maximum among:
  1. 1.
    $|G^{\prec}(\alpha)|$;
  2. 2.
    the largest integer $0\leq i\leq|Q|$ such that, if $i\geq 1$, then $Q[i]\in G^{*}(\alpha)$;
  3. 3.
    the smallest integer $0\leq j\leq|Q|$ such that, for every $0<k<|\alpha|$ such that $g_{k}>f_{k}$, we have $\mathtt{in}(Q[1,j],s(\alpha,k))\geq g_{k}$.
We are now ready to generalize the Burrows-Wheeler Transform and the FM-index to Wheeler GDFAs. Let $\mathcal{A}=(Q,E,s,F)$ be a GDFA. We say that $\mathcal{A}$ is a $r$-GDFA if all edge labels have length at most $r$. Fix $1\leq i\leq r$. Let $E(i)=\{(u^{\prime},u,\rho)\in E\;|\;\rho\in\Sigma^{i}\}$ be the number of edges labeled with a string of length $i$, and let $\Sigma(i)=\{\rho\in\Sigma^{i}\;|\;(\exists u^{\prime},u\in Q)((u^{\prime},u,\rho)\in E(i)\}$ be the set of all strings of length $i$ labeling some edge. Let $e_{i}=|E(i)|$ and $\sigma_{i}=|\Sigma(i)|$; we have $\sigma_{i}\leq\min\{\sigma^{i},e_{i}\}$. The $i$-outdegree ($i$-indegree, respectively) of a state is equal to the number of edges in $E(i)$ leaving (reaching, respectively) the state. The sum of the $i$-outdegrees of all the states and the sum of the $i$-indegrees of all the states are both equal to $e_{i}$. Lastly, $\sum_{i=1}^{r}e_{i}=e$ and the total length of all edge labels is $\mathfrak{e}=\sum_{i=1}^{r}e_{i}i$.
Let us define the Burrows-Wheeler Transform of a Wheeler GDFA (see Figure 5 for an example).
Definition 4.10 (Burrows-Wheeler Transform of a Wheeler GDFA).
Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler GDFA. The Burrows-Wheeler Transform $BWT(\mathcal{A})$ of $\mathcal{A}$ consists of the following strings.
  1. 1.
    For every $1\leq i\leq r$, the bit string $\mathtt{OUT}_{i}\in\{0,1\}^{e_{i}+n}$ that stores the $i$-outdegrees in unary. More precisely, (i) $\mathtt{OUT}_{i}$ contains exactly $n$ characters equal to $1$, (ii) $\mathtt{OUT}_{i}$ contains exactly $e_{i}$ characters equal to 00, and (iii) the number of zeros between the $(\ell-1)$-th character equal to one (or the beginning of the sequence if $\ell=1$) and the $\ell$-th character equal to $1$ yields the $i$-outdegree of $Q[\ell]$.
  2. 2.
    For every $1\leq i\leq r$, the bit string $\mathtt{IN}_{i}\in\{0,1\}^{e_{i}+n}$ that stores the $i$-indegrees in unary. More precisely, (i) $\mathtt{IN}_{i}$ contains exactly $n$ characters equal to $1$, (ii) $\mathtt{IN}_{i}$ contains exactly $e_{i}$ characters equal to 00, and (iii) the number of zeros between the $(\ell-1)$-th character equal to one (or the beginning of the sequence if $\ell=1$) and the $\ell$-th character equal to $1$ yields the $i$-indegree of $Q[\ell]$.
  3. 3.
    For every $1\leq i\leq r$, the string $\mathtt{LAB}_{i}\in(\Sigma^{i})^{e_{i}}$ that stores the edge labels of length $i$ (with their multiplicities). More precisely, we sort of all edges in $E(i)$ by the index of the start states (w.r.t to $\preceq_{\mathcal{A}}$). Edges with the same start state are sorted by label. Lastly, we obtain $\mathtt{LAB}_{i}$ by concatenating the labels of all the edges following this edge order.
  4. 4.
    The bit string $\mathtt{FIN}\in\{0,1\}^{n}$ that marks the final states, that is, the $i$-th bit of $\mathtt{FIN}$ is equal to 1 if and only if $Q[i]\in F$.
We can now prove that the BWT of a Wheeler GDFA $\mathcal{A}$ is a valid encoding of $\mathcal{A}$, just like the BWT of a string is a valid encoding of the string.
Theorem 4.11.
Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler GDFA. If we only know $BWT(\mathcal{A})$, then we can retrieve $\mathcal{A}$ (up to isomorphism). In other words, $BWT(\mathcal{A})$ is an encoding of $\mathcal{A}$.
Proof 4.12.
Consider states $Q[1],\dots,Q[n]$. By Lemma 4.2 we have $s=Q[1]$, and by using $\mathtt{FIN}$ we can retrieve the set of all final states $F$. We only have to show that we can retrieve $E$. In other words, for every $1\leq i^{\prime},i\leq n$ and for every $\rho\in\Sigma^{+}$ we must determine whether $(Q[i^{\prime}],Q[i],\rho)\in E$. It will suffice to retrieve the set $E(i)$ for every $1\leq i\leq r$, because $E$ is the (disjoint) union of the $E(i)$’s. From $\mathtt{LAB}_{i}$ we can retrieve the labels all edges in $E(i)$, with their multiplicities. From $\mathtt{IN}_{i}$ we can retrieve the $i$-indegree of each $Q[\ell]$. By Lemma 4.2, for every $\rho\in\Sigma^{i}$ labeling some edge reaching some node $Q[\ell]$ and for every $\rho^{\prime}\in\Sigma^{i}$ labeling some edge reaching $Q[\ell+1]$ it must be $\rho\preceq\rho^{\prime}$. Since we know the labels of all edges in $E(i)$ with multiplicities and we know the $i$-indegrees, then we can retrieve the labels of all edges entering each $Q[\ell]$, with multiplicities. From $\mathtt{OUT}_{i}$ we can retrieve the $i$-outdegrees of each $Q[\ell]$, and the order used in the definition of $\mathtt{LAB}_{i}$ implies that we can retrieve the labels of all edges leaving each $Q[\ell]$. Since we know the labels of all edges entering each $Q[\ell]$ and we know the labels of all edges leaving each $Q[\ell]$, then for every $\rho\in\Sigma^{i}$ we know the set of all states reached by an edge labeled $\rho$, with multiplicities, and the set of all states having an outgoing edge labeled $\rho$. By Lemma 4.2, for every $1\leq j<k\leq n$, if $Q[j]$ is reached by an edge labeled $\rho$ leaving the state $Q[j^{\prime}]$ and $Q[k]$ is reached by an edge labeled $\rho$ leaving the state $Q[k^{\prime}]$, then it must be $j^{\prime}<k^{\prime}$. As a consequence, we can retrieve the set $E(i)_{\rho}$ of all edges labeled $\rho$ for every $\rho\in\Sigma^{i}$, and so we can retrieve $E(i)$, because $E(i)$ is the (disjoint) union of the $E(i)_{\rho}$’s.
We are ready to prove the main theorem of this section (the full proof is in the appendix). Recall that $n$ is the number of vertices, $e$ is the number of edges, and $\mathfrak{e}$ is the total length of all edge labels.
Theorem 4.13 (FM-index of Wheeler GDFAs).
Let $\mathcal{A}$ be a Wheeler $r$-GDFA. Then, we can encode $\mathcal{A}$ by using $\mathfrak{e}\log\sigma(1+o(1))+O(e+rn)$ bits so that later on, given a pattern $\alpha\in\Sigma^{*}$ of length $m$, we can solve the SMLG problem on $\mathcal{A}$ in $O(rm(\log r+\log\log\sigma))$ time. Within the same time bound we can also decide whether $\alpha\in\mathcal{L(A)}$.
Proof 4.14.
(Sketch) By Lemma 4.6, we only need to compute $|G^{\prec}(\alpha)|$ and $|G^{\prec}_{\dashv}(\alpha)|$. In $O(m)$ steps, we recursively compute $|G^{\prec}(p(\alpha,k))|$ and $|G_{\dashv}(p(\alpha,k))|$ for every $0\leq k\leq|\alpha|$. By Lemma 4.8 and Lemma 4.9, we can reduce this problem to the problem of solving a number of elementary queries, such as computing $f_{k}=\mathtt{out}(Q[1,|G^{\prec}(p(\alpha,|\alpha|-k))|],s(\alpha,k))$ for every $0<k<|\alpha|$. But we only need to compute $f_{k}$ for $10<k<r+$ 1, because otherwise $|s(\alpha,k)|>r$ and all edge labels have length at most $r$. In general, we need to solve every elementary query at most $O(r)$ time. By augmenting the Burrows-Wheeler-Transform of $\mathcal{A}$ with compact data structures for rank/select queries and succinct dictionaries, we can solve every elementary query in $O(\log r+\log\log\sigma)$ time.
5 Conclusions and Future Work
In this paper, we considered the model of generalized automata, and we introduced the set $\mathcal{W(A)}$. We showed that $\mathcal{W(A)}$ plays the same role of $\operatorname{Pref}(\mathcal{L(A)})$ in conventional NFAs: the set $\mathcal{W(A)}$ can be used to derive a Myhill-Nerode theorem, and it represents the starting point for extending the FM-index to generalized automata.
Further lines of research include extending the Burrows-Wheeler Transform and the FM-index to arbitrary GNFAs. Indeed, the Burrows-Wheeler Transform and the FM-index were recently generalized from Wheeler NFAs to arbitrary NFAs by means of the so-called co-lex orders [14, 16] and co-lex relations [12]. However, we remark that the efficient time bounds for the SMLG problem that we derived in this paper cannot hold for arbitrary GNFAs due to the (conditional) lower bounds by Equi et al. that we recalled in the introduction.
Giammaresi and Montalbano described an effective procedure for computing a minimal GDFA equivalent to a given GDFA [27, 26], but we do not know if there exists an efficient algorithm for minimizing a GDFA. The Myhill-Nerode theorem for generalized automata implies that for every minimal GDFA $\mathcal{A}$ there exists a set $\mathcal{W}\subseteq\Sigma^{*}$ such that $\mathcal{A}$ is isomorphic to the minimal $\mathcal{W}$-GDFA recognizing $\mathcal{L(A)}$. At the same time, given a $\mathcal{W}$-GDFA recognizing $\mathcal{L}$, we may build the minimal $\mathcal{W}$-GDFA recognizing $\mathcal{L}$ by extending Hopcroft’s algorithm [30] to GDFAs. If we could prove that, for every admissible $\mathcal{W}\subseteq\Sigma^{*}$, the number of states of the minimal $\mathcal{W}$-GDFA recognizing $\mathcal{L}$ is comparable to the number of states of a minimal GDFA recognizing $\mathcal{L}$, then we would obtain a fast algorithm that significantly reduces the number of states of a GDFA without changing the recognized language.
This paper leaves many questions of theoretical interest open. The class of Wheeler languages is the class of all regular languages that are recognized by some Wheeler NFA [4]. Wheeler languages enjoy several properties: for example, they admit a characterization in terms of convex equivalence relations [4]. In addition, every Wheeler language is also recognized by some DFA, and, in particular, there exists a unique state-minimal DFA recognizing a given Wheeler language [2]. The main limitation of Wheeler languages is that they represent only a small subclass of regular languages: for example, a unary language (that is, a language over an alphabet of size one) is Wheeler if and only if it is either finite or co-finite [4]. The intuitive reason why most regular languages are not Wheeler is that the definition of a Wheeler NFA $\mathcal{A}$ implies strong requirements on the set $\operatorname{Pref}(\mathcal{L(A)})$ (which lead to the characterization in terms of convex equivalence relations). However, when we switch to GNFAs, it is $\mathcal{W(A)}$ that plays the role of $\operatorname{Pref}(\mathcal{L(A)})$, and it may hold $\mathcal{W(A)}\subsetneqq\operatorname{Pref}(\mathcal{L(A)})$, thus now the same requirement only apply to a smaller subset. The natural question is whether Wheeler GNFAs extend the class of Wheeler languages. The answer is affirmative: there exists a regular $\mathcal{L}$ language such that $\mathcal{L}$ is not Wheeler (that is, no Wheeler NFA recognizes $\mathcal{L}$), but $\mathcal{L}$ is recognized by a Wheeler GDFA. Define =$\mathcal{L}=\{a^{2n}\;|\;n\geq 0\}$. Then, $\mathcal{L}$ is not Wheeler [4], but $\mathcal{L}$ is recognized by the GDFA consisting of a single state, both initial and final, with a self-loop labeled $aa$. As a consequence, the class of all languages recognized by a Wheeler GNFA is strictly larger than the class of Wheeler languages. We can call the languages in this new class generalized Wheeler languages: the next step is to understand which properties of Wheeler languages are still true and how it is possible to characterize this new class.
Appendix A Proofs from Section 3
Statement of Lemma 3.6. Let $\mathcal{W}\subseteq\Sigma^{*}$ and let $\sim$ be an equivalence relation on $\mathcal{W}$. For every $\alpha,\beta\in\mathcal{W}$, let:
$\alpha\sim_{r}\beta\Longleftrightarrow(\forall\phi\in\Sigma^{*})((\alpha\phi\in\mathcal{W}\Longleftrightarrow\beta\phi\in\mathcal{W})\land(\alpha\phi\in\mathcal{W}\Longrightarrow\alpha\phi\sim\beta\phi)).$
Then $\sim_{r}$ is an equivalence relation on $\mathcal{W}$, it is right-invariant and it is the coarsest right-invariant equivalence relation on $\mathcal{W}$ refining $\sim$. We say that $\sim_{r}$ is the right-invariant refinement of $\sim$.
Proof A.1.
It is immediate to check that $\sim_{r}$ is an equivalence relation. Let us prove that $\sim_{r}$ is right-invariant. Assume that $\alpha,\beta\in\mathcal{W}$ and $\phi\in\Sigma^{*}$ are such that $\alpha\sim_{r}\beta$ and $\alpha\phi\in\mathcal{W}$. We must prove that $\beta\phi\in\mathcal{W}$ and $\alpha\phi\sim_{r}\beta\phi$. Now, $\beta\phi\in\mathcal{W}$ follows immediately from $\alpha\sim_{r}\beta$ and $\alpha\phi\in\mathcal{W}$. Moreover, $\alpha\phi\sim\beta\phi$ also follows from $\alpha\sim_{r}\beta$ and $\alpha\phi\in\mathcal{W}$. Next, fix $\psi\in\Sigma^{*}$ such that $\alpha\phi\psi\in\mathcal{W}$. We must prove that $\beta\phi\psi\in\mathcal{W}$ and $\alpha\phi\psi\sim\beta\phi\psi$. This follows from $\alpha\sim_{r}\beta$ and $\alpha\phi\psi\in\mathcal{W}$. Now, $\sim_{r}$ refines $\sim$, because for every $\alpha,\beta\in\mathcal{W}$, if $\alpha\sim_{r}\beta$, by letting $\phi$ be the empty string we conclude $\alpha\sim\beta$. Lastly, we want to prove that $\sim_{r}$ is the coarsest right-invariant equivalence relation of $\mathcal{W}$ refining $\sim$. Let $\sim_{*}$ be a right-invariant equivalence relation on $\mathcal{W}$ refining $\sim$. Assume that for some $\alpha,\beta\in\mathcal{W}$ we have $\alpha\sim_{*}\beta$. We must prove that $\alpha\sim_{r}\beta$. Let $\phi\in\Sigma^{*}$ be such that $\alpha\phi\in\mathcal{W}$. We must prove that $\beta\phi\in\mathcal{W}$ and $\alpha\phi\sim\beta\phi$. Since $\sim_{*}$ is right-invariant, from $\alpha\sim_{*}\beta$ and $\alpha\phi\in\mathcal{W}$ we obtain $\beta\phi\in\mathcal{W}$ and $\alpha\phi\sim_{*}\beta\phi$, which implies $\alpha\phi\sim\beta\phi$ because $\sim_{*}$ refines $\sim$.
Statement of Lemma 3.12. Let $\mathcal{A}=(Q,E,s,F)$ be a GNFA. Then, $\equiv_{\mathcal{A}}$ is right-invariant, it refines $\equiv_{\mathcal{L(A)},\mathcal{W(A)}}$, it has finite index and $\mathcal{L(A)}$ is the union of some $\equiv_{\mathcal{A}}$-classes.
Proof A.2.
First, $\equiv_{\mathcal{A}}$ is right-invariant because it is a right-invariant refinement by definition.
Let us prove that $\mathcal{L(A)}$ is the union of some $\equiv_{\mathcal{A}}$-classes. It will suffice to show that $\mathcal{L(A)}$ is the union of some $\sim_{\mathcal{A}}$-classes, because $\equiv_{\mathcal{A}}$ is a refinement of $\sim_{\mathcal{A}}$. Let $\alpha,\beta\in\mathcal{W(A)}$ such that $\alpha\sim_{\mathcal{A}}\beta$, that is, $I_{\alpha}=I_{\beta}$. We must prove that $\alpha\in\mathcal{L(A)}$ if and only if $\beta\in\mathcal{L(A)}$. We have $\alpha\in\mathcal{L(A)}$ if and only if $I_{\alpha}\cap F\not=\emptyset$, if and only if $I_{\beta}\cap F\not=\emptyset$, if and only if $\beta\in\mathcal{L(A)}$.
Let us prove that $\equiv_{\mathcal{A}}$ refines $\equiv_{\mathcal{L(A)},\mathcal{W(A)}}$. Since $\mathcal{L(A)}$ is the union of some $\equiv_{\mathcal{A}}$-classes, then $\equiv_{\mathcal{A}}$ refines $\sim_{\mathcal{L(A)},\mathcal{W(A)}}$. Then, $\equiv_{\mathcal{A}}$ also refines $\equiv_{\mathcal{L(A)},\mathcal{W(A)}}$, because $\equiv_{\mathcal{A}}$ is right-invariant and $\equiv_{\mathcal{L(A)},\mathcal{W(A)}}$ is the coarsest right-invariant equivalence relation refining $\sim_{\mathcal{L(A)},\mathcal{W(A)}}$.
Finally, let us prove that $\sim_{\mathcal{A}}$ has finite index. Let $\bar{\mathcal{A}}$ be the NFA with $\epsilon$-transitions equivalent to $\mathcal{A}$ that we built in Section 2. Let $\alpha,\beta\in\mathcal{W(A)}$ be such that $\alpha\sim_{\bar{\mathcal{A}}}\beta$. Let us prove that it must be $\alpha\equiv_{\mathcal{A}}\beta$. By the definition of right-invariant refinement, we must prove that for every $\phi\in\Sigma^{*}$ we have $\alpha\phi\in\mathcal{W(A)}$ if and only if $\beta\phi\in\mathcal{W(A)}$, and, if $\alpha\phi\in\mathcal{W(A)}$, then $\alpha\phi\sim_{\mathcal{A}}\beta\phi$. Since $\bar{\mathcal{A}}$ is an NFA with $\epsilon$-transitions, we have that $\sim_{\bar{\mathcal{A}}}$ is right-invariant (see the discussion before Definition 3.10), so since $\alpha\phi\in\mathcal{W(A)}\subseteq\operatorname{Pref}(\mathcal{L(A)})$, we obtain $\alpha\phi\sim_{\bar{\mathcal{A}}}\beta\phi$, which implies $I^{\bar{\mathcal{A}}}_{\alpha\phi}=I^{\bar{\mathcal{A}}}_{\beta\phi}$, which (by the construction of $\bar{\mathcal{A}}$) implies $I_{\alpha\phi}^{\mathcal{A}}=I_{\alpha\phi}^{\bar{\mathcal{A}}}\cap Q=I_{\beta\phi}^{\bar{\mathcal{A}}}\cap Q=I_{\beta\phi}^{\mathcal{A}}$. As a consequence, we have $\alpha\phi\in\mathcal{W(A)}$ if and only if $I_{\alpha\phi}^{\mathcal{A}}\not=\emptyset$, if and only if $I_{\beta\phi}^{\mathcal{A}}\not=\emptyset$, if and only if $\beta\phi\in\mathcal{W(A)}$, and, if $\alpha\phi\in\mathcal{W(A)}$, then $\alpha\phi\sim_{\mathcal{A}}\beta\phi$. Since for every $\alpha,\beta\in\mathcal{W(A)}$ we have that $\alpha\sim_{\bar{\mathcal{A}}}\beta$ implies $\alpha\equiv_{\mathcal{A}}\beta$, in order to prove that $\equiv_{\mathcal{A}}$ has finite index it will suffice to prove that $\sim_{\bar{\mathcal{A}}}$ has finite index. This follows from the observation that every $\sim_{\bar{\mathcal{A}}}$-class can be associated with a distinct nonempty subset of the set $\bar{Q}$ of all states in $\bar{\mathcal{A}}$ (by means of the well-defined mapping $[\alpha]_{\sim_{\bar{\mathcal{A}}}}\mapsto I^{\bar{\mathcal{A}}}_{\alpha})$, so the index is bounded by $12^{|\bar{Q}|}-$ 1.
Statement of Lemma 3.13. Let $\mathcal{L}\subseteq\Sigma^{*}$ and let $\mathcal{W}\subseteq\Sigma^{*}$ be a locally bounded set such that $\mathcal{L}\cup\{\epsilon\}\subseteq\mathcal{W}\subseteq\operatorname{Pref}(\mathcal{L})$. Assume that $\mathcal{L}$ is the union of some classes of a right-invariant equivalence relation $\sim$ on $\mathcal{W}$ of finite index. Then, $\mathcal{L}$ is recognized by a $\mathcal{W}$-GDFA $\mathcal{A_{\sim}}=(Q_{\sim},E_{\sim},s_{\sim},F_{\sim})$ such that:
  1. 1.
    $|Q_{\sim}|$ is equal to the index of $\sim$.
  2. 2.
    $\equiv_{\mathcal{A_{\sim}}}$ and $\sim$ are the same equivalence relation.
Moreover, if $\mathcal{B}$ is a $\mathcal{W}$-GDFA that recognizes $\mathcal{L}$, then $\mathcal{A_{\equiv_{\mathcal{B}}}}$ is isomorphic to $\mathcal{B}$.
Proof A.3.
Define $\mathcal{A_{\sim}}=(Q_{\sim},E_{\sim},s_{\sim},F_{\sim})$ as follows.
  • $Q_{\sim}=\{[\alpha]_{\sim}\;|\;\alpha\in\mathcal{W}\}$.
  • $s_{\sim}=[\epsilon]_{\sim}$.
  • $E_{\sim}=\{([\alpha]_{\sim},[\alpha\rho]_{\sim},\rho)\;|\;\alpha\in\mathcal{W},\rho\in\mathcal{K}(T_{\alpha})\}$.
  • $F_{\sim}=\{[\alpha]_{\sim}\;|\;\alpha\in\mathcal{L}\}$.
Let us prove that $\mathcal{A}_{\sim}$ is a well-defined GDFA. First, the number of states is finite because $\sim$ has finite index. Next, $\epsilon\in\mathcal{W}$, so $s_{\sim}$ is well-defined. The set $F_{\sim}$ is well-defined because (i) if $\alpha\in\mathcal{L}$, then $\alpha\in\mathcal{W}$, so $[\alpha]_{\sim}$ is well-defined, and (ii) $\mathcal{L}$ if the union of some $\sim$-classes, so if for some $\alpha,\alpha^{\prime}\in\mathcal{W}$ we have $\alpha\sim\alpha^{\prime}$, then $\alpha\in\mathcal{L}$ if and only if $\alpha^{\prime}\in\mathcal{L}$. Let us prove that $E_{\sim}$ is well-defined. First, notice that $\rho\in T_{\alpha}$, so $\alpha\rho\in\mathcal{W}$ and $[\alpha\rho]_{\sim}$ is well-defined. Moreover, if $\alpha^{\prime}\in\mathcal{W}$ is such that $\alpha\sim\alpha^{\prime}$, then $\alpha^{\prime}\rho\in\mathcal{W}$ and $\alpha\rho\sim\alpha^{\prime}\rho$, because $\sim$ is right-invariant. In particular, we have $T_{\alpha}=T_{\alpha}^{\prime}$ and so $\mathcal{K}(T_{\alpha})=\mathcal{K}(T_{\alpha}^{\prime})$, which shows that $E_{\sim}$ is well-defined. Moreover, $E_{\sim}$ is a finite set because $\mathcal{W}$ is locally bounded, so $\mathcal{K}(T_{\alpha})$ is finite for every $\alpha\in\mathcal{W}$. Let us prove by induction on $|\alpha|$ that for every $\alpha\in\mathcal{W}$ state $[\alpha]_{\sim}$ is reachable from the initial state. If $|\alpha|=0$, then $\alpha=\epsilon$ and so $[\alpha]_{\sim}$ is the initial state. Now assume $|\alpha|>0$. Let $\alpha^{\prime}$ be the longest strict prefix of $\alpha$ such that $\alpha^{\prime}\in\mathcal{W}$, which must exist because $\epsilon$ is a strict prefix of $\alpha$, being $|\alpha|>0$, and $\epsilon\in\mathcal{W}$. Let $\rho\in\Sigma^{+}$ be such that $\alpha=\alpha^{\prime}\rho$. Let us prove that $([\alpha^{\prime}]_{\sim},[\alpha]_{\sim},\rho)\in E_{\sim}$. Since $\alpha=\alpha^{\prime}\rho$, we only have to show that $\rho\in\mathcal{K}(T_{\alpha^{\prime}})$. We have $\rho\in T_{\alpha^{\prime}}$ because $\alpha=\alpha^{\prime}\rho\in\mathcal{W}$, and we have $\rho\in\mathcal{K}(T_{\alpha^{\prime}})$ because if $\rho^{\prime}\in\Sigma^{+}$ is a strict prefix of $\rho$, then $\rho^{\prime}\not\in T_{\alpha}$, otherwise we would have $\alpha^{\prime}\rho\in\mathcal{W}$, and $\alpha^{\prime}\rho^{\prime}$ would be a strict prefix of $\alpha$ longer than $\alpha^{\prime}$, which contradicts the maximality of $\alpha^{\prime}$. By the inductive hypothesis, $[\alpha^{\prime}]_{\sim}$ is reachable from the initial state, so $[\alpha]_{\sim}$ is also reachable from the initial state because $([\alpha^{\prime}]_{\sim},[\alpha]_{\sim},\rho)\in E_{\sim}$. Let us prove every state is co-reachable. Fix $\alpha\in\mathcal{W}$. We will prove that $[\alpha]_{\sim}$ is either final, or it allows reaching a final state. In particular, $\alpha\in\operatorname{Pref}(\mathcal{L})$, so there exists $\rho\in\Sigma^{*}$ such that $\alpha\rho\in\mathcal{L}$, and so $\alpha\rho\in\mathcal{W}$. Let $\rho_{1},\dots,\rho_{s}$ be all distinct prefixes of $\rho$ such that $\alpha\rho_{i}\in\mathcal{W}$ for every $1\leq i\leq s$, where $\rho_{i}$ is a strict prefix of $\rho_{i+1}$ for every $11\leq i\leq s-$ 1. Note that $s\geq 1$, $\rho_{1}=\epsilon$ and $\rho_{s}=\rho$. The same argument used to prove reachability shows that $([\alpha\rho_{i}]_{\sim},[\alpha\rho_{i+1}],\tau_{i})\in E_{\sim}$, where $\tau_{i}\in\Sigma^{+}$ is such that $\rho_{i+1}=\rho_{i}\tau_{i}$ for every $11\leq i\leq s-$ 1. Since $\alpha\rho_{s}=\alpha\rho\in\mathcal{L}$ and so $[\alpha\rho_{s}]_{\sim}$ is a final state, then $[\alpha\rho_{i}]_{\sim}$ is either final or it allows reaching a final state for every $1\leq i\leq s$, and the conclusion follows because $\alpha=\alpha\rho_{1}$. Lastly, for every $\alpha\in\mathcal{W}$, every $\rho\in\Sigma^{+}$ labels at most one edge leaving $[\alpha]_{\sim}$ because we have shown that the definition of $E_{\sim}$ does not depend on the choice of representatives in the equivalence classes, and the set of all strings labeling an edge leaving $[\alpha]_{\sim}$ is $\mathcal{K}(T_{\alpha})$, which is prefix-free, being a prefix-free kernel.
Next, we want to prove that for every $\alpha\in\Sigma^{*}$ and for every $\beta\in\mathcal{W}$ we have:
$(\alpha\in\mathcal{W(A_{\sim})})\land(I_{\alpha}=[\beta]_{\sim})\Longleftrightarrow(\alpha\in\mathcal{W})\land(\alpha\sim\beta).$ (1)
We proceed by induction on $|\alpha|$. If $|\alpha|=0$, then $\alpha=\epsilon$. Notice that $\epsilon\in\mathcal{W(A_{\sim})}$ because $\epsilon\in I_{s_{\sim}}$, and $\epsilon\in\mathcal{W}$. Moreover, for every $\beta\in\mathcal{W}$ we have $I_{\epsilon}=[\beta]_{\sim}\Longleftrightarrow[\beta]_{\sim}=s_{\sim}\Longleftrightarrow\epsilon\sim\beta$.
Now, assume $|\alpha|>0$. Note that the inductive hypothesis implies that for every $\alpha^{\prime}\in\Sigma^{*}$ such that $|\alpha^{\prime}|<|\alpha|$ we have $\alpha^{\prime}\in\mathcal{W(A_{\sim})}\Longleftrightarrow\alpha^{\prime}\in\mathcal{W}$, and, if $\alpha^{\prime}\in\mathcal{W(A_{\sim})}$, then $I_{\alpha^{\prime}}=[\alpha^{\prime}]_{\sim}$. Indeed, (normal-$\Rightarrow$) follows by choosing any $\beta\in\mathcal{W}$ such that $I_{\alpha^{\prime}}=[\beta]_{\sim}$ (which exists because $\alpha^{\prime}\in\mathcal{W(A_{\sim})}$) in Equation 1; (normal-$\Leftarrow$) and $I_{\alpha^{\prime}}=[\alpha^{\prime}]_{\sim}$ follow by choosing $\beta=\alpha^{\prime}$ in Equation 1.
Let $\alpha^{\prime}$ be the longest strict prefix of $\alpha$ such that $\alpha^{\prime}\in\mathcal{W(A_{\sim})}$, which must exist because $\epsilon$ is a strict prefix of $\alpha$, being $|\alpha|>0$, and $\epsilon\in\mathcal{W(A_{\sim})}$. Since we have shown that a string shorter than $\alpha$ is in $\mathcal{W(A_{\sim})}$ if and only if it is in $\mathcal{W}$, then $\alpha^{\prime}$ is also the longest strict prefix of $\alpha$ such that $\alpha^{\prime}\in\mathcal{W}$. Moreover, it must be $I_{\alpha^{\prime}}=[\alpha^{\prime}]_{\sim}$. Write $\alpha=\alpha^{\prime}\rho$, with $\rho\in\Sigma^{+}$.
  1. 1.
    Assume that $(\alpha\in\mathcal{W(A_{\sim})})\land(I_{\alpha}=[\beta]_{\sim})$. We must prove that $(\alpha\in\mathcal{W})\land(\alpha\sim\beta)$. Since $\alpha^{\prime},\alpha\in\mathcal{W(A_{\sim})}$, $\alpha^{\prime}$ is the longest strict prefix of $\alpha$ in $\mathcal{W(A_{\sim})}$, $\alpha=\alpha^{\prime}\rho$, and $I_{\alpha^{\prime}}=[\alpha^{\prime}]_{\sim}$, then by Remark 3.3 we obtain that $([\alpha^{\prime}]_{\sim},[\beta]_{\sim},\rho)\in E_{\sim}$. The definition of $E_{\sim}$ implies $\alpha=\alpha^{\prime}\rho\in\mathcal{W}$ and $\alpha\sim\beta$.
  2. 2.
    Assume that $(\alpha\in\mathcal{W})\land(\alpha\sim\beta)$. We must prove that $(\alpha\in\mathcal{W(A_{\sim})})\land(I_{\alpha}=[\beta]_{\sim})$. Let us prove that $([\alpha^{\prime}]_{\sim},[\alpha]_{\sim},\rho)\in E_{\sim}$. Since $\alpha^{\prime},\alpha\in\mathcal{W}$, we only have to show that $\rho\in\mathcal{K}(T_{\alpha^{\prime}})$. Since $\alpha=\alpha^{\prime}\rho$, we have $\rho\in T_{\alpha^{\prime}}$. If $\rho^{\prime}\in\Sigma^{+}$ is a strict prefix of $\rho$, then it cannot hold $\alpha^{\prime}\rho^{\prime}\in\mathcal{W}$ by the maximality of $\alpha^{\prime}$, so $\rho\in\mathcal{K}(T_{\alpha^{\prime}})$. From $I_{\alpha^{\prime}}=[\alpha^{\prime}]_{\sim}$, $([\alpha^{\prime}]_{\sim},[\alpha]_{\sim},\rho)\in E_{\sim}$ and $\alpha=\alpha^{\prime}\rho$ we obtain $\alpha\in\mathcal{W(A_{\sim})}$ and $I_{\alpha}=[\alpha]_{\sim}$. Since $\alpha\sim\beta$, we conclude $I_{\alpha}=[\beta]_{\sim}$.
This concludes the proof of Equation 1. In particular, by the same argument used at the beginning of the inductive step, we obtain $\mathcal{W(A_{\sim})}=\mathcal{W}$, which proves that $\mathcal{A_{\sim}}$ is a $\mathcal{W}$-GDFA, and for every $\alpha\in\mathcal{W(A_{\sim})}=\mathcal{W}$:
$I_{\alpha}=[\alpha]_{\sim}.$ (2)
From the definition of $F_{\sim}$ and Equation 1 we obtain:
$\mathcal{L(A_{\sim})}=\{\alpha\in\mathcal{W(A_{\sim})}\;|\;I_{\alpha}=[\beta]_{\sim}\text{ for some $\beta\in\mathcal{L}$}\}=\{\alpha\in\mathcal{W}\;|\;\alpha\sim\beta\text{ for some $\beta\in\mathcal{L}$}\}=\mathcal{L}$
where in the last equality we have ()($\subseteq$)( ⊆ ) because $\mathcal{L}$ is the union of some $\sim$-classes, and ()superset-of-or-equals($\supseteq$)( ⊇ ) because $\mathcal{L}\subseteq\mathcal{W}$. This proves that $\mathcal{A_{\sim}}$ recognizes $\mathcal{L}$. Moreover:
  1. 1.
    The number of states of $\mathcal{A}_{\sim}$ is clearly equal to the index of $\sim$.
  2. 2.
    By Equation 2, for every $\alpha,\beta\in\mathcal{W}=\mathcal{W(A_{\sim})}$ we have $\alpha\equiv_{A_{\sim}}\beta\Longleftrightarrow I_{\alpha}=I_{\beta}\Longleftrightarrow[\alpha]_{\sim}=[\beta]_{\sim}\Longleftrightarrow\alpha\sim\beta$, so $\equiv_{A_{\sim}}$ and $\sim$ are the same equivalence relation.
Finally, suppose that $\mathcal{B}=(Q_{\mathcal{B}},E_{\mathcal{B}},s_{\mathcal{B}},F_{\mathcal{B}})$ is a $\mathcal{W}$-GDFA that recognizes $\mathcal{L}$. Notice that by Lemma 3.12 we have that $\equiv_{\mathcal{B}}$ is right-invariant, it has finite index and $\mathcal{L}$ is the union of some $\equiv_{\mathcal{B}}$-classes, so $\mathcal{A_{\equiv_{\mathcal{B}}}}$ is well-defined, and $\mathcal{W(\mathcal{A_{\equiv_{\mathcal{B}}}})}=\mathcal{W}=\mathcal{W(B)}$. Let $\phi:Q_{\equiv_{\mathcal{B}}}\to Q_{\mathcal{B}}$ be the function sending the state $[\alpha]_{\equiv_{\mathcal{B}}}$ of $Q_{\equiv_{\mathcal{B}}}$ into the state $I_{\alpha}^{\mathcal{B}}$ of $Q_{\mathcal{B}}$, that is, $\phi([\alpha]_{\equiv_{\mathcal{B}}})=I_{\alpha}^{\mathcal{B}}$ for every $\alpha\in\mathcal{W}$. Notice that $\phi$ is well-defined and injective because by the definition of $\equiv_{\mathcal{B}}$ we have $[\alpha]_{\equiv_{\mathcal{B}}}=[\beta]_{\equiv_{\mathcal{B}}}\Longleftrightarrow\alpha\equiv_{\mathcal{B}}\beta\Longleftrightarrow I_{\alpha}^{\mathcal{B}}=I_{\beta}^{\mathcal{B}}$ for every $\alpha,\beta\in\mathcal{W}$. Let us prove that $\phi$ determines an isomorphism between $\mathcal{A_{\equiv_{\mathcal{B}}}}$ and $\mathcal{B}$. First, $\phi$ is surjective because every state of $\mathcal{B}$ is reachable from the initial state. Next, $\phi(s_{\equiv_{\mathcal{B}}})=\phi([\epsilon]_{\equiv_{\mathcal{B}}})=I_{\epsilon}^{\mathcal{B}}=s_{\mathcal{B}}$. Moreover for every $\alpha\in\mathcal{W}$ we have $[\alpha]_{\equiv_{\mathcal{B}}}\in F_{\equiv_{\mathcal{B}}}\Longleftrightarrow\alpha\in\mathcal{L}\Longleftrightarrow I_{\alpha}^{\mathcal{B}}\in F_{\mathcal{B}}$. Finally, by Remark 3.3, for every $\alpha,\beta\in\mathcal{W}$ and for every $\rho\in\Sigma^{+}$:
$\begin{split}([\alpha]_{\equiv_{\mathcal{B}}},[\beta]_{\equiv_{\mathcal{B}}},\rho)\in E_{\equiv_{\mathcal{B}}}&\Longleftrightarrow\rho\in\mathcal{K}(T_{\alpha})\land\alpha\rho\equiv_{\mathcal{B}}\beta\Longleftrightarrow\\ &\Longleftrightarrow\alpha\rho\in\mathcal{W}\land(I_{\alpha}^{\mathcal{B}},I_{\alpha\rho}^{\mathcal{B}},\rho)\in E_{\mathcal{B}}\land I_{\alpha\rho}^{\mathcal{B}}=I_{\beta}^{\mathcal{B}}\Longleftrightarrow\\ &\Longleftrightarrow(I_{\alpha}^{\mathcal{B}},I_{\beta}^{\mathcal{B}},\rho)\in E_{\mathcal{B}}\Longleftrightarrow(\phi([\alpha]_{\equiv_{\mathcal{B}}}),\phi([\beta]_{\equiv_{\mathcal{B}}}),\rho)\in E_{\mathcal{B}}\end{split}$
and we conclude that $\mathcal{A_{\equiv_{\mathcal{B}}}}$ and $\mathcal{B}$ are isomorphic.
Appendix B Proofs from Section 4
Statement of Lemma 4.2. Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler GDFA. Then:
  1. 1.
    $s$ comes first in the total order $\preceq_{\mathcal{A}}$.
  2. 2.
    For every $(u^{\prime},u,\rho),(v^{\prime},v,\rho^{\prime})\in E$, if $u\prec_{\mathcal{A}}v$ and $\rho^{\prime}$ is not a strict suffix of $\rho$, then $\rho\preceq\rho^{\prime}$.
  3. 3.
    For every $(u^{\prime},u,\rho),(v^{\prime},v,\rho)\in E$, if $u\prec_{\mathcal{A}}v$, then $u^{\prime}\prec_{\mathcal{A}}v^{\prime}$.
Proof B.1.
  1. 1.
    Let $u\in Q\setminus\{s\}$. We must prove that $s\prec_{\mathcal{A}}u$. Since $\preceq_{\mathcal{A}}$ is a total order, we only have to prove that if $s$ and $u$ are $\preceq_{\mathcal{A}}$-comparable, then it must be $s\prec_{\mathcal{A}}u$. This follows from the fact that for every $\alpha\in I_{u}$ we have $\epsilon\prec\alpha$, and $\epsilon\in I_{s}$.
  2. 2.
    If $\rho=\rho^{\prime}$ or $\rho$ is a strict suffix of $\rho^{\prime}$, then the conclusion is immediate, so we can assume that $\rho\not=\rho^{\prime}$, $\rho$ is not a strict suffix of $\rho^{\prime}$ and $\rho^{\prime}$ is not a strict suffix of $\rho$. Let $\alpha^{\prime}\in I_{u^{\prime}}$ and $\beta^{\prime}\in I_{v^{\prime}}$. Then, $\alpha^{\prime}\rho\in I_{u}$ and $\beta^{\prime}\rho^{\prime}\in I_{v}$. Since $u\prec_{\mathcal{A}}v$, then $\alpha^{\prime}\rho\prec\beta^{\prime}\rho^{\prime}$, Since $\rho\not=\rho^{\prime}$, $\rho$ is not a strict suffix of $\rho^{\prime}$ and $\rho^{\prime}$ is not a strict suffix of $\rho$, then we conclude $\rho\prec\rho^{\prime}$.
  3. 3.
    Let $\alpha^{\prime}\in I_{u^{\prime}}$ and $\beta^{\prime}\in I_{v^{\prime}}$. We must prove that $\alpha^{\prime}\prec\beta^{\prime}$. From $(u^{\prime},u,\rho),(v^{\prime},v,\rho)\in E$ we obtain $\alpha^{\prime}\rho\in I_{u}$ and $\beta^{\prime}\rho\in I_{v}$, so from $u\prec_{\mathcal{A}}v$ we obtain $\alpha^{\prime}\rho\prec\beta^{\prime}\rho$ and thus $\alpha^{\prime}\prec\beta^{\prime}$.
Statement of Lemma 4.6. Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler GDFA, and let $\alpha\in\Sigma^{*}$. Then:
  1. 1.
    $G^{\prec}(\alpha)\cap G_{\dashv}(\alpha)=\emptyset$.
  2. 2.
    $G_{\dashv}(\alpha)$ is $\preceq_{\mathcal{A}}$-convex.
  3. 3.
    If $u,v\in Q$ are such that $u\prec_{\mathcal{A}}v$ and $v\in G^{\prec}(\alpha)$, then $u\in G^{\prec}(\alpha)$. In other words, $G^{\prec}(\alpha)=Q[1,|G^{\prec}(\alpha)|]$.
  4. 4.
    If $u,v\in Q$ are such that $u\prec_{\mathcal{A}}v$ and $v\in G^{\prec}_{\dashv}(\alpha)$, then $u\in G^{\prec}_{\dashv}(\alpha)$. In other words, $G^{\prec}_{\dashv}(\alpha)=Q[1,|G^{\prec}_{\dashv}(\alpha)|]$.
Proof B.2.
  1. 1.
    If $u\in G_{\dashv}(\alpha)$, then there exists $\beta\in I_{u}$ such that $\alpha\dashv\beta$. In particular, $\alpha\preceq\beta$, so $u\not\in G^{\prec}(\alpha)$.
  2. 2.
    Assume that $u,v,z\in Q$ are such that $u\prec_{\mathcal{A}}v\prec_{\mathcal{A}}z$ and $u,z\in G_{\dashv}(\alpha)$. We must prove that $v\in G_{\dashv}(\alpha)$. Since $u,z\in G_{\dashv}(\alpha)$, then there exist $\beta\in I_{u}$ and $\delta\in I_{z}$ such that $\alpha\dashv\beta$ and $\alpha\dashv\delta$. Fix any $\gamma\in I_{v}$; we only have to prove that $\alpha\dashv\gamma$. From $u\prec_{\mathcal{A}}v\prec_{\mathcal{A}}z$ we obtain $\beta\prec\gamma\prec\delta$. As a consequence, from $\alpha\dashv\beta$ and $\alpha\dashv\delta$ we conclude $\alpha\dashv\gamma$.
  3. 3.
    Let $\beta\in I_{u}$. We must prove that $\beta\prec\alpha$. Fix any $\gamma\in I_{v}$. Since $u\prec_{\mathcal{A}}v$, we have $\beta\prec\gamma$. From $v\in G^{\prec}(\alpha)$ we obtain $\gamma\prec\alpha$, so we conclude $\beta\prec\alpha$.
  4. 4.
    Since $v\in G^{\prec}_{\dashv}(\alpha)$ , we have either $v\in G^{\prec}(\alpha)$ or $v\in G_{\dashv}(\alpha)$. If $v\in G^{\prec}(\alpha)$, then $u\in G^{\prec}(\alpha)$ by the previous point and so $u\in G^{\prec}_{\dashv}(\alpha)$. Now assume that $v\in G_{\dashv}(\alpha)$. If $u\in G_{\dashv}(\alpha)$, then $u\in G^{\prec}_{\dashv}(\alpha)$ and we are done, so we can assume $u\not\in G_{\dashv}(\alpha)$. Let us prove that it must be $u\in G^{\prec}(\alpha)$, which again implies $u\in G^{\prec}_{\dashv}(\alpha)$. Fix $\beta\in I_{u}$; we must prove that $\beta\prec\alpha$. Since $v\in G_{\dashv}(\alpha)$, then there exists $\gamma\in\Sigma^{*}$ such that $\gamma\alpha\in I_{v}$. From $u\prec_{\mathcal{A}}v$ we obtain $\beta\prec\gamma\alpha$. Since $u\not\in G_{\dashv}(\alpha)$ implies $\lnot(\alpha\dashv\beta)$, from $\beta\prec\gamma\alpha$ we conclude $\beta\prec\alpha$.
Lemma B.3.
[…]
Proof B.4.
()($\Rightarrow$)( ⇒ ). Let us prove (i). Pick $\rho\in\lambda(u)$. Then, there exists $u^{\prime}\in Q$ such that $(u^{\prime},u,\rho)\in E$. We must prove that $\rho\prec\alpha$. Let $\beta^{\prime}\in I_{u^{\prime}}$; from $(u^{\prime},u,\rho)\in E$ we obtain $\beta^{\prime}\rho\in I_{u}$, so from $u\in G^{\prec}(\alpha)$ we obtain $\beta^{\prime}\rho\prec\alpha$, which implies $\rho\prec\alpha$. Let us prove (ii). Assume that $u^{\prime}\in Q$ is such that that $(u^{\prime},u,s(\alpha,k))\in E$ for some $0<k<|\alpha|$. We must prove that $u^{\prime}\in G^{\prec}(p(\alpha,|\alpha|-k))$. Pick $\gamma^{\prime}\in I_{u^{\prime}}$. We must prove that $\gamma^{\prime}\prec p(\alpha,|\alpha|-k)$. Since $(u^{\prime},u,s(\alpha,k))\in E$, we have $\gamma^{\prime}s(\alpha,k)\in I_{u}$, so from $u\in G^{\prec}(\alpha)$ we obtain $\gamma^{\prime}s(\alpha,k)\prec\alpha$, which implies $\gamma^{\prime}\prec p(\alpha,|\alpha|-k)$.
[…]
Statement of Lemma 4.8. Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler GDFA, and let $\alpha\in\Sigma^{*}$. For $0<k<|\alpha|$, let $f_{k}=\mathtt{out}(Q[1,|G^{\prec}(p(\alpha,|\alpha|-k))|],s(\alpha,k))$. Then, $|G^{\prec}(\alpha)|$ is the largest integer $0\leq j\leq|Q|$ such that (i) $\mathtt{in}(Q[1,j],s(\alpha,k))\leq f_{k}$, for every $0<k<|\alpha|$, and (ii) if $j\geq 1$, then $\max_{\lambda(Q[j])}\prec\alpha$.
Proof B.5.
First, note that by Lemma 4.6 we have $G^{\prec}(\alpha)=Q[1,|G^{\prec}(\alpha)|]$ and, for every $0<k<\alpha$, we have $G^{\prec}(p(\alpha,|\alpha|-k))=Q[1,|G^{\prec}(p(\alpha,|\alpha|-k)|]$. Let $j^{*}$ be the largest integer $0\leq j\leq|Q|$ such that (i) $\mathtt{in}(Q[1,j],s(\alpha,k))\leq f_{k}$, for every $0<k<|\alpha|$, and (ii) if $j\geq 1$, then $\max_{\lambda(Q[j])}\prec\alpha$. We want to prove that $|G^{\prec}(\alpha)|=j^{*}$.
()($\leq$)( ≤ ) The conclusion is immediate if $|G^{\prec}(\alpha)|=0$, so we can assume $|G^{\prec}(\alpha)|\geq 1$. It will suffice to prove that $\mathtt{in}(Q[1,|G^{\prec}(\alpha)|],s(\alpha,k))\leq f_{k}$, for every $0<k<|\alpha|$, and $\max_{\lambda(Q[|G^{\prec}(\alpha)|])}\prec\alpha$. This follows from Lemma B.3 and the definition of $f_{k}$.
()($\geq$)( ≥ ) We only have to prove that if $|G^{\prec}(\alpha)|+1\leq j\leq|Q|$, then either $\mathtt{in}(Q[1,j],s(\alpha,k))>f_{k}$ for some $0<k<|\alpha|$ or $a\preceq\max_{\lambda(Q[j])}$. Assume that $\max_{\lambda(Q[j])}\prec\alpha$. Since $Q[j]\not\in G^{\prec}(\alpha)$, by Lemma B.3 for some $0<k<|\alpha|$ there exists $v^{\prime}\in Q$ such that $(v^{\prime},Q[j],s(\alpha,k))\in E$ and $v^{\prime}\not\in Q[1,|G^{\prec}(p(\alpha,|\alpha|-k))|]$. We will conclude that $\mathtt{in}(Q[1,j],s(\alpha,k))>f_{k}$ if we show that if $u^{\prime},u\in Q$ are such that $u^{\prime}\in Q[1,|G^{\prec}(p(\alpha,|\alpha|-k))|]$ and $(u^{\prime},u,s(\alpha,k))\in E$, then $u\in Q[1,j]$. Suppose for sake of contradiction that $u\not\in Q[1,j]$. This implies that $Q[j]\prec_{\mathcal{A}}u$, so from $(v^{\prime},Q[j],s(\alpha,k)),(u^{\prime},u,s(\alpha,k))\in E$ and Lemma 4.2 we obtain $v^{\prime}\prec_{\mathcal{A}}u^{\prime}$. Since $u^{\prime}\in Q[1,|G^{\prec}(p(\alpha,|\alpha|-k))|]$, we conclude $v^{\prime}\in Q[1,|G^{\prec}(p(\alpha,|\alpha|-k))|]$, a contradiction.
Statement of Lemma 4.9. Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler GDFA, and let $\alpha\in\Sigma^{*}$. For $0<k<|\alpha|$, let $f_{k}=\mathtt{out}(Q[1,|G^{\prec}(p(\alpha,|\alpha|-k))|],s(\alpha,k))$ and $g_{k}=\mathtt{out}(Q[1,|G^{\prec}_{\dashv}(p(\alpha,|\alpha|-k))|],s(\alpha,k))$. Then, $g_{k}\geq f_{k}$ for every $0<k<|\alpha|$. Moreover, $|G^{\prec}_{\dashv}(\alpha)|$ is equal to the maximum among:
  1. 1.
    $|G^{\prec}(\alpha)|$;
  2. 2.
    the largest integer $0\leq i\leq|Q|$ such that, if $i\geq 1$, then $Q[i]\in G^{*}(\alpha)$;
  3. 3.
    the smallest integer $0\leq j\leq|Q|$ such that, for every $0<k<|\alpha|$ such that $g_{k}>f_{k}$, we have $\mathtt{in}(Q[1,j],s(\alpha,k))\geq g_{k}$.
Proof B.6.
First, note that by Lemma 4.6 we have $G^{\prec}(\alpha)=Q[1,|G^{\prec}(\alpha)|]$, $G^{\prec}_{\dashv}(\alpha)=Q[1,|G^{\prec}_{\dashv}(\alpha)|]$ and, for every $0<k<\alpha$, we have $G^{\prec}(p(\alpha,|\alpha|-k))=Q[1,|G^{\prec}(p(\alpha,|\alpha|-k)|]$ and $G^{\prec}_{\dashv}(p(\alpha,|\alpha|-k))=Q[1,|G^{\prec}_{\dashv}(p(\alpha,|\alpha|-k)|]$. For every $0<k<\alpha$ we have $g_{k}\geq f_{k}$ because $G^{\prec}(p(\alpha,|\alpha|-k)\subseteq G^{\prec}_{\dashv}(p(\alpha,|\alpha|-k)$. Let $i^{*}$ be the largest integer $0\leq i\leq|Q|$ such that, if $i\geq 1$, then $Q[i]\in G^{*}(\alpha)$, and let $j^{*}$ be the smallest integer $0\leq j\leq|Q|$ such that $\mathtt{in}(Q[1,j],s(\alpha,k))\geq g_{k}$ for every $0<k<|\alpha|$ such that $g_{k}>f_{k}$. Let $l=\max(|G^{\prec}(\alpha)|,i^{*},j^{*})$. We want to prove that $l=|G^{\prec}_{\dashv}(\alpha)|$.
()($\leq$)( ≤ ) We have to prove that $|G^{\prec}(\alpha)|\leq|G^{\prec}_{\dashv}(\alpha)|$, $i^{*}\leq|G^{\prec}_{\dashv}(\alpha)|$ and $j^{*}\leq|G^{\prec}_{\dashv}(\alpha)|$. The inequality $|G^{\prec}(\alpha)|\leq|G^{\prec}_{\dashv}(\alpha)|$ follows from $G^{\prec}(\alpha)\subseteq G^{\prec}_{\dashv}(\alpha)$. Let us prove that $i^{*}\leq|G^{\prec}_{\dashv}(\alpha)|$. If $i^{*}=0$ we are done, so we can assume $i^{*}\geq 1$. We know that $Q[i^{*}]\in G^{*}(\alpha)$, so $Q[i^{*}]\in G_{\dashv}(\alpha)$ and $Q[i^{*}]\in G^{\prec}_{\dashv}(\alpha)$, which implies $i^{*}\leq|G^{\prec}_{\dashv}(\alpha)|$. Let us prove that $j^{*}\leq|G^{\prec}_{\dashv}(\alpha)|$. If $j^{*}=0$ we are done, so we can assume $j^{*}\geq 1$. We know that there exists $0<k<|\alpha|$, with $g_{k}>f_{k}$, such that $\mathtt{in}(Q[1,j^{*}-1],s(\alpha,k))<g_{k}$ but $\mathtt{in}(Q[1,j^{*}],s(\alpha,k))\geq g_{k}$. We are only left with showing that there exists $u^{\prime}\in G_{\dashv}(p(\alpha,|\alpha|-k))$ such that $(u^{\prime},Q[j^{*}],s(\alpha,k))\in E$, because this will imply $Q[j^{*}]\in G_{\dashv}(\alpha)$, so $Q[j^{*}]\in G^{\prec}_{\dashv}(\alpha)$ and $j^{*}\leq|G^{\prec}_{\dashv}(\alpha)|$. Suppose for sake of contradiction that for every $u^{\prime}\in Q$ such that $(u^{\prime},Q[j^{*}],s(\alpha,k))\in E$ we have $u^{\prime}\not\in G_{\dashv}(p(\alpha,|\alpha|-k))$. First, let us prove that there can not exist $u^{\prime}_{1}\in Q[1,|G^{\prec}(p(\alpha,|\alpha|-k))|]$ and $u^{\prime}_{2}\in Q[|G^{\prec}_{\dashv}(p(\alpha,|\alpha|-k))|+1,|Q|]$ such that $(u^{\prime}_{1},Q[j^{*}],s(\alpha,k))\in E$ and $(u^{\prime}_{2},Q[j^{*}],s(\alpha,k))\in E$. Suppose for sake of contradiction that such $u^{\prime}_{1}$ and $u^{\prime}_{2}$ exist. We will obtain a contradiction by showing that there exists no $u^{\prime}_{3}\in G_{\dashv}(p(\alpha,|\alpha|-k))$ such that for some $u_{3}\in Q$ we have $(u^{\prime}_{3},u_{3},s(\alpha,k))\in E$, because this would imply $g_{k}=f_{k}$, which contradicts $g_{k}>f_{k}$. Indeed (i) it cannot be $u_{3}=Q[j]$ because we know that for every $u^{\prime}\in Q$ such that $(u^{\prime},Q[j^{*}],s(\alpha,k))\in E$ we have $u^{\prime}\not\in G_{\dashv}(p(\alpha,|\alpha|-k))$, (ii) it cannot be $u_{3}\prec_{\mathcal{A}}Q[j]$ because by Lemma 4.2 we would obtain $u^{\prime}_{3}\prec_{\mathcal{A}}u^{\prime}_{1}$, and since $u^{\prime}_{1}\in Q[1,|G^{\prec}(p(\alpha,|\alpha|-k))|]$, we would conclude $u^{\prime}_{3}\in Q[1,|G^{\prec}(p(\alpha,|\alpha|-k))|]$, and (iii) it cannot be $Q[j]\prec_{\mathcal{A}}u_{3}$ because by Lemma 4.2 we would obtain $u^{\prime}_{2}\prec_{\mathcal{A}}u^{\prime}_{3}$, and since $u^{\prime}_{2}\in Q[|G^{\prec}_{\dashv}(p(\alpha,|\alpha|-k))|+1,|Q|]$, we would conclude $u^{\prime}_{3}\in Q[|G^{\prec}_{\dashv}(p(\alpha,|\alpha|-k))|+1,|Q|]$. As a consequence, one of the following two cases must occur:
  1. 1.
    For every $u^{\prime}_{1}\in Q$ such that $(u^{\prime}_{1},Q[j^{*}],s(\alpha,k))\in E$ we have $u^{\prime}_{1}\in Q[1,|G^{\prec}(p(\alpha,|\alpha|-k))|]$. Let us prove that if $v^{\prime},v\in Q$ are such that $v\in Q[1,j^{*}]$ and $(v^{\prime},v,s(\alpha,k))\in E$, then $v^{\prime}\in Q[1,|G^{\prec}(p(\alpha,|\alpha|-k))|]$. This will imply $\mathtt{in}(Q[1,j^{*}],s(\alpha,k))\leq f_{k}<g_{k}$, the desired contradiction. If $v=Q[j^{*}]$, the conclusion follows because we know that for every $u^{\prime}_{1}\in Q$ such that $(u^{\prime}_{1},Q[j^{*}],s(\alpha,k))\in E$ we have $u^{\prime}_{1}\in Q[1,|G^{\prec}(p(\alpha,|\alpha|-k))|]$. Now, assume $v\prec_{\mathcal{A}}Q[j]$. Since $\mathtt{in}(Q[1,j^{*}-1],s(\alpha,k))<g_{k}$ but $\mathtt{in}(Q[1,j^{*}],s(\alpha,k))\geq g_{k}$, then there exists $u^{*}_{1}\in Q$ such that $(u^{*}_{1},Q[j^{*}],s(\alpha,k))\in E$, and we know that it must be $u^{*}_{1}\in Q[1,|G^{\prec}(p(\alpha,|\alpha|-k))|]$. From $v\prec_{\mathcal{A}}Q[j]$ and Lemma 4.2 we obtain $v^{\prime}\prec_{\mathcal{A}}u^{*}_{1}$, and since $u^{*}_{1}\in Q[1,|G^{\prec}(p(\alpha,|\alpha|-k))|]$, we conclude $v^{\prime}\in Q[1,|G^{\prec}(p(\alpha,|\alpha|-k))|]$.
  2. 2.
    For every $u^{\prime}_{2}\in Q$ such that $(u^{\prime}_{2},Q[j^{*}],s(\alpha,k))\in E$ we have $u^{\prime}_{2}\in Q[|G^{\prec}_{\dashv}(p(\alpha,|\alpha|-k))|+1,|Q|]$. Let us prove that if $v^{\prime},v\in Q$ are such that $v^{\prime}\in Q[1,|G^{\prec}_{\dashv}((p(\alpha,|\alpha|-k))|]$ and $(v^{\prime},v,s(\alpha,k))\in E$, then $v\in Q[1,j^{*}-1]$. This will imply $\mathtt{in}(Q[1,j^{*}-1],s(\alpha,k))\geq g_{k}$, the desired contradiction. It cannot be $v=Q[j^{*}]$, because this would imply $v^{\prime}\in Q[|G^{\prec}_{\dashv}(p(\alpha,|\alpha|-k))|+1,|Q|]$. Now, assume for sake of contradiction that $Q[j^{*}]\prec_{\mathcal{A}}v$. Since $\mathtt{in}(Q[1,j^{*}-1],s(\alpha,k))<g_{k}$ but $\mathtt{in}(Q[1,j^{*}],s(\alpha,k))\geq g_{k}$, then there exists $u^{*}_{2}\in Q$ such that $(u^{*}_{2},Q[j^{*}],s(\alpha,k))\in E$, and we know that it must be $u^{*}_{2}\in Q[|G^{\prec}_{\dashv}(p(\alpha,|\alpha|-k))|+1,|Q|]$. From $Q[j^{*}]\prec_{\mathcal{A}}v$ and Lemma 4.2 we obtain $u^{*}_{2}\prec_{\mathcal{A}}v^{\prime}$, and since $u^{*}_{2}\in Q[|G^{\prec}_{\dashv}(p(\alpha,|\alpha|-k))|+1,|Q|]$ , we conclude $v^{\prime}\in Q[|G^{\prec}_{\dashv}(p(\alpha,|\alpha|-k))|+1,|Q|]$, a contradiction.
()($\geq$)( ≥ ) We only have to prove that if $l+1\leq j\leq|Q|$, then $Q[j]\not\in G^{\prec}_{\dashv}(\alpha)$. Since $j\geq l+1>|G^{\prec}(\alpha)|$, we have $Q[j]\not\in G^{\prec}(\alpha)$, hence we are left with showing that $Q[j]\not\in G_{\dashv}(\alpha)$. Assume for sake of contradiction that $Q[j]\in G_{\dashv}(\alpha)$. It cannot be $Q[j]\in G^{*}(\alpha)$, otherwise $j\leq i^{*}\leq l$, a contradiction. Hence, $Q[j]\in G_{\dashv}(\alpha)\setminus G^{*}(\alpha)$. This means that for some $0<k<|\alpha|$ there exists $u^{\prime}\in Q$ such that $(u^{\prime},Q[j],s(\alpha,k))\in E$ and $u^{\prime}\in G_{\dashv}(p(\alpha,|\alpha|-k))$, which implies $g_{k}>f_{k}$. Let us prove that $\mathtt{in}(Q[1,j-1],s(\alpha,k))<g_{k}$. Assume that $(v^{\prime},v,s(\alpha,k))\in E$ is such that $v\in Q[1,j-1]$. Since $v\prec_{\mathcal{A}}Q[j]$, from Axiom 2 we obtain $v^{\prime}\prec_{\mathcal{A}}u^{\prime}$, and since $u^{\prime}\in G_{\dashv}(p(\alpha,|\alpha|-k))$ and so $u^{\prime}\in Q[1,|G^{\prec}_{\dashv}(p(\alpha,|\alpha|-k))|]$ we obtain $v^{\prime}\in Q[1,|G^{\prec}_{\dashv}(p(\alpha,|\alpha|-k))|]$. This implies that $\mathtt{in}(Q[1,j-1],s(\alpha,k))\leq g_{k}$, and it must be $\mathtt{in}(Q[1,j-1],s(\alpha,k))<g_{k}$ because $(u^{\prime},Q[j],s(\alpha,k))\in E$ and $u^{\prime}\in G_{\dashv}(p(\alpha,|\alpha|-k))$. We conclude $j\leq j^{*}\leq l$, a contradiction.
Now, we want to show how to build the FM-index of a Wheeler GDFA (Theorem 4.13). First, we will present a lemma which shows that in Lemma 4.8 and Lemma 4.9, when we have a statement such as "for every $k$…", we can always assume that $k$ takes at most $r$ values, if $\mathcal{A}$ is a $r$-GDFA.
Lemma B.7.
Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler $r$-GDFA, and let $\alpha\in\Sigma^{*}$. For $0<k<\min\{r+1,|\alpha|\}$, let $f_{k}=\mathtt{out}(Q[1,|G^{\prec}(p(\alpha,|\alpha|-k))|],s(\alpha,k))$ and $g_{k}=\mathtt{out}(Q[1,|G^{\prec}_{\dashv}(p(\alpha,|\alpha|-k))|],s(\alpha,k))$. Then, $g_{k}\geq f_{k}$ for every $0<k<\min\{r+1,|\alpha|\}$. Moreover:
  1. 1.
    $|G^{\prec}(\alpha)|$ is the largest integer $0\leq j\leq|Q|$ such that (i) $\mathtt{in}(Q[1,j],s(\alpha,k))\leq f_{k}$, for every $0<k<\min\{r+1,|\alpha|\}$, and (ii) every $\rho^{\prime}\in\Sigma^{k}$ for which there exists $1\leq\ell\leq j$ with $\rho^{\prime}\in\lambda(Q[\ell])$ satisfies $\rho^{\prime}\preceq w_{\alpha,k}$ with $w_{\alpha,k}$ being the largest string in $\Sigma^{k}$ smaller than $\alpha$, for every $10<k<r+$ 1.
  2. 2.
    $|G^{\prec}_{\dashv}(\alpha)|$ is equal to the maximum among:
    1. (a)
      $|G^{\prec}(\alpha)|$;
    2. (b)
      the largest integer $0\leq i\leq|Q|$ such that, if $i\geq 1$, then for some $10<k<r+$ 1 there exists $\rho^{\prime}\in\Sigma^{k}\cap\lambda(Q[h])$ with $\alpha\dashv\rho^{\prime}$.
    3. (c)
      the smallest integer $0\leq j\leq|Q|$ such that, for every $0<k<\min\{r+1,|\alpha|\}$ such that $g_{k}>f_{k}$, we have $\mathtt{in}(Q[1,j],s(\alpha,k))\geq g_{k}$.
Proof B.8.
  1. 1.
    We want to prove that the statement of the corollary is equivalent to the statement of Lemma 4.8. For every $\min\{r+1,|\alpha|\}\leq k<|\alpha|$ we have $|s(\alpha,k)|>r$, so $f_{k}=0$ and $\mathtt{in}(Q,s(\alpha,k))=0$ because $\mathcal{A}$ is a $r$-GDFA. This implies that condition (i) in the statement correctly replaces condition (i) in Lemma 4.8. We are only left with showing that condition (ii) in the statement correctly replaces condition (ii) in Lemma 4.8. This follows immediately from Lemma 4.2.
  2. 2.
    Let us prove that the three conditions in the statement are equivalent to the three conditions in Lemma 4.9.
    1. (a)
      The first condition is the same.
    2. (b)
      The second condition is equivalent because $\mathcal{A}$ is a $r$-GDFA.
    3. (c)
      The third condition is equivalent because if $\min\{r+1,|\alpha|\}\leq k<|\alpha|$, then $|s(\alpha,k)|>r$, so $g_{k}=f_{k}=0$.
Let us recall the following lemmas for compressed data structures (we follow the approach in [14]). Recall that the zero-order entropy of a sequence $S\in\Sigma^{n}$ of length $n$ over alphabet $\Sigma$ is $H_{0}(S)=\sum_{c\in\Sigma}\frac{|S|_{c}}{n}\log_{2}\frac{n}{|S|_{c}}$, where $|S|_{c}$ denotes the number of occurrences of character $c$ in $S$.
Lemma B.9 (Rank/Select, see [14]).
Let $S\in\Sigma^{n}$ be a string over an integer alphabet $\Sigma=[0,\sigma-1]$ of size $\sigma\leq n$. Then, there exists a data structure of $nH_{0}(S)(1+o(1))+O(n)$ bits supporting the following operations in time $O(\log\log\sigma)$:
  • Access: $S[i]$, for any $1\leq i\leq n$.
  • Rank: $S.rank(i,c)=|\{j\in\{1,\dots,i\}\ |\ S[j]=c\}|$, for any $1\leq i\leq n$ and $c\in\Sigma$.
  • Select: $S.select(i,c)$ equals the integer $j$ such that $S[j]=c$ and $S.rank(j,c)=i$, for any $1\leq i\leq S.rank(n,c)$ and $c\in\Sigma$.
Lemma B.10 (Fully-indexable succinct dictionaries, see [14]).
A set $A=\{x_{1},\dots,x_{n}\}\subseteq[1,u]$ of cardinality $n$ can be represented with a data structure of $n\log(u/n)+O(n)$ bits so that the following operations can be implemented in $O(\log\log(u/n))$ time:
  • Rank: $A.rank(x)=|\{y\in A\ |\ y\leq x\}|$, for any $1\leq x\leq u$.
  • Select: $A.select(i)=x$ such that $x\in A$ and $A.rank(x)=i$, for any $1\leq i\leq|A|$.
In particular, in $O(\log\log(u/n))$ time we can also solve the following queries:
  • Predecessor: the largest element of $A$ smaller than or equal to $x$, if it exists. For any $1\leq x\leq u$, $A.pred(x)=A.select(A.rank(x))$ if $A.rank(x)>0$, and $A.pred(x)=\bot$ otherwise.
  • Strict-Successor: the smallest element of $A$ strictly greater than $x$, if it exists. For any $1\leq x\leq u$, $A.succ(x)=A.select(A.rank(x)+1)$ if $A.rank(x)<|A|$, and $A.succ(x)=\bot$ otherwise.
  • Membership: For any $1\leq x\leq u$, $x\in A$ if and only if $x=A.pred(x)$.
Statement of Theorem 4.13. Let $\mathcal{A}$ be a Wheeler $r$-GDFA. Then, we can encode $\mathcal{A}$ by using $\mathfrak{e}\log\sigma(1+o(1))+O(e+rn)$ bits so that later on, given a pattern $\alpha\in\Sigma^{*}$ of length $m$, we can solve the SMLG problem on $\mathcal{A}$ in $O(rm(\log r+\log\log\sigma))$ time. Within the same time bound we can also decide whether $\alpha\in\mathcal{L(A)}$.
Proof B.11.
We will use some well-known properties of the entropy $H_{0}(S)$ of a string $S\in\Sigma^{h}$: the quantity $hH_{0}(S)$ is a lower bound to the length of any encoding of $S$ that encodes each character independently from the others via a prefix code of the alphabet $\Sigma$, and in particular $H_{0}(S)\leq\log_{2}\sigma$, where $\sigma=|\Sigma|$. We store the following data structures.
  • One fully-indexable succinct dictionary (Lemma B.10) on each $\Sigma(i)$. The total number of required bits is $\sum_{i=1}^{r}(\sigma_{i}\log(\sigma^{i}/\sigma_{i})+O(\sigma_{i}))\leq\sum_{i=1}^{r}(e_{i}\log(\sigma^{i}/\sigma_{i})+O(e_{i}))=\mathfrak{e}\log\sigma-(\sum_{i=1}^{r}e_{i}\log\sigma_{i})+O(e)$. As a consequence, we can solve rank, select, predecessor, strict-successor and membership queries on each dictionary in $O(\log\log(\sigma^{i}/\sigma_{i}))\subseteq O(\log\log\sigma^{r})=O(\log r+\log\log\sigma)$ time.
  • The bit string $\texttt{FIN}\in\{0,1\}^{n}$ of Definition 4.10 represented by the data structure of Lemma B.9. The number of required bits is $nH_{0}(\texttt{FIN})(1+o(1))+O(n)\subseteq O(n)$.
  • For every $1\leq i\leq r$, the bit string $\mathtt{OUT}_{i}\in\{0,1\}^{e_{i}+n}$ of Definition 4.10 represented by the data structure of Lemma B.9. The total number of required bits is $\sum_{i=1}^{r}((e_{i}+n)H_{0}(\mathtt{OUT}_{i})(1+o(1))+(e_{i}+n)\cdot O(1))=O(e+nr)$. As a consequence, we can solve rank and select queries on each $\mathtt{OUT}_{i}$ in $O(1)$ time.
  • For every $1\leq i\leq r$, the bit string $\mathtt{IN}_{i}\in\{0,1\}^{e_{i}+n}$ of Definition 4.10 represented by the data structure of Lemma B.9. The total number of required bits is again $O(e+nr)$. As a consequence, we can solve rank and select queries on each $\mathtt{IN}_{i}$ in $O(1)$ time.
  • For every $1\leq i\leq r$, the string $\mathtt{LAB}_{i}\in(\Sigma^{i})^{e_{i}}$ of Definition 4.10 represented by the data structure of Lemma B.9. The total number of required bits is $\sum_{i=1}^{r}(e_{i}H_{0}(\mathtt{LAB}_{i})(1+o(1))+O(e_{i}))\leq(\sum_{i=1}^{r}(e_{i}\log\sigma_{i})(1+o(1)))+O(e)\leq(\sum_{i=1}^{r}e_{i}\log\sigma_{i})+\mathfrak{e}\log\sigma\cdot o(1)+O(e)$. As a consequence, we can solve rank and select queries on each $\mathtt{LAB}_{i}$ in $O(\log\log\sigma^{i})\subseteq O(\log\log\sigma^{r})=O(\log r+\log\log\sigma)$ time.
  • For every $1\leq i\leq r$, a bit string $\mathtt{AUX}_{i}\in\{0,1\}^{e_{i}}$ represented by the data structure of Lemma B.9 which stores auxiliary information as follows. Sort all edges in $E(i)$ by end state and, if the end state is the same, by label. Build a string $\mathtt{AUX}^{\prime}_{i}\in(\Sigma^{i})^{e_{i}}$ by concatenating all labels of the sorted edges. Note that by Lemma 4.2, if $11\leq k\leq e_{i}-$ 1, then $\mathtt{AUX}^{\prime}_{i}[k]\preceq\mathtt{AUX}^{\prime}_{i}[k+1]$. Finally, $\mathtt{AUX}_{i}\in\{0,1\}^{e_{i}}$ is the bit string such that $\mathtt{AUX}[k]=1$ if and only if $k=1$ or $\mathtt{AUX}^{\prime}_{i}[k]\neq\mathtt{AUX}^{\prime}_{i}[k-1]$, for every $1\leq k\leq e_{i}$. The total number of required bits is $O(e)$. As a consequence, we can solve rank and select queries on each $\mathtt{AUX}_{i}$ in $O(1)$ time.
By adding up the space of all components (note that the terms $-\sum_{i=1}^{r}e_{i}\log\sigma_{i}$ in the dictionaries $\Sigma(i)$’s and $\sum_{i=1}^{r}e_{i}\log\sigma_{i}$ in the sequences $\mathtt{LAB}_{i}$’s cancel out), we conclude that our data structures take at most $\mathfrak{e}\log\sigma(1+o(1))+O(e+nr)$ bits. In particular, we are storing (an augmented variant of the) Burrows-Wheeler Transform of $\mathcal{A}$, so by Theorem 4.11 we are storing an encoding of $\mathcal{A}$.
In order to solve the SMLG problem on input $\alpha\in\Sigma^{m}$, we need to compute $G^{\prec}_{\dashv}(\alpha)$. By Lemma 4.6, we only need to compute $|G^{\prec}(\alpha)|$ and $|G_{\dashv}(\alpha)|$. In $O(m)$ steps, we will recursively compute $|G^{\prec}(p(\alpha,k))|$ and $|G_{\dashv}(p(\alpha,k))|$ for every $0\leq k\leq|\alpha|$. Note that, if $k=0$, then $|G^{\prec}(\epsilon)|=0$ and $|G_{\dashv}(\epsilon)|=|Q|=n$. Now, let us show how to compute $|G^{\prec}(\alpha)|$ and $|G_{\dashv}(\alpha)|$, assuming that we have already computed $|G^{\prec}(p(\alpha,k))|$ and $|G_{\dashv}(p(\alpha,k))|$ for every $0\leq k<|\alpha|$. From Lemma B.7, we obtain that we only need to solve each of the following queries at most $O(r)$ times:
  • (op1) for any $1\leq k\leq|Q|$, for every $1\leq i\leq r$ and for any $\rho\in\Sigma^{i}$, compute $\mathtt{out}(Q[1,k],\rho)$;
  • (op2) for any $1\leq i\leq r$, for any $\rho\in\Sigma^{i}$ and for any $h\geq 0$, find the largest integer $0\leq k\leq|Q|$ such that $\mathtt{in}(Q[1,k],\rho)\leq h$;
  • (op3) for any $1\leq i\leq r$, for any $\rho\in\Sigma^{i}$ and for any $z\geq 1$, find the smallest integer $1\leq t\leq|Q|$ such that $\mathtt{in}(Q[1,t],\rho)\geq z$, if it exists, otherwise report that it does not exist.
  • (op4) for any $1\leq i\leq r$ and for any $\rho\in\Sigma^{i}$, find the largest integer $0\leq h\leq|Q|$ such that $\rho^{\prime}\preceq\rho$ for every $\rho^{\prime}\in\Sigma^{i}$ for which there exists $1\leq j\leq h$ with $\rho^{\prime}\in\lambda(Q[j])$.
  • (op5) for any $\rho\in\Sigma^{*}$, with $|\rho|\leq r$, and for any $|\rho|\leq i\leq r$, find the largest integer $0\leq h\leq|Q|$ such that, if $h\geq 1$, then there exists $\rho^{\prime}\in\Sigma^{i}\cap\lambda(Q[h])$ with $\rho\dashv\rho^{\prime}$.
As a consequence, we are only left with showing that we can solve each of these queries in $O(\log r+\log\log\sigma)$ time.
(op1). The number of edges in $E(i)$ leaving a state in $Q[1,k]$ can be obtained by computing $d=\mathtt{OUT}_{i}.rank(\mathtt{OUT}_{i}.select(k,1),0)$. As a consequence, we have $\mathtt{out}(Q[1,k],\rho)=\mathtt{LAB}_{i}.rank(d,\rho)$.
(op2). The number of edges labeled $\rho$ is $t=\mathtt{LAB}_{i}.rank(e_{i},\rho)$. If $t\leq h$, then the largest $k$ with the desired properties is $k=|Q|$. Now assume that $t>h$. The number of strings smaller than or equal to $\rho$ occurring in $\Sigma(i)$ is $\Sigma(i).rank(\rho)$ and the $h+1$-th edge labeled $\rho$ in the ordering given by $\mathtt{AUX}^{\prime}_{i}$ is the $g$-th edge, where $g=\mathtt{AUX}.rank(\Sigma(i).rank(\rho),1)+h$. This edge reaches state $Q[f]$, where $f=\mathtt{IN}_{i}.rank(\mathtt{IN}_{i}.select(g,0),1)+1$, and so the largest $k$ with the desired properties if $f-1$.
(op3). We simply use operation (op2) to compute the largest integer $0\leq k\leq|Q_{i}|$ such that $\mathtt{in}(Q[1,k],\rho)\leq z-1$. If $k=|Q_{i}|$, then the desired integer does not exist, otherwise it is equal to $k+1$.
(op4). We first check whether $\Sigma(i).succ(\rho)$ is defined. If it is not defined, then the largest $h$ with the desired properties is $|Q|$. Now assume that $\Sigma(i).succ(\rho)$ is defined. Then by Lemma 4.2 the largest integer with the desired property is simply the largest integer $0\leq k\leq|Q_{i}|$ such that $\mathtt{in}(Q[1,k],\Sigma(i).succ(\rho))\leq 0$, which can be computed using (op2).
(op5). Let $c\in\Sigma$ be the largest character in (Σ,)normal-$(\Sigma,\preceq)$. Let $a=\Sigma(i).pred(c^{i-|\rho|}\rho)$ If $a$ is not suffixed by $\rho$ or it is not defined, then the largest $h$ with the desired properties is $h=0$. Otherwise, by Lemma 4.2 we have that $h$ is the largest integer such that $a\in\lambda(Q[h])$. The number of edges labeled $a$ is $d=\mathtt{LAB}_{i}.rank(e_{i},\rho)$, so $h$ is the smallest integer $t$ such that $in(Q[1,t],\rho)\geq d$, which can be computed using (op3).
Now, let us prove that in $O(rm(\log r+\log\log\sigma))$ time we can also decide whether $\alpha\in\mathcal{L(A)}$. We claim that the same algorithm works, as long as we artificially start with $|G^{\prec}(\epsilon)|=0$ and $|G_{\dashv}(\epsilon)|=1$ (instead of $|G^{\prec}(\epsilon)|=0$ and $|G_{\dashv}(\epsilon)|=n$). Indeed, consider the GDFA $\mathcal{A}$’ obtained from $A$ by adding a new initial state $s^{\prime}$ and adding the edge $(s^{\prime},s,\#)$, where $s$ is the old initial state and #normal-#\## is a new character, smaller than every character in $\Sigma$. Then, $\mathcal{A}^{\prime}$ is also Wheeler (note that $\preceq_{\mathcal{A}}$ is contained in $\preceq_{\mathcal{A}}^{\prime}$), and deciding whether $\alpha\in\mathcal{L(A)}$ is equivalent to solving the SMLG problem on input $\mathcal{A}$’ and #$\#\alpha$ (and checking whether the unique reached state is in $F$, if it exists), which after one step yields $G^{\prec}(\#)=\{s^{\prime},s\}$ and $G_{\dashv}(\#)=\{s\}$, so the conclusion follows after checking whether the unique reached state is in $F$ by using $\mathtt{FIN}$.
Appendix C The FM-index: from GDFAs to GNFAs without $\epsilon$-transitions
The aim of this section is to show that Theorem 4.13 also holds for GNFAs without $\epsilon$-transitions, thus leading to Theorem 1.3. To this end, we will extend the notion of Wheelerness to GNFAs without $\epsilon$-transitions.
Let us recall the definition of Wheeler NFA. An NFA $\mathcal{A}$ is Wheeler if there exists a total order $\leq$ on $Q$ such that (i) $s$ comes first in the total order, (ii) for every $(u^{\prime},u,a),(v^{\prime},v,b)\in E$, if $u<v$, then $a\preceq b$ and (iii) for every $(u^{\prime},u,a),(v^{\prime},v,a)\in E$, if $u<v$, then $u^{\prime}\leq v^{\prime}$. Alanko et al. [3, Lemma 2.3] proved that, if $u<v$ in the total order, then $(\forall\alpha\in I_{u})(\forall\beta\in I_{v})((\{\alpha,\beta\}\not\subseteq I_{u}\cap I_{v})\to(\alpha\prec\beta))$. Let us see how to extend Definition 4.1 to GNFAs without $\epsilon$-transitions. We will first generalize the definition of $\preceq_{\mathcal{A}}$ from GDFAs to GNFAs without $\epsilon$-transitions drawing inspiration from Alanko et al.’s result.
Let $\mathcal{A}=(Q,E,s,F)$ be a GNFA without $\epsilon$-transitions. Let $\preceq_{\mathcal{A}}$ be the relation on $Q$ such that, for every $u,v\in Q$, we have $u\preceq_{\mathcal{A}}v$ if and only if $(\forall\alpha\in I_{u})(\forall\beta\in I_{v})((\{\alpha,\beta\}\not\subseteq I_{u}\cap I_{v})\to(\alpha\prec\beta))$. If $\mathcal{A}$ is a GDFA, then $\preceq_{\mathcal{A}}$ reduces to the definition given in Section 4, because for every $u,v\in Q$ with $u\not=v$ we have $I_{u}\cap I_{v}=\emptyset$ (see Remark 3.3). We have seen that, if $\mathcal{A}$ is a GDFA, then $\preceq_{\mathcal{A}}$ is a partial order. If $\mathcal{A}$ is an arbitrary GNFA without $\epsilon$-transitions, in general $\preceq_{\mathcal{A}}$ is only a preorder, that is, it is a reflexive and transitive relation, but it need not be antisymmetric (see Figure 8).
We can now give the following definition.
Definition C.1.
Let $\mathcal{A}=(Q,E,s,F)$ be a GNFA without $\epsilon$-transitions. We say that $\mathcal{A}$ is Wheeler if there exists a total order $\leq$ on $Q$ such that:
  1. 1.
    For every $u,v\in Q$, if $u\leq v$, then $u\preceq_{\mathcal{A}}v$.
  2. 2.
    For every $(u^{\prime},u,\rho),(v^{\prime},v,\rho^{\prime})\in E$, if $u<v$ and $\rho^{\prime}$ is not a strict suffix of $\rho$, then $\rho\preceq\rho^{\prime}$.
  3. 3.
    For every $(u^{\prime},u,\rho),(v^{\prime},v,\rho)\in E$, if $u<v$, then $u^{\prime}\leq v^{\prime}$.
    We say that $\leq$ is a Wheeler order on $\mathcal{A}$.
Note that if $\leq$ is a Wheeler order, then $s$ comes first by Property 1, because $\epsilon\in I_{s}$ and for every $u\in Q\setminus\{s\}$ we have $\epsilon\not\in I_{u}$. As a consequence, Lemma 4.2 is true also for GNFAs without $\epsilon$-transitions (if in the statement of Lemma 4.2 we replace $\preceq_{\mathcal{A}}$ with a Wheeler order $\leq$).
If $\mathcal{A}$ is an NFA, then Definition C.1 reduces to the definition of Wheeler NFA, because by Alanko et al.’s result Property 1 follows from Properties 2, 3 and the fact that $s$ comes first in a Wheeler order. If $\mathcal{A}$ is a GDFA, then Definition C.1 reduces to the definition of Wheeler GDFA (Definition 4.1) by Lemma 4.2.
Let us present some preliminary remarks (see Figure 8). In general, a Wheeler order on a GNFA is not unique (this is already true for NFAs). Moreover, Property 2, 3 in Definition C.1 do not follow from Property 1 (this is already true for NFAs). Lastly, Property 1 does not follow from Properties 2, 3 and the requirement that $s$ must come first (while we have just seen that, if $\mathcal{A}$ is an NFA, then Property 1 follows from Properties 2, 3 and the requirement that $s$ must come first).
The problem of deciding whether a GDFA is Wheeler is tractable (see Lemma 4.3), but it becomes NP-hard on GNFAs, because it is already NP-hard on NFAs [28]. There are natural way of defining a Wheeler order on a GNFA. For example, it is easy to check that a Wheeler order on a GNFA without $\epsilon$-transitions is induced by any Wheeler order on the equivalent NFA (without $\epsilon$-transitions) defined in Section 2.
[…]
Figure 8: Left: The GNFA $\mathcal{A}$ is such that $\preceq_{\mathcal{A}}$ is not antisymmetric and both total orders in which $u_{1}$ comes first are Wheeler (in particular, a Wheeler order need not be unique). Center: The total order $\leq$ given by $u_{1}<u_{2}<u_{3}<u_{4}<u_{5}$ is such that $\leq$ satisfies Properties 1 in Definition C.1, but it does not satisfy Properties 2, 3. Right: The total order $\leq$ given by $u_{1}<u_{2}<u_{3}<u_{4}$ is such that $\leq$ satisfies Properties 2, 3 in Definition C.1 and the initial state comes first, but it does not satisfy Properties 1.
We will now explain how our results can be generalized from Wheeler GDFAs to Wheeler GNFAs without $\epsilon$-transitions. Lemma 4.6 is still true if we replace $\preceq_{\mathcal{A}}$ with any total order $\leq$ on the GNFA. This follows from how $\preceq_{\mathcal{A}}$ is defined on GNFAs. For example, let us prove Property 3 in Lemma 4.6. We know that $u,v\in Q$ are such that $u<v$ and $v\in G^{\prec}(\alpha)$, and we must prove that $u\in G^{\prec}(\alpha)$. Since $u<v$ and $\leq$ is a Wheeler order, then $u\preceq_{\mathcal{A}}v$. Let $\beta\in I_{u}$; we must prove that $\beta\prec\alpha$. If $\beta\in I_{v}$, then from $v\in G^{\prec}(\alpha)$ we immediately conclude $\beta\prec\alpha$. Now assume that $\beta\not\in I_{v}$, and pick any $\gamma\in I_{v}$. From $u\preceq_{\mathcal{A}}v$, $\beta\in I_{u}\setminus I_{v}$ and $\gamma\in I_{v}$ we obtain $\beta\prec\gamma$. From $\gamma\in I_{v}$ and $v\in G^{\prec}(\alpha)$ we obtain $\gamma\prec\alpha$, so we conclude $\beta\prec\alpha$ and we are done. Lastly, one can readily check that the proofs of all remaining results in Section 4 still hold true (if everywhere we replace $\preceq_{\mathcal{A}}$ with a Wheeler order $\leq$), because we have shown that Lemma 4.2 is also true for GNFAs without $\epsilon$-transitions.