### v3

# A Myhill-Nerode Theorem for Generalized Automata, with Applications to Pattern Matching and Compression

Abstract.
The model of generalized automata, introduced by Eilenberg in 1974, allows representing a regular language more concisely than conventional automata by allowing edges to be labeled not only with characters, but also strings. Giammaresi and Montalbano introduced a notion of determinism for generalized automata [STACS 1995]. While generalized deterministic automata retain many properties of conventional deterministic automata, the uniqueness of a minimal generalized deterministic automaton is lost.
In the first part of the paper, we show that the lack of uniqueness can be explained by introducing a set $\mathcal{W(A)}$ associated with a generalized automaton $\mathcal{A}$. If $\mathcal{A}$ is a conventional automaton, the set $\mathcal{W(A)}$ is always trivially equal to the set of all prefixes of the language recognized by the automaton, but this need not be true for generalized automata. By fixing $\mathcal{W(A)}$, we can derive for the first time a full Myhill-Nerode theorem for generalized automata, which contains the textbook Myhill-Nerode theorem for conventional automata as a degenerate case.
In the second part of the paper, we show that the set $\mathcal{W(A)}$ leads to applications for pattern matching and data compression. Wheeler automata [TCS 2017, SODA 2020] are a popular class of automata that can be compactly stored using $e\log\sigma(1+o(1))+O(e)$ bits ($e$ being the number of edges, $\sigma$ being the size of the alphabet) in such a way that pattern matching queries can be solved in $O(m\log\log\sigma)$ time ($m$ being the length of the pattern). In the paper, we show how to extend these results to generalized automata. More precisely, a Wheeler generalized automata can be stored using $\mathfrak{e}\log\sigma(1+o(1))+O(e)$ bits so that pattern matching queries can be solved in $O(m\log\log\sigma)$ time, where $\mathfrak{e}$ is the total length of all edge labels.
Key words and phrases:
Generalized automata, Myhill-Nerode theorem, minimal automaton, Burrows-Wheeler transform, FM-index.
1. Introduction
The class of regular languages can be defined starting from non-deterministic finite automata (NFAs). In his monumental work [Eil74] on automata theory (which dates back to 1974), Eilenberg proposed a natural generalization of NFAs where edges can be labeled not only with characters but with (possibly empty) finite strings, the so-called generalized non-deterministic finite automata (GNFAs). While classical automata are only a special case of generalized automata, it is immediate to realize that generalized automata can only recognize regular languages, because it is well-known that epsilon transitions do not add expressive power [HMU06], and a string-labeled edge can be decomposed into a path of edges labeled only with characters. However, generalized automata can represent regular languages more concisely than classical automata. A standard measure of the complexity of a regular language is the minimum number of states of some automaton recognizing the language, and generalized automata may have fewer states than conventional automata. In generalized automata, we assume that both the number of states and the number of edges are finite, but the number of edges cannot be bounded by some function of the number of states and the size of the finite alphabet (and so edge labels may be arbitrarily long). As a consequence, in principle it is not clear whether the problem of determining the minimum number of states of some generalized automaton recognizing a given language is decidable. In [Has91], Hashiguchi showed that the problem is decidable by proving that there must exist a state-minimal generalized automaton for which the lengths of edge labels can be bounded by a function that only depends on the size of the syntactic monoid recognizing the language.
An NFA is a deterministic finite automaton (DFA) if no state has two distinct outgoing edges with the same label. This local notion of determinism extends to global determinism, that is, given a string $\alpha$, there exists at most one path labeled $\alpha$ that can be followed starting from the initial state. However, this is not true for generalized automata (see Figure 1). When considering generalized automata, we must add the additional requirement that no state has two distinct outgoing edges such that one edge label is a prefix of the other edge label. By adding this requirement, we retrieve global determinism, thus obtaining generalized deterministic finite automata (GDFAs).
11start223344ababccaabcbc
Figure 1. No state has two distinct outgoing edges with the same label, but there are two distinct paths labeled abcabc from the initial state.
For every regular language, there exists a unique deterministic automaton recognizing the language and having the minimum number of states among all deterministic automata recognizing the language, the minimal DFA of the language. More generally, a classical textbook result in automata theory is the Myhill-Nerode theorem. Let $\operatorname{Pref}(\mathcal{L})$ be the set of all strings prefixing at least one string in the language $\mathcal{L}$. We have the following result.
Theorem 1 (Myhill-Nerode theorem).
Let $\mathcal{L}\subseteq\Sigma^{*}$. The following are equivalent:
  1. (1)
    $\mathcal{L}$ is recognized by an NFA.
  2. (2)
    The Myhill-Nerode equivalence $\equiv_{\mathcal{L}}$ has finite index.
  3. (3)
    There exists a right-invariant equivalence relation $\sim$ on $\operatorname{Pref}(\mathcal{L})$ of finite index such that $\mathcal{L}$ is the union of some $\sim$-classes.
  4. (4)
    $\mathcal{L}$ is recognized by a DFA.
Moreover, if one of the above statements is true (and so all the above statements are true), then there exists a unique minimal DFA recognizing $\mathcal{L}$ (that is, two DFAs recognizing $\mathcal{L}$ having the minimum number of states among all DFAs recognizing $\mathcal{L}$ must be isomorphic).
The problem of studying the notion of determinism in the setting of generalized automata was approached by Giammaresi and Montalbano [GM99, GM95]. The notion of isomorphism can be naturally extended to GDFAs (intuitively, two GDFAs are isomorphic if they are the same GDFA up to renaming the states), and the natural question is whether one can analogously define the minimal GDFA of a regular language. This is not possible: in general, there can exist two or more non-isomorphic state-minimal GDFAs recognizing a given regular language. Consider the two distinct GDFAs in Figure 2. It is immediate to check that the two (non-isomorphic) GDFAs recognize the same language, and it can be shown that no GDFA with fewer than three states can recognize the same language [GM99].
[…]
Figure 2. Two state-minimal GDFAs recognizing the same regular language.
The non-uniqueness of a state-minimal GDFAs seems to imply a major difference in the behavior of generalized automata compared to conventional automata, so it looks like there is no hope to derive a structural result like the Myhill-Nerode theorem in the model of generalized automata. It is natural to wonder whether the lack of uniqueness should be interpreted as a weakness of the model of generalized automata, or rather as a consequence of some deeper property. Consider a conventional automaton $\mathcal{A}$ that recognizes a language $\mathcal{L}$. As is typical in automata theory, we can assume that all states are reachable from the initial state, and all states are either final or allow reaching a final state. Then, the set $\mathcal{W(A)}$ of all strings that can be read starting from the initial state and reaching some state is exactly equal to $\operatorname{Pref}(\mathcal{L})$. However, this is no longer true in the model of generalized automata: typically, we do not have $\mathcal{W(A)}=\operatorname{Pref}(\mathcal{L})$, but only $\mathcal{W(A)}\subseteq\operatorname{Pref}(\mathcal{L})$. For example, consider Figure 2. In both automata we have $a^{3}\in\operatorname{Pref}(\mathcal{L})$, but we have $a^{3}\in\mathcal{W(A)}$ only for the GDFA on the left.
Given $\mathcal{W}\subseteq\operatorname{Pref}(\mathcal{L})$, we say that a GNFA $\mathcal{A}$ recognizing $\mathcal{L}$ is a $\mathcal{W}$-GNFA if $\mathcal{W(A)}=\mathcal{W}$. We will show that, if $\mathcal{L}$ is recognized by a $\mathcal{W}$-GDFA, then there exists a unique state-minimal $\mathcal{W}$-GDFA recognizing $\mathcal{L}$. In particular, our result will imply the uniqueness of the minimal automaton for standard DFAs, because for DFAs it must necessarily be $\mathcal{W}=\operatorname{Pref}(\mathcal{L})$.
We will actually prove much more. We will show that, once we fix $\mathcal{W}$, then nondeterminism and determinism still have the same expressive power, and it is possible to derive a characterization in terms of equivalence relations. In other words, we will prove a Myhill-Nerode theorem for generalized automata. To this end, we will introduce the notion of locally bounded set (Definition 3), which we can use to prove the following result.
Theorem 2 (Myhill-Nerode theorem for generalized automata).
Let $\mathcal{L}\subseteq\Sigma^{*}$ and let $\mathcal{W}\subseteq\Sigma^{*}$ be a locally bounded set such that $\mathcal{L}\cup\{\epsilon\}\subseteq\mathcal{W}\subseteq\operatorname{Pref}(\mathcal{L})$. The following are equivalent:
  1. (1)
    $\mathcal{L}$ is recognized by a $\mathcal{W}$-GNFA.
  2. (2)
    The Myhill-Nerode equivalence $\equiv_{\mathcal{L},\mathcal{W}}$ has finite index.
  3. (3)
    There exists a right-invariant equivalence relation $\sim$ on $\mathcal{W}$ of finite index such that $\mathcal{L}$ is the union of some $\sim$-classes.
  4. (4)
    $\mathcal{L}$ is recognized by a $\mathcal{W}$-GDFA.
Moreover, if one of the above statements is true (and so all the above statements are true), then there exists a unique minimal $\mathcal{W}$-GDFA recognizing $\mathcal{L}$ (that is, two $\mathcal{W}$-GDFAs recognizing $\mathcal{L}$ having the minimum number of states among all $\mathcal{W}$-GDFAs recognizing $\mathcal{L}$ must be isomorphic).
In particular, we will show that there is no loss of generality in assuming that $\mathcal{W}\subseteq\Sigma^{*}$ is a locally bounded set such that $\mathcal{L}\cup\{\epsilon\}\subseteq\mathcal{W}\subseteq\operatorname{Pref}(\mathcal{L})$, because these are necessary conditions for the existence of a $\mathcal{W}$-GNFA. We conclude that our Myhill-Nerode theorem for GNFAs provides the first structural result for the model of generalized automata.
In the second part of the paper, we show that the set $\mathcal{W(A)}$ sheds new light on the String Matching in Labeled Graphs (SMLG) problem. The SMLG problem has a fascinating history that dates back more than thirty years. Loosely speaking, the SMLG problem can be defined as follows: given a directed graph whose nodes or edges are labeled with nonempty strings and given a pattern string, decide whether the pattern can be read by following a path on the graph and concatenating the labels. The SMLG problem is a natural generalization of the classical pattern matching problem on texts (which requires determining whether a pattern occurs in a text) because texts can be seen as graphs consisting of a single path. The pattern matching problem on text can be efficiently solved in $O(n+m)$ time ($n$ being the length of the text, $m$ being the length of the pattern) by using the Knuth-Morris-Pratt algorithm [KMP77]. The SMLG problem is more challenging, and the complexity can be affected by the specific variant of the pattern matching problem under consideration or the class of graphs to which the problem is restricted. For example, in the (approximate) variant where one allows errors in the graph, the problem becomes NP-hard [ALL00], so generally errors are only allowed in the pattern. The SMLG problem was studied extensively during the nineties [MW92, Aku93, PK95, ALL00, RM17, Nav00]; Amir et al. showed how to solve the (exact) SMLG problem on arbitrary graphs in $O(\mathfrak{e}+me)$ time [ALL00], where $e$ is the number of edges in the graph, $m$ is the length of the pattern, and $\mathfrak{e}$ is the total length of all labels in the graph. Recently, the SMLG problem has been back in the spotlight. Equi et al. [EMTG23] showed that, on arbitrary graphs, for every $\epsilon>0$ the SMLG problem cannot be solved in $O(me^{1-\epsilon})$ or $O(m^{1-\epsilon}e)$ time, unless the Orthogonal Vectors hypothesis fails. In applications (especially in bioinformatics) we often need faster algorithms, so the SMLG problem has been restricted to classes of graphs on which it can be solved more efficiently. For example, Elastic Founder graphs can be used to represent multiple sequence alignments (MSA), a central model of biological evolution, and on Elastic Founder graphs the SMLG problem can be solved in linear time under a number of assumptions which only have a limited impact on the generality of the model [ENA+23, RM22].
The pattern matching problem on texts has been revolutionized by the invention of the Burrows-Wheeler Transform [BW94] and the FM-index [FM00, FM05], which allow solving pattern matching queries efficiently on compressed text, thus establishing a new paradigm in bioinformatics (where the huge increase of genomic data requires the development of space-efficient algorithms) [SD10]. Recently, these ideas were extended to NFAs. In particular, Wheeler NFAs are a popular class of automata on which the SMLG problem can be solved in linear time, while only storing a compact representation of the Wheeler NFA [GMS17, ADPP20]. A special case of Wheeler NFAs are de Bruijn graphs [BOSS12], which are used to perform Eulerian sequence assembly [IW95, PTW01, BNA+12]. Wheeler NFAs are also of relevant theoretical interest: for example, the powerset construction applied to a Wheeler NFA leads to a linear blow-up in the number of states of the equivalent DFA, and the equivalent DFA is Wheeler [ADPP21]; on arbitrary NFAs, the blow-up can be exponential.
The missing step is to determine whether it is possible to generalize the Burrows-Wheeler Transform and the FM-index to GNFAs, so that the resulting data structures can also be applied to Elastic Founder graphs and other classes of graphs where labels can be arbitrary strings. Indeed, in data compression it is common to consider edge-labeled graphs where one compresses unary paths in the graph to save space, and each path is replaced by a single edge labeled with the concatenation of all labels. For example, some common data structures that are stored using this mechanism are Patricia trees, suffix trees and pangenomes [Nav16, BBB+22, MBCT23]. We say that a GNFA is an $r$-GNFA if all edge labels have length at most $r$ (so a GNFA is a conventional NFA if an only if it is a $1$-GNFA). Let $m$, $e$ and $\mathfrak{e}$ as above and let $\sigma=|\Sigma|$. In Section 4, we will extend the notion of Wheelerness to GNFAs. The key ingredient will be the same set $\mathcal{W(A)}$ that we use in our Myhill-Nerode theorem: we will consider a partial order $\preceq_{\mathcal{A}}$ which sorts the set of all states with respect to the co-lexicographic order of the strings in $\mathcal{W(A)}$ (See Definition 4.2). We will then prove the following result.
Theorem 3 (FM-index of generalized automata).
Let $\mathcal{A}$ be a Wheeler $r$-GNFA, with $\sigma\leq\mathfrak{e}^{O(1)}$ and $r=O(1)$. Then, we can encode $\mathcal{A}$ by using $\mathfrak{e}\log\sigma(1+o(1))+O(e)$ bits so that later on, given a pattern $\alpha\in\Sigma^{*}$ of length $m$, we can solve the SMLG problem on $\mathcal{A}$ in $O(m\log\log\sigma)$ time. Within the same time bound, we can also decide whether $\alpha$ is recognized by $\mathcal{A}$.
If $r=1$ (that is, if $\mathcal{A}$ is a conventional Wheeler NFA), we conclude that we can encode $\mathcal{A}$ by using $e\log\sigma(1+o(1))+O(e)$ bits in such a way that we can solve the SMLG problem in $O(m\log\log\sigma)$ time; in other words, we retrieve the time and space bounds already known for Wheeler automata [GMS17, CDPP23]. If $r=O(1)$, we can still solve pattern matching queries in linear time (for constant alphabets), thus breaking the lower bound by Equi et al. while only storing a compressed representation of $\mathcal{A}$.
2. Preliminaries
Let $\Sigma$ be a finite alphabet, and let $\Sigma^{*}$ the set of all finite strings on $\Sigma$. We denote by $\epsilon$ the empty string and by $\Sigma^{+}$ the set $\Sigma^{*}\setminus\{\epsilon\}$ of all nonempty finite strings on $\Sigma$. If $\mathcal{L}\subseteq\Sigma^{*}$, let $\operatorname{Pref}(\mathcal{L})$ be the set of all prefixes of some string in $\mathcal{L}$. Note that if $\mathcal{L}\not$$\emptyset$, then $\epsilon\in\operatorname{Pref}(\mathcal{L)}$. We say that $\mathcal{L}\subseteq\Sigma^{*}$ is prefix-free if no string in $\mathcal{L}$ is a strict prefix of another string in $\mathcal{L}$. Note that if $\mathcal{L}$ is prefix-free and $\epsilon\in\mathcal{L}$, then $\mathcal{L}=\{\epsilon\}$. If $\mathcal{L}\subseteq\Sigma^{*}$, the prefix-free kernel of $\mathcal{L}$ is the set $\mathcal{K}(\mathcal{L})$ of all strings in $\mathcal{L}$ whose strict prefixes are all not in $\mathcal{L}$. Note that $\mathcal{K}(\mathcal{L})$ is always prefix-free, and $\mathcal{L}$ is prefix-free if and only if $\mathcal{L}=\mathcal{K}(\mathcal{L})$.
Let us recall the definition of generalized automaton [GM95, GM99].
{defi}
A generalized non-deterministic finite automaton (GNFA) is a 4-tuple $\mathcal{A}=(Q,E,s,F)$, where $Q$ is a finite set of states, $E\subseteq Q\times Q\times\Sigma^{*}$ is a finite set of string-labeled edges, $s\in Q$ is the initial state and $F\subseteq Q$ is a set of final states. Moreover, we assume that, for every $u\in Q$, (i) $u$ is reachable from the initial state and (ii) $u$ is co-reachable, that is, $u$ is either final or allows reaching a final state.
A generalized deterministic finite automaton (GDFA) is a GNFA such that, for every $u\in Q$, (i) no edge leaving $u$ is labeled with $\epsilon$, (ii) distinct edges leaving $u$ have distinct labels, and (iii) the set of all strings labeling some edge leaving $u$ is prefix-free.
The assumption that every state is reachable and co-reachable is standard in automata theory because all states that do not satisfy this requirement can be removed without changing the recognized language. A conventional NFA (DFA, respectively) is a GNFA (GDFA, respectively) where all edges are labeled with characters from $\Sigma$. Note that we explicitly require a GNFA to have finitely many edges (in conventional NFAs, the finiteness of the number of states automatically implies the finiteness of the number of edges because the alphabet is finite). If we allowed a GNFA to have infinitely many edges, then any nonempty (possibly non-regular) language would be recognized by a GNFA with two states, where the first state is initial, the second state is final, all edges go from the first state to the second state, and a string labels an edge if and only if it is in the language. By requiring a GNFA to have finitely many edges, the class of recognized languages is exactly the class of regular languages, because it is easy to transform a GNFA into a NFA with $\epsilon$-transitions (that is, an NFA where edges can also be labeled with the empty string $\epsilon$) that recognizes the same language by proceeding as follows: for every edge $(u^{\prime},u,\rho)\in E$, with $\rho=r_{1},\dots,r_{|\rho|}\in\Sigma^{+}$, where $r_{1},\dots,r_{|\rho|}\in\Sigma$ and $|\rho|\geq 2$, we delete the edge $(u^{\prime},u,\rho)$, we add $|\rho|-1$ new states $z_{1},\dots,z_{|\rho|-1}$, and then we add the edges $(u^{\prime},z_{1},r_{1})$, $(z_{1},z_{2},r_{2})$, $\dots$, $(z_{|\rho|-1},u,r_{|\rho|})$ (none of the new states is made initial or final).
Let us introduce some notation that will be helpful in the paper.
{defi}
Let $\mathcal{A}=(Q,E,s,F)$ be a GNFA.
  • For every $\alpha\in\Sigma^{*}$, let $I_{\alpha}$ be the set of all states that can be reached from the initial state by following edges whose labels, when concatenated, yield $\alpha$. In other words, for every $u\in Q$ we have $u\in I_{\alpha}$ if and only if there exist $t\geq 0$, $u_{1},u_{2},\dots,u_{t}\in Q$ and $\alpha_{1},\alpha_{2},\dots,\alpha_{t}\in\Sigma^{*}$ such that (i) $(s,u_{1},\alpha_{1}),(u_{1},u_{2},\alpha_{2}),(u_{2},u_{3},\alpha_{3}),\dots,(u_{t-1},u_{t},\alpha_{t})\in E$, (ii) $\alpha=\alpha_{1}\alpha_{2}\alpha_{3}\dots\alpha_{t}$ and $u_{t}=u$. Note that $\epsilon\in I_{s}$.
  • Let $\mathcal{L(A)}$ be the language recognized by $\mathcal{A}$, that is, $\mathcal{L(A)}=\{\alpha\in\Sigma^{*}\;|\;I_{\alpha}\cap F\not$$\emptyset\}$.
  • For every $u\in Q$, let $I_{u}$ be the set of all strings that can be read from the initial state to $u$ by concatenating edge labels, that is, $I_{u}=\{\alpha\in\Sigma^{*}\;|\;u\in I_{\alpha}\}$. Note that for every $u\in Q$ we have $\emptyset\subsetneqq I_{u}\subseteq\operatorname{Pref}(\mathcal{L(A)})$ because every state is reachable and co-reachable.
When $\mathcal{A}$ is not clear from the context, we write $I_{\alpha}^{\mathcal{A}}$ and $I_{u}^{\mathcal{A}}$.
3. Generalized Automata: The Myhill-Nerode Theorem
The Myhill-Nerode theorem for conventional automata (Theorem 1) provides some algebraic properties that $\operatorname{Pref}(\mathcal{L})$ must satisfy for $\mathcal{L}\subseteq\Sigma^{*}$ to be a regular language. Intuitively, the link between the algebraic characterization and the automata characterization of regular languages is that, given an NFA $\mathcal{A}=(Q,E,s,F)$ that recognizes $\mathcal{L}$, we have $\bigcup_{u\in Q}I_{u}=\operatorname{Pref}(\mathcal{L)}$: indeed, if $\alpha\in\operatorname{Pref}(\mathcal{L})$, one can read $\alpha$ on $\mathcal{A}$ starting from the initial state. However, if $\mathcal{A}=(Q,E,s,F)$ is a GNFA that recognizes $\mathcal{L}$, then we only have $\bigcup_{u\in Q}I_{u}\subseteq\operatorname{Pref}(\mathcal{L)}$, because if $\alpha\in\operatorname{Pref}(\mathcal{L})$, then one can read $\alpha$ on $\mathcal{A}$ starting from the initial state, but it may happen that we only read a strict prefix of the label of the last edge, if the label is a string of length at least two.
Let us give the following definition.
{defi}
Let $\mathcal{A}=(Q,E,s,F)$ be a GNFA. Define:
$\mathcal{W(A)}=\bigcup_{u\in Q}I_{u}.$
We say that $\mathcal{A}$ is a $\mathcal{W(A)}$-GNFA.
Note that for every $\alpha\in\Sigma^{*}$ we have $I_{\alpha}\not$$\emptyset$ if and only if $\alpha\in\mathcal{W(A)}$. Moreover, $\mathcal{L(A)}\cup\{\epsilon\}\subseteq\mathcal{W(A)}\subseteq\operatorname{Pref}(\mathcal{L(A)})$, because (i) if $\alpha\in\mathcal{L(A)}$, then $I_{\alpha}\cap F\not$$\emptyset$ and in particular $\alpha\in\mathcal{W(A)}$, (ii) $\epsilon\in I_{s}$, (iii) $I_{u}\subseteq\operatorname{Pref}(\mathcal{L(A)})$ for every $u\in Q$. Let us prove an additional property of $\mathcal{W(A)}$. Pick $\alpha\in\mathcal{W(A)}$, and consider the set $T_{\alpha}=\{\rho\in\Sigma^{+}\;|\;\alpha\rho\in\mathcal{W(A)}\}$. Consider the prefix-free kernel $\mathcal{K}(T_{\alpha})$. If $\rho\in\mathcal{K}(T_{\alpha})$, then $I_{\alpha\rho}\not$$\emptyset$, but for every $\rho^{\prime}\in\Sigma^{+}$ being a strict nonempty prefix of $\rho$ we have $I_{\alpha\rho^{\prime}}=\emptyset$. This implies that $|\rho|\leq r$, where $r$ is the maximum of the lengths of all edge labels. We conclude that $\mathcal{K}(T_{\alpha})$ must be finite because $\Sigma$ is finite. This leads to the following definition.
{defi}
  • Let $\mathcal{W}\subseteq\Sigma^{*}$. We say that $\mathcal{W}$ is locally bounded if for every $\alpha\in\mathcal{W}$ we have that $\mathcal{K}(T_{\alpha})$ is finite, where $T_{\alpha}=\{\rho\in\Sigma^{+}\;|\;\alpha\rho\in\mathcal{W}\}$.
  • Let $\mathcal{A}=(Q,E,s,F)$ be a GNFA, and let $\mathcal{W}\subseteq\Sigma^{*}$ be a locally bounded set such that $\mathcal{L(A)}\cup\{\epsilon\}\subseteq\mathcal{W}\subseteq\operatorname{Pref}(\mathcal{L(A)})$. We say that $\mathcal{A}$ is a $\mathcal{W}$-GNFA if $\mathcal{W(A)}=\mathcal{W}$.
Remark 4.
Let $\mathcal{A}=(Q,E,s,F)$ be a GDFA. Let $\alpha\in\mathcal{W(A)}$. If $\alpha=\epsilon$, then $I_{\epsilon}=\{s\}$ because no edge is labeled with $\epsilon$. If $|\alpha|>1$, then there exists (i) a prefix $\alpha_{1}\in\Sigma^{*}$ of $\alpha$ and (ii) $u_{1}\in Q$ such that $(s,u_{1},\alpha_{1})\in E$, and since $\mathcal{A}$ is a GDFA, we have (i) $\alpha_{1}\in\Sigma^{+}$ and (ii) both $\alpha_{1}$ and $u_{1}$ are unique. In particular, $\alpha_{1}\in\mathcal{W(A)}$. If $\alpha_{1}$ is a strict prefix of $\alpha$, we can repeat the argument starting from $u_{1}$. We conclude that for every $\alpha\in\mathcal{W(A)}$ we have $|I_{\alpha}|=1$. As a consequence, if $u,v\in Q$ are distinct, then $I_{u}\cap I_{v}=\emptyset$. In the following, if $\mathcal{A}$ is a GDFA and $\alpha\in\mathcal{W(A)}$, we will identify $I_{\alpha}$ and the state being its unique element.
Moreover, our argument shows that, if $\mathcal{A}$ is a GDFA, then, for every $\alpha\in\mathcal{W(A)}$ such that $|\alpha|>0$, the longest strict prefix of $\alpha$ in $\mathcal{W(A)}$ is the unique strict prefix $\alpha^{\prime}$ of $\alpha$ in $\mathcal{W(A)}$ such that, letting $\rho\in\Sigma^{+}$ be the string for which $\alpha=\alpha^{\prime}\rho$, we have $(I_{\alpha^{\prime}},I_{\alpha},\rho)\in E$. This implies that if $\mathcal{A}$ is a GDFA and $\alpha\in\mathcal{W(A)}$, then $\mathcal{K}(T_{\alpha})=\{\rho\in\Sigma^{+}\;|\;\alpha\rho\in\mathcal{W(A)},(I_{\alpha},I_{\alpha\rho},\rho)\in E\}$.
In the classical Myhill-Nerode theorem, we consider equivalence relations defined on $\operatorname{Pref}(\mathcal{L})$. In our setting, we will need to define equivalence relations on subsets of $\operatorname{Pref}(\mathcal{L})$. This leads to the following general definition.
{defi}
Let $\mathcal{W}\subseteq\Sigma^{*}$ and let $\sim$ be an equivalence relation on $\mathcal{W}$. We say that $\sim$ is right-invariant if:
$(\forall\alpha,\beta\in\mathcal{W})(\forall\phi\in\Sigma^{*})(\alpha\sim\beta\Longrightarrow((\alpha\phi\in\mathcal{W}\Longleftrightarrow\beta\phi\in\mathcal{W})\land(\alpha\phi\in\mathcal{W}\Longrightarrow\alpha\phi\sim\beta\phi))).$
Remark 5.
Notice that the property defining a right-invariant equivalence relation is trivially true if $\phi$ is the empty string, so it can be rephrased as follows:
$(\forall\alpha,\beta\in\mathcal{W})(\forall\phi\in\Sigma^{+})(\alpha\sim\beta\Longrightarrow((\phi\in T_{\alpha}\Longleftrightarrow\phi\in T_{\beta})\land(\phi\in T_{\alpha}\Longrightarrow\alpha\phi\sim\beta\phi))).$
Let us prove that $\sim$ is right-invariant if and only if:
$\begin{split}&(\forall\alpha,\beta\in\mathcal{W})(\forall\phi\in\Sigma^{+})\\ &(\alpha\sim\beta\Longrightarrow((\phi\in\mathcal{K}(T_{\alpha})\Longleftrightarrow\phi\in\mathcal{K}(T_{\beta}))\land(\phi\in\mathcal{K}(T_{\alpha})\Longrightarrow\alpha\phi\sim\beta\phi))).\end{split}$
($\Longrightarrow$) Let $\alpha,\beta\in\mathcal{W}$ such that $\alpha\sim\beta$, and let $\phi\in\mathcal{K}(T_{\alpha})$. We must prove that $\phi\in\mathcal{K}(T_{\beta})$ and $\alpha\phi\sim\beta\phi$. Since $\phi\in\mathcal{K}(T_{\alpha})\subseteq T_{\alpha}$, we immediately obtain $\phi\in T_{\beta}$ and $\alpha\phi\sim\beta\phi$, so we only have to prove that $\phi\in\mathcal{K}(T_{\beta})$. Since for every $\phi^{\prime}\in\Sigma^{+}$ we have $\phi^{\prime}\in T_{\alpha}$ if and only if $\phi^{\prime}\in T_{\beta}$, then $T_{\alpha}=T_{\beta}$, and so $\mathcal{K}(T_{\alpha})=\mathcal{K}(T_{\beta})$. As a consequence, from $\phi\in\mathcal{K}(T_{\alpha})$ we conclude $\phi\in\mathcal{K}(T_{\beta})$.
[…]
In general, an equivalence relation is not right-invariant. Let us show how to define a canonical right-invariant equivalence relation starting from any equivalence relation.
Lemma 6.
Let $\mathcal{W}\subseteq\Sigma^{*}$ and let $\sim$ be an equivalence relation on $\mathcal{W}$. For every $\alpha,\beta\in\mathcal{W}$, let:
$\alpha\sim_{r}\beta\Longleftrightarrow(\forall\phi\in\Sigma^{*})((\alpha\phi\in\mathcal{W}\Longleftrightarrow\beta\phi\in\mathcal{W})\land(\alpha\phi\in\mathcal{W}\Longrightarrow\alpha\phi\sim\beta\phi)).$
Then $\sim_{r}$ is an equivalence relation on $\mathcal{W}$, it is right-invariant and it is the coarsest right-invariant equivalence relation on $\mathcal{W}$ refining $\sim$. We say that $\sim_{r}$ is the right-invariant refinement of $\sim$.
Proof 3.1.
It is immediate to check that $\sim_{r}$ is an equivalence relation. Let us prove that $\sim_{r}$ is right-invariant. Assume that $\alpha,\beta\in\mathcal{W}$ and $\phi\in\Sigma^{*}$ are such that $\alpha\sim_{r}\beta$ and $\alpha\phi\in\mathcal{W}$. We must prove that $\beta\phi\in\mathcal{W}$ and $\alpha\phi\sim_{r}\beta\phi$. Now, $\beta\phi\in\mathcal{W}$ follows immediately from $\alpha\sim_{r}\beta$ and $\alpha\phi\in\mathcal{W}$. Moreover, $\alpha\phi\sim\beta\phi$ also follows from $\alpha\sim_{r}\beta$ and $\alpha\phi\in\mathcal{W}$. Next, fix $\psi\in\Sigma^{*}$ such that $\alpha\phi\psi\in\mathcal{W}$. We must prove that $\beta\phi\psi\in\mathcal{W}$ and $\alpha\phi\psi\sim\beta\phi\psi$. This follows from $\alpha\sim_{r}\beta$ and $\alpha\phi\psi\in\mathcal{W}$. Moreover, $\sim_{r}$ refines $\sim$ because for every $\alpha,\beta\in\mathcal{W}$, if $\alpha\sim_{r}\beta$, by letting $\phi$ be the empty string we conclude $\alpha\sim\beta$. Lastly, we want to prove that $\sim_{r}$ is the coarsest right-invariant equivalence relation of $\mathcal{W}$ refining $\sim$. Let $\sim_{*}$ be a right-invariant equivalence relation on $\mathcal{W}$ refining $\sim$. Assume that for some $\alpha,\beta\in\mathcal{W}$ we have $\alpha\sim_{*}\beta$. We must prove that $\alpha\sim_{r}\beta$. Let $\phi\in\Sigma^{*}$ be such that $\alpha\phi\in\mathcal{W}$. We must prove that $\beta\phi\in\mathcal{W}$ and $\alpha\phi\sim\beta\phi$. Since $\sim_{*}$ is right-invariant, from $\alpha\sim_{*}\beta$ and $\alpha\phi\in\mathcal{W}$ we obtain $\beta\phi\in\mathcal{W}$ and $\alpha\phi\sim_{*}\beta\phi$, which implies $\alpha\phi\sim\beta\phi$ because $\sim_{*}$ refines $\sim$.
The Myhill-Nerode equivalence plays a major role in the classical Myhill-Nerode theorem. Let us show how we can extend it when $\mathcal{W}$ is not necessarily equal to $\operatorname{Pref}(\mathcal{L})$.
{defi}
Let $\mathcal{L},\mathcal{W}\subseteq\Sigma^{*}$. The Myhill-Nerode equivalence on $\mathcal{L}$ and $\mathcal{W}$ is the equivalence relation $\equiv_{\mathcal{L},\mathcal{W}}$ on $\mathcal{W}$ defined as the right-invariant refinement of $\sim_{\mathcal{L},\mathcal{W}}$, where $\sim_{\mathcal{L},\mathcal{W}}$ is the equivalence relation on $\mathcal{W}$ such that for every $\alpha,\beta\in\mathcal{W}$:
$\alpha\sim_{\mathcal{L},\mathcal{W}}\beta\Longleftrightarrow(\alpha\in\mathcal{L}\Longleftrightarrow\beta\in\mathcal{L}).$
If $\mathcal{W}=\operatorname{Pref}(\mathcal{L})$, then we retrieve the classical Myhill-Nerode equivalence relation for $\mathcal{L}$. Let us describe some elementary properties of $\equiv_{\mathcal{L},\mathcal{W}}$.
Lemma 7.
Let $\mathcal{L},\mathcal{W}\subseteq\Sigma^{*}$. Then $\equiv_{\mathcal{L},\mathcal{W}}$ is right-invariant, and $\mathcal{L}$ is the union of some $\equiv_{\mathcal{L},\mathcal{W}}$-classes.
Proof 3.2.
First, $\equiv_{\mathcal{L},\mathcal{W}}$ is right-invariant because it is a right-invariant refinement by definition. Moreover, $\mathcal{L}$ is the union of some $\sim_{\mathcal{L},\mathcal{W}}$-classes, and so also of some $\equiv_{\mathcal{L},\mathcal{W}}$-classes because $\equiv_{\mathcal{L},\mathcal{W}}$ refines $\sim_{\mathcal{L},\mathcal{W}}$.
Let $\mathcal{A}$ be a conventional NFA, and define the equivalence relation $\sim_{\mathcal{A}}$ on $\operatorname{Pref}(\mathcal{L(A)})$ as follows: for every $\alpha,\beta\in\operatorname{Pref}(\mathcal{L(A)})$, let $\alpha\sim_{\mathcal{A}}\beta$ if and only if $I_{\alpha}=I_{\beta}$. This equivalence relation is an intermediate tool in the Myhill-Nerode theorem for conventional automata, and it can be also defined for a generalized automata $\mathcal{A}$ by considering the equivalence relation $\sim_{\mathcal{A}}$ on $\mathcal{W(A)}$ such that for every $\alpha,\beta\in\mathcal{W(A)}$ we have $\alpha\sim_{\mathcal{A}}\beta$ if and only if $I_{\alpha}=I_{\beta}$. If $\mathcal{A}$ is an NFA (or an NFA with $\epsilon$-transitions), then $\sim_{\mathcal{A}}$ is right-invariant, because for every $\alpha\in\operatorname{Pref}(\mathcal{L(A)})$ and for every prefix $\alpha^{\prime}$ of $\alpha$, any path from the initial state to a state in $I_{\alpha}$ must go through a state in $I_{\alpha^{\prime}}$. However, in general this property is not true if $\mathcal{A}$ is a GNFA, so $\sim_{\mathcal{A}}$ need not be right-invariant if $\mathcal{A}$ is a GNFA (see Figure 3). Since right-invariance is crucial in the Myhill-Nerode theorem, we will consider the right-invariant refinement of $\sim_{\mathcal{A}}$.
11start223344a,ba,bccacac
Figure 3. A GNFA $\mathcal{A}$ such that $\sim_{\mathcal{A}}$ is not right-invariant. Indeed, we have $a,b,ac,bc\in\mathcal{W(A)}$ and $a\sim_{\mathcal{A}}b$, but $ac\not\sim_{\mathcal{A}}bc$.
{defi}
Let $\mathcal{A}=(Q,E,s,F)$ be a GNFA. Let $\equiv_{\mathcal{A}}$ be the right-invariant refinement of $\sim_{\mathcal{A}}$, where $\sim_{\mathcal{A}}$ is the equivalence relation on $\mathcal{W(A)}$ such that for every $\alpha,\beta\in\mathcal{W(A)}$:
$\alpha\sim_{\mathcal{A}}\beta\Longleftrightarrow I_{\alpha}=I_{\beta}.$
Remark 8.
Let us prove that if $\mathcal{A}$ is GDFA, then $\sim_{\mathcal{A}}$ is right-invariant. Let $\alpha,\beta\in\mathcal{W}$ be such that $I_{\alpha}=I_{\beta}$, and let $\phi\in\mathcal{K}(T_{\alpha})$. By Remark 5, we only have to prove that $\phi\in\mathcal{K}(T_{\beta})$ and $I_{\alpha\phi}=I_{\beta\phi}$. By Remark 4, we have $\alpha\phi\in\mathcal{W(A)}$ and $(I_{\alpha},I_{\alpha\phi},\phi)\in E$. Hence $(I_{\beta},I_{\alpha\phi},\phi)\in E$, we obtain $I_{\alpha\phi}=I_{\beta\phi}$, and again by Remark 4 we conclude $\phi\in\mathcal{K}(T_{\beta})$. Notice that, in fact, the generalized automaton $\mathcal{A}$ in Figure 3 is not a GDFA.
Since $\equiv_{\mathcal{A}}$ is the right-invariant refinement of $\sim_{\mathcal{A}}$, we conclude that, if $\mathcal{A}$ is a GDFA, then $\equiv_{\mathcal{A}}$ and $\sim_{\mathcal{A}}$ are the same equivalence relation.
Let us study the properties of $\equiv_{\mathcal{A}}$.
Lemma 9.
Let $\mathcal{A}=(Q,E,s,F)$ be a GNFA. Then, $\equiv_{\mathcal{A}}$ is right-invariant, it refines $\equiv_{\mathcal{L(A)},\mathcal{W(A)}}$, it has finite index, and $\mathcal{L(A)}$ is the union of some $\equiv_{\mathcal{A}}$-classes.
Proof 3.3.
First, $\equiv_{\mathcal{A}}$ is right-invariant because it is a right-invariant refinement by definition.
Let us prove that $\mathcal{L(A)}$ is the union of some $\equiv_{\mathcal{A}}$-classes. It will suffice to show that $\mathcal{L(A)}$ is the union of some $\sim_{\mathcal{A}}$-classes, because $\equiv_{\mathcal{A}}$ is a refinement of $\sim_{\mathcal{A}}$. Let $\alpha,\beta\in\mathcal{W(A)}$ such that $\alpha\sim_{\mathcal{A}}\beta$, that is, $I_{\alpha}=I_{\beta}$. We must prove that $\alpha\in\mathcal{L(A)}$ if and only if $\beta\in\mathcal{L(A)}$. We have $\alpha\in\mathcal{L(A)}$ if and only if $I_{\alpha}\cap F\not$$\emptyset$, if and only if $I_{\beta}\cap F\not$$\emptyset$, if and only if $\beta\in\mathcal{L(A)}$.
Let us prove that $\equiv_{\mathcal{A}}$ refines $\equiv_{\mathcal{L(A)},\mathcal{W(A)}}$. Since $\mathcal{L(A)}$ is the union of some $\equiv_{\mathcal{A}}$-classes, then $\equiv_{\mathcal{A}}$ refines $\sim_{\mathcal{L(A)},\mathcal{W(A)}}$. Then, $\equiv_{\mathcal{A}}$ also refines $\equiv_{\mathcal{L(A)},\mathcal{W(A)}}$, because $\equiv_{\mathcal{A}}$ is right-invariant and $\equiv_{\mathcal{L(A)},\mathcal{W(A)}}$ is the coarsest right-invariant equivalence relation refining $\sim_{\mathcal{L(A)},\mathcal{W(A)}}$.
Lastly, let us prove that $\equiv_{\mathcal{A}}$ has finite index. Let $\bar{\mathcal{A}}$ be the NFA with $\epsilon$-transitions for which $\mathcal{L}(\bar{\mathcal{A}})=\mathcal{L}(\mathcal{A})$ that can be constructed from $\mathcal{A}$ by proceedings as explained in Section 2. Let $\alpha,\beta\in\mathcal{W(A)}$ be such that $\alpha\sim_{\bar{\mathcal{A}}}\beta$. Let us prove that it must be $\alpha\equiv_{\mathcal{A}}\beta$. By the definition of right-invariant refinement, we must prove that for every $\phi\in\Sigma^{*}$ we have $\alpha\phi\in\mathcal{W(A)}$ if and only if $\beta\phi\in\mathcal{W(A)}$, and, if $\alpha\phi\in\mathcal{W(A)}$, then $\alpha\phi\sim_{\mathcal{A}}\beta\phi$. Since $\bar{\mathcal{A}}$ is an NFA with $\epsilon$-transitions, we have that $\sim_{\bar{\mathcal{A}}}$ is right-invariant (see the discussion before Definition 3). From $\alpha\phi\in\mathcal{W(A)}\subseteq\operatorname{Pref}(\mathcal{L(A)})$ we obtain $\alpha\phi\sim_{\bar{\mathcal{A}}}\beta\phi$ and so $I^{\bar{\mathcal{A}}}_{\alpha\phi}=I^{\bar{\mathcal{A}}}_{\beta\phi}$. By the construction of $\bar{\mathcal{A}}$, we obtain $I_{\alpha\phi}^{\mathcal{A}}=I_{\alpha\phi}^{\bar{\mathcal{A}}}\cap Q=I_{\beta\phi}^{\bar{\mathcal{A}}}\cap Q=I_{\beta\phi}^{\mathcal{A}}$. As a consequence, we have $\alpha\phi\in\mathcal{W(A)}$ if and only if $I_{\alpha\phi}^{\mathcal{A}}\not$$\emptyset$, if and only if $I_{\beta\phi}^{\mathcal{A}}\not$$\emptyset$, if and only if $\beta\phi\in\mathcal{W(A)}$, and, if $\alpha\phi\in\mathcal{W(A)}$, then $\alpha\phi\sim_{\mathcal{A}}\beta\phi$. This proves that $\alpha\equiv_{\mathcal{A}}\beta$. Since for every $\alpha,\beta\in\mathcal{W(A)}$ we have that $\alpha\sim_{\bar{\mathcal{A}}}\beta$ implies $\alpha\equiv_{\mathcal{A}}\beta$, in order to prove that $\equiv_{\mathcal{A}}$ has finite index it will suffice to prove that $\sim_{\bar{\mathcal{A}}}$ has finite index. To this end, observe that every $\sim_{\bar{\mathcal{A}}}$-class can be associated with a distinct nonempty subset of the set $\bar{Q}$ of all states in $\bar{\mathcal{A}}$ (via the well-defined mapping $\alpha]_{\sim_{\bar{\mathcal{A}}}}\mapsto I^{\bar{\mathcal{A}}}_{\alpha})$, so the index of $\sim_{\bar{\mathcal{A}}}$ is bounded by $2^{|\bar{Q}|}-1$.
Given two GNFAs $\mathcal{A}=(Q,E,s,F)$ and $\mathcal{A}^{\prime}=(Q^{\prime},E^{\prime},s^{\prime},F)$, we say that $\mathcal{A}$ and $\mathcal{A}^{\prime}$ are isomorphic if there exists a bijection $\phi:Q\mapsto Q^{\prime}$ such that (i) for every $u,v\in Q$ and for every $\rho\in\Sigma^{*}$ we have $(u,v,\rho)\in E$ if and only if $(\phi(u),\phi(v),\rho)\in E^{\prime}$, (ii) $\phi(s)=s^{\prime}$ and (iii) for every $u\in Q$ we have $u\in F$ if and only if $\phi(u)\in F$. In particular, for two isomorphic GNFAs $\mathcal{A}$ and $\mathcal{A}^{\prime}$, we have that $\mathcal{A}$ is a GDFA if and only if $\mathcal{A}^{\prime}$ is a GDFA.
The following lemma is crucial to derive our Myhill-Nerode theorem for generalized automata.
Lemma 10.
Let $\mathcal{L}\subseteq\Sigma^{*}$ and let $\mathcal{W}\subseteq\Sigma^{*}$ be a locally bounded set such that $\mathcal{L}\cup\{\epsilon\}\subseteq\mathcal{W}\subseteq\operatorname{Pref}(\mathcal{L})$. Assume that $\mathcal{L}$ is the union of some classes of a right-invariant equivalence relation $\sim$ on $\mathcal{W}$ of finite index. Then, $\mathcal{L}$ is recognized by a $\mathcal{W}$-GDFA $\mathcal{A_{\sim}}=(Q_{\sim},E_{\sim},s_{\sim},F_{\sim})$ such that:
  1. (1)
    $|Q_{\sim}|$ is equal to the index of $\sim$.
  2. (2)
    $\equiv_{\mathcal{A_{\sim}}}$ and $\sim$ are the same equivalence relation.
Moreover, if $\mathcal{B}$ is a $\mathcal{W}$-GDFA that recognizes $\mathcal{L}$, then $\mathcal{A_{\equiv_{\mathcal{B}}}}$ is isomorphic to $\mathcal{B}$.
Proof 3.4.
Define $\mathcal{A_{\sim}}=(Q_{\sim},E_{\sim},s_{\sim},F_{\sim})$ as follows.
  • $Q_{\sim}=\{[\alpha]_{\sim}\;|\;\alpha\in\mathcal{W}\}$.
  • $s_{\sim}=[\epsilon]_{\sim}$.
  • $E_{\sim}=\{([\alpha]_{\sim},[\alpha\rho]_{\sim},\rho)\;|\;\alpha\in\mathcal{W},\rho\in\mathcal{K}(T_{\alpha})\}$.
  • $F_{\sim}=\{[\alpha]_{\sim}\;|\;\alpha\in\mathcal{L}\}$.
Let us prove that $\mathcal{A}_{\sim}$ is a well-defined GDFA. First, the number of states is finite because $\sim$ has finite index. Next, $\epsilon\in\mathcal{W}$, so $s_{\sim}$ is well defined. The set $F_{\sim}$ is well defined because (i) if $\alpha\in\mathcal{L}$, then $\alpha\in\mathcal{W}$, so $\alpha]_{\sim}$ is well defined, and (ii) $\mathcal{L}$ if the union of some $\sim$-classes, so if for some $\alpha,\alpha^{\prime}\in\mathcal{W}$ we have $\alpha\sim\alpha^{\prime}$, then $\alpha\in\mathcal{L}$ if and only if $\alpha^{\prime}\in\mathcal{L}$. Let us prove that $E_{\sim}$ is well defined. Pick $\alpha\in\mathcal{W}$ and $\rho\in\mathcal{K}(T_{\alpha})$. First, notice that $\rho\in T_{\alpha}$, so $\alpha\rho\in\mathcal{W}$ and $\alpha\rho]_{\sim}$ is well defined. Moreover, if $\alpha^{\prime}\in\mathcal{W}$ is such that $\alpha\sim\alpha^{\prime}$, then $\alpha^{\prime}\rho\in\mathcal{W}$ and $\alpha\rho\sim\alpha^{\prime}\rho$, because $\sim$ is right-invariant. Since $\alpha\sim\alpha^{\prime}$ and $\sim$ is right-invariant, we have $T_{\alpha}=T_{\alpha^{\prime}}$ and so $\mathcal{K}(T_{\alpha})=\mathcal{K}(T_{\alpha^{\prime}})$, which shows that $E_{\sim}$ is well defined. Moreover, $E_{\sim}$ is a finite set because $\mathcal{W}$ is locally bounded, so $\mathcal{K}(T_{\alpha})$ is finite for every $\alpha\in\mathcal{W}$. Let us prove by induction on $|\alpha|$ that, for every $\alpha\in\mathcal{W}$, the state $\alpha]_{\sim}$ is reachable from the initial state. If $|\alpha|=0$, then $\alpha=\epsilon$ and so $\alpha]_{\sim}$ is the initial state. Now assume $|\alpha|>0$. Let $\alpha^{\prime}$ be the longest strict prefix of $\alpha$ such that $\alpha^{\prime}\in\mathcal{W}$, which must exist because $\epsilon$ is a strict prefix of $\alpha$ (we have $|\alpha|>0$) and $\epsilon\in\mathcal{W}$. Let $\rho\in\Sigma^{+}$ be such that $\alpha=\alpha^{\prime}\rho$. Let us prove that $([\alpha^{\prime}]_{\sim},[\alpha]_{\sim},\rho)\in E_{\sim}$. Since $\alpha=\alpha^{\prime}\rho$, we only have to show that $\rho\in\mathcal{K}(T_{\alpha^{\prime}})$. We have $\rho\in T_{\alpha^{\prime}}$ because $\alpha=\alpha^{\prime}\rho\in\mathcal{W}$, and we have $\rho\in\mathcal{K}(T_{\alpha^{\prime}})$ because if $\rho^{\prime}\in\Sigma^{+}$ is a strict prefix of $\rho$, then $\rho^{\prime}\not\in T_{\alpha}$, otherwise we would have $\alpha^{\prime}\rho\in\mathcal{W}$, and $\alpha^{\prime}\rho^{\prime}$ would be a strict prefix of $\alpha$ longer than $\alpha^{\prime}$, which contradicts the maximality of $\alpha^{\prime}$. By the inductive hypothesis, $\alpha^{\prime}]_{\sim}$ is reachable from the initial state, so $\alpha]_{\sim}$ is also reachable from the initial state because $([\alpha^{\prime}]_{\sim},[\alpha]_{\sim},\rho)\in E_{\sim}$. Let us prove every state is co-reachable. Fix $\alpha\in\mathcal{W}$. We must prove that $\alpha]_{\sim}$ is either final, or it allows reaching a final state. In particular, $\alpha\in\operatorname{Pref}(\mathcal{L})$, so there exists $\rho\in\Sigma^{*}$ such that $\alpha\rho\in\mathcal{L}$, and so $\alpha\rho\in\mathcal{W}$. Let $\rho_{1},\dots,\rho_{s}$ be all distinct prefixes of $\rho$ such that $\alpha\rho_{i}\in\mathcal{W}$ for every $1\leq i\leq s$, where $\rho_{i}$ is a strict prefix of $\rho_{i+1}$ for every $1\leq i\leq s-1$. Note that $s\geq 1$, $\rho_{1}=\epsilon$ and $\rho_{s}=\rho$. The same argument used to prove that every state is reachable from the initial state shows that $([\alpha\rho_{i}]_{\sim},[\alpha\rho_{i+1}],\tau_{i})\in E_{\sim}$, where $\tau_{i}\in\Sigma^{+}$ is such that $\rho_{i+1}=\rho_{i}\tau_{i}$ for every $1\leq i\leq s-1$. Since $\alpha\rho_{s}=\alpha\rho\in\mathcal{L}$ and so $\alpha\rho_{s}]_{\sim}$ is a final state, then $\alpha\rho_{i}]_{\sim}$ is either final or it allows reaching a final state for every $1\leq i\leq s$, and the conclusion follows because $\alpha=\alpha\rho_{1}$. Lastly, for every $\alpha\in\mathcal{W}$, every $\rho\in\Sigma^{+}$ labels at most one edge leaving $\alpha]_{\sim}$ because we have shown that the definition of $E_{\sim}$ does not depend on the choice of the representatives in the equivalence classes, and the set of all strings labeling an edge leaving $\alpha]_{\sim}$ is $\mathcal{K}(T_{\alpha})$, which is prefix-free, being a prefix-free kernel. This concludes the proof that $\mathcal{A}_{\sim}$ is a well-defined GDFA.
Next, we want to prove that for every $\alpha\in\Sigma^{*}$ and for every $\beta\in\mathcal{W}$ we have:
$(\alpha\in\mathcal{W(A_{\sim})})\land(I_{\alpha}=[\beta]_{\sim})\Longleftrightarrow(\alpha\in\mathcal{W})\land(\alpha\sim\beta).$ (1)
We proceed by induction on $|\alpha|$. If $|\alpha|=0$, then $\alpha=\epsilon$. Notice that $\epsilon\in\mathcal{W(A_{\sim})}$ because $\epsilon\in I_{s_{\sim}}$, and $\epsilon\in\mathcal{W}$. Moreover, for every $\beta\in\mathcal{W}$ we have $I_{\epsilon}=[\beta]_{\sim}\Longleftrightarrow[\beta]_{\sim}=s_{\sim}\Longleftrightarrow\epsilon\sim\beta$.
[…]
Let $\alpha^{\prime}$ be the longest strict prefix of $\alpha$ such that $\alpha^{\prime}\in\mathcal{W(A_{\sim})}$, which must exist because $\epsilon$ is a strict prefix of $\alpha$, being $|\alpha|>0$, and $\epsilon\in\mathcal{W(A_{\sim})}$. Since we know that a string shorter than $\alpha$ is in $\mathcal{W(A_{\sim})}$ if and only if it is in $\mathcal{W}$, then $\alpha^{\prime}$ is also the longest strict prefix of $\alpha$ such that $\alpha^{\prime}\in\mathcal{W}$. Moreover, we know that $I_{\alpha^{\prime}}=[\alpha^{\prime}]_{\sim}$. Write $\alpha=\alpha^{\prime}\rho$, with $\rho\in\Sigma^{+}$.
$(\Longrightarrow)$ Assume that $(\alpha\in\mathcal{W(A_{\sim})})\land(I_{\alpha}=[\beta]_{\sim})$. We must prove that $(\alpha\in\mathcal{W})\land(\alpha\sim\beta)$. Since $\alpha^{\prime},\alpha\in\mathcal{W(A_{\sim})}$, $\alpha^{\prime}$ is the longest strict prefix of $\alpha$ such such that $\alpha^{\prime}\in\mathcal{W(A_{\sim})}$ and $\alpha=\alpha^{\prime}\rho$, then $\rho\in\mathcal{K}(T_{\alpha})$, so by Remark 4 we have $(I_{\alpha^{\prime}},I_{\alpha},\rho)\in E_{\sim}$. Since $I_{\alpha^{\prime}}=[\alpha^{\prime}]_{\sim}$ and $I_{\alpha}=[\beta]_{\sim}$, we have $([\alpha^{\prime}]_{\sim},[\beta]_{\sim},\rho)\in E_{\sim}$. The definition of $E_{\sim}$ implies $\alpha=\alpha^{\prime}\rho\in\mathcal{W}$ and $\alpha\sim\beta$.
[…]
This concludes the proof of Equation 1. In particular, by the same argument used at the beginning of the inductive step, we obtain $\mathcal{W(A_{\sim})}=\mathcal{W}$, which proves that $\mathcal{A_{\sim}}$ is a $\mathcal{W}$-GDFA, and by the same argument we also obtain that for every $\alpha\in\mathcal{W(A_{\sim})}=\mathcal{W}$:
$I_{\alpha}=[\alpha]_{\sim}.$ (2)
From the definition of $F_{\sim}$ and Equation 1 we obtain:
$\begin{split}\mathcal{L(A_{\sim})}&=\{\alpha\in\mathcal{W(A_{\sim})}\;|\;I_{\alpha}=[\beta]_{\sim}\text{ for some $\beta\in\mathcal{L}$}\}\\ &=\{\alpha\in\mathcal{W}\;|\;\alpha\sim\beta\text{ for some $\beta\in\mathcal{L}$}\}=\mathcal{L}\end{split}$
where in the last equality we have $(\subseteq)$ because $\mathcal{L}$ is the union of some $\sim$-classes, and $(\supseteq)$ because $\mathcal{L}\subseteq\mathcal{W}$. This proves that $\mathcal{A_{\sim}}$ recognizes $\mathcal{L}$. Moreover:
  1. (1)
    The number of states of $\mathcal{A}_{\sim}$ is equal to the index of $\sim$ by the definition of $Q_{\sim}$.
  2. (2)
    By Equation 2, for every $\alpha,\beta\in\mathcal{W}=\mathcal{W(A_{\sim})}$ we have $\alpha\equiv_{A_{\sim}}\beta\Longleftrightarrow I_{\alpha}=I_{\beta}\Longleftrightarrow[\alpha]_{\sim}=[\beta]_{\sim}\Longleftrightarrow\alpha\sim\beta$, so $\equiv_{A_{\sim}}$ and $\sim$ are the same equivalence relation.
Lastly, suppose that $\mathcal{B}=(Q_{\mathcal{B}},E_{\mathcal{B}},s_{\mathcal{B}},F_{\mathcal{B}})$ is a $\mathcal{W}$-GDFA that recognizes $\mathcal{L}$. Notice that by Lemma 9 we have that $\equiv_{\mathcal{B}}$ is right-invariant, $\equiv_{\mathcal{B}}$ has finite index, and $\mathcal{L}$ is the union of some $\equiv_{\mathcal{B}}$-classes, so $\mathcal{A_{\equiv_{\mathcal{B}}}}$ is well defined, and $\mathcal{W(\mathcal{A_{\equiv_{\mathcal{B}}}})}=\mathcal{W}=\mathcal{W(B)}$. Let $\phi:Q_{\equiv_{\mathcal{B}}}\mapsto Q_{\mathcal{B}}$ be the function sending the state $\alpha]_{\equiv_{\mathcal{B}}}$ of $Q_{\equiv_{\mathcal{B}}}$ into the state $I_{\alpha}^{\mathcal{B}}$ of $Q_{\mathcal{B}}$, that is, $\phi([\alpha]_{\equiv_{\mathcal{B}}})=I_{\alpha}^{\mathcal{B}}$ for every $\alpha\in\mathcal{W}$. Notice that $\phi$ is well defined and injective because Remark 8 implies that $\equiv_{\mathcal{B}}$ and $\sim_{\mathcal{B}}$ are the same equivalence relation, so we have $\alpha]_{\equiv_{\mathcal{B}}}=[\beta]_{\equiv_{\mathcal{B}}}\Longleftrightarrow\alpha\equiv_{\mathcal{B}}\beta\Longleftrightarrow\alpha\sim_{\mathcal{B}}\beta\Longleftrightarrow I_{\alpha}^{\mathcal{B}}=I_{\beta}^{\mathcal{B}}$ for every $\alpha,\beta\in\mathcal{W}$. Let us prove that $\phi$ determines an isomorphism between $\mathcal{A_{\equiv_{\mathcal{B}}}}$ and $\mathcal{B}$. First, $\phi$ is surjective because every state of $\mathcal{B}$ is reachable from the initial state. Next, $\phi(s_{\equiv_{\mathcal{B}}})=\phi([\epsilon]_{\equiv_{\mathcal{B}}})=I_{\epsilon}^{\mathcal{B}}=s_{\mathcal{B}}$. Moreover, for every $\alpha\in\mathcal{W}$ we have $\alpha]_{\equiv_{\mathcal{B}}}\in F_{\equiv_{\mathcal{B}}}\Longleftrightarrow\alpha\in\mathcal{L}\Longleftrightarrow I_{\alpha}^{\mathcal{B}}\in F_{\mathcal{B}}$. Finally, by Remark 4, for every $\alpha,\beta\in\mathcal{W}$ and for every $\rho\in\Sigma^{+}$:
$\begin{split}([\alpha]_{\equiv_{\mathcal{B}}},[\beta]_{\equiv_{\mathcal{B}}},\rho)\in E_{\equiv_{\mathcal{B}}}&\Longleftrightarrow\alpha\rho\in\mathcal{W}\land\rho\in\mathcal{K}(T_{\alpha})\land\alpha\rho\equiv_{\mathcal{B}}\beta\\ &\Longleftrightarrow\alpha\rho\in\mathcal{W}\land\rho\in\mathcal{K}(T_{\alpha})\land(I_{\alpha}^{\mathcal{B}},I_{\alpha\rho}^{\mathcal{B}},\rho)\in E_{\mathcal{B}}\land I_{\alpha\rho}^{\mathcal{B}}=I_{\beta}^{\mathcal{B}}\\ &\Longleftrightarrow(I_{\alpha}^{\mathcal{B}},I_{\beta}^{\mathcal{B}},\rho)\in E_{\mathcal{B}}\Longleftrightarrow(\phi([\alpha]_{\equiv_{\mathcal{B}}}),\phi([\beta]_{\equiv_{\mathcal{B}}}),\rho)\in E_{\mathcal{B}}\end{split}$
and we conclude that $\mathcal{A_{\equiv_{\mathcal{B}}}}$ and $\mathcal{B}$ are isomorphic.
Remark 11.
In the statement of Lemma 10 we cannot remove the assumption that $\mathcal{W}$ is locally bounded, because we have shown that if $\mathcal{A}$ is a GNFA, then $\mathcal{W(A)}$ is locally bounded. However, if $\mathcal{W}$ is not locally bounded, then $\mathcal{A}_{\sim}$ is still a well-defined automaton with finitely many states, but it has infinitely many edges. For example, $\mathcal{W}=\{\epsilon\}\cup a^{*}b$ is not locally bounded because (i) $T_{\epsilon}=a^{*}b$ and (ii) $\mathcal{K}(T_{\epsilon})=a^{*}b$ is an infinite set. If $\mathcal{L}=a^{*}b$, then $\mathcal{L}\cup\{\epsilon\}\subseteq\mathcal{W}\subseteq\operatorname{Pref}(\mathcal{L})$. Moreover, $\equiv_{\mathcal{L},\mathcal{W}}$ has finite index (the equivalence classes are $\{\epsilon\}$ and $a^{*}b$), and by Lemma 7 we know that $\equiv_{\mathcal{L},\mathcal{W}}$ is right-invariant and $\mathcal{L}$ is the union of some $\equiv_{\mathcal{L},\mathcal{W}}$-classes. We conclude that $\mathcal{A_{\equiv_{\mathcal{L},\mathcal{W}}}}$ is well defined, but it has infinitely many edges (see Figure 4).
[…]
Figure 4. An example where $\mathcal{W}$ is not locally bounded.
We can now prove our Myhill-Nerode theorem for generalized automata.
Theorem 12 (Myhill-Nerode theorem for generalized automata).
Let $\mathcal{L}\subseteq\Sigma^{*}$ and let $\mathcal{W}\subseteq\Sigma^{*}$ be a locally bounded set such that $\mathcal{L}\cup\{\epsilon\}\subseteq\mathcal{W}\subseteq\operatorname{Pref}(\mathcal{L})$. The following are equivalent:
  1. (1)
    $\mathcal{L}$ is recognized by a $\mathcal{W}$-GNFA.
  2. (2)
    The Myhill-Nerode equivalence $\equiv_{\mathcal{L},\mathcal{W}}$ has finite index.
  3. (3)
    There exists a right-invariant equivalence relation $\sim$ on $\mathcal{W}$ of finite index such that $\mathcal{L}$ is the union of some $\sim$-classes.
  4. (4)
    $\mathcal{L}$ is recognized by a $\mathcal{W}$-GDFA.
Moreover, if one of the above statements is true (and so all the above statements are true), then there exists a unique minimal $\mathcal{W}$-GDFA recognizing $\mathcal{L}$ (that is, two $\mathcal{W}$-GDFAs recognizing $\mathcal{L}$ having the minimum number of states among all $\mathcal{W}$-GDFAs recognizing $\mathcal{L}$ must be isomorphic). The minimal $\mathcal{W}$-GDFA recognizing $\mathcal{L}$ is $\mathcal{A_{\equiv_{\mathcal{L},\mathcal{W}}}}$ as defined in Lemma 10, where $\equiv_{\mathcal{L},\mathcal{W}}$ is the Myhill-Nerode equivalence on $\mathcal{L}$ and $\mathcal{W}$.
Proof 3.5.
$(1)\Longrightarrow(2)$ Let $\mathcal{A}$ be a $\mathcal{W}$-GDFA recognizing $\mathcal{L}$. By Lemma 9 we have that $\equiv_{\mathcal{A}}$ has finite index and it refines $\equiv_{\mathcal{L},\mathcal{W}}$, so also $\equiv_{\mathcal{L},\mathcal{W}}$ has finite index.
$(2)\Longrightarrow(3)$ By Lemma 7 the desired equivalence relation is $\equiv_{\mathcal{L},\mathcal{W}}$.
$(3)\Longrightarrow(4)$ It follows from Lemma 10.
$(4)\Longrightarrow(1)$ Every $\mathcal{W}$-GDFA is a $\mathcal{W}$-GNFA.
Now, assume that one of the above statements is true (and so all the above statements are true). Let us prove that the minimal $\mathcal{W}$-GDFA recognizing $\mathcal{L}$ is $\mathcal{A_{\equiv_{\mathcal{L},\mathcal{W}}}}$ as defined in Lemma 10. First, $\mathcal{A_{\equiv_{\mathcal{L},\mathcal{W}}}}$ is well-defined $\mathcal{W}$-GDFA recognizing $\mathcal{L}$ because (i) $\equiv_{\mathcal{L},\mathcal{W}}$ is right-invariant and $\mathcal{L}$ is the union of some $\equiv_{\mathcal{L},\mathcal{W}}$-classes by Lemma 7, and (ii) $\equiv_{\mathcal{L},\mathcal{W}}$ has finite index by one of the statements that we assume to be true. Let $\mathcal{B}$ be any $\mathcal{W}$-GDFA recognizing $\mathcal{L}$ non-isomorphic to $\mathcal{A_{\equiv_{\mathcal{L},\mathcal{W}}}}$. We must prove that the number of states on $\mathcal{A_{\equiv_{\mathcal{L},\mathcal{W}}}}$ is smaller than the number of states of $\mathcal{B}$. By Lemma 10, $\mathcal{A}_{\equiv_{\mathcal{B}}}$ is isomorphic to $\mathcal{B}$. We know that $\equiv_{\mathcal{B}}$ is a refinement of $\equiv_{\mathcal{L},\mathcal{W}}$ by Lemma 9, and it must be a strict refinement of $\equiv_{\mathcal{L},\mathcal{W}}$, otherwise $\mathcal{A_{\equiv_{\mathcal{L},\mathcal{W}}}}$ would be equal to $\mathcal{A}_{\equiv_{\mathcal{B}}}$, which is isomorphic to $\mathcal{B}$, a contradiction. We conclude that the index of $\equiv_{\mathcal{L},\mathcal{W}}$ is smaller than the index of $\equiv_{\mathcal{B}}$, so again by Lemma 10 the number of states of $\mathcal{A_{\equiv_{\mathcal{L},\mathcal{W}}}}$ is smaller than the number of states of $\mathcal{A}_{\equiv_{\mathcal{B}}}$ and so of $\mathcal{B}$.
The Myhill-Nerode theorem for conventional automata (Theorem 1) is a special case of Theorem 12, because if $\mathcal{A}$ is an NFA, then $\mathcal{W(A)}=\operatorname{Pref}(\mathcal{L(A)})$. Moreover, Theorem 12 is consistent with the example in Figure 2, because, calling $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ the two GDFAs in Figure 2, we have shown that $\mathcal{W}(\mathcal{A}_{1})\not=\mathcal{W}(\mathcal{A}_{2})$.
4. Generalized Automata: Pattern Matching and Compression
In this section, we prove Theorem 3 and some related results. In order to present the main ideas, it will suffice to consider GDFAs. The more general case of GNFAs will be considered in Appendix A.
[…]
  • $\mathtt{OUT}_{1}=001011$
  • $\mathtt{OUT}_{2}=001101$
  • $\mathtt{IN}_{1}=100101$
  • $\mathtt{IN}_{2}=101001$
  • $\mathtt{LAB}_{1}=(b,c,b)$
  • $\mathtt{LAB}_{2}=(ab,ac,bc)$
  • $\mathtt{FIN}=011$
Figure 5. Left: A Wheeler GDFA $\mathcal{A}$. The states are numbered following the total order $\preceq_{\mathcal{A}}$. Right: The Burrows-Wheeler Transform (BWT) of $\mathcal{A}$ (see Definition 4.3).
4.1. Preliminary Definitions
To introduce Wheeler generalized automata, we will first present some preliminary definitions. Following the article introduction, we can naturally define the SMLG problem for GNFAs.
{defi}
Let $\mathcal{A}$ be a GNFA. The String Matching in Labeled Graphs (SMLG) problem for GNFAs is defined as follows: build a data structure that encodes $\mathcal{A}$ such that, given a string $\alpha$, we can efficiently compute the set of all states reached by a path suffixed by $\alpha$.
Let $V$ be a set. We say that a (binary) relation $\leq$ on $V$ is a partial order if $\leq$ is reflexive, antisymmetric and transitive. A partial order $\leq$ is a total order if for every $u,v\in V$ we have $(u\leq v)\lor(v\leq u)$. We say that $U\subseteq V$ is $\leq$-convex if for every $u,v,z\in V$, if $u\leq v\leq z$ and $u,z\in U$, then $v\in U$. For every $u,v\in V$, we write $u<v$ if $(u\leq v)\land(u\not=v)$.
The most important data structures for solving pattern matching queries on compressed strings (such as the suffix array [MW92], the Burrows-Wheeler transform [BW94] and the FM-index [FM05]) are closely related to the idea of sorting strings. Consequently, as customary in the literature on Wheeler automata [ADPP20, CP21], we assume that there exists a fixed total order $\preceq$ on the alphabet $\Sigma$ (in our examples, we always assume $a\prec b\prec c\prec\dots$), and $\preceq$ is extended co-lexicographically to $\Sigma^{*}$ (that is, for every $\alpha,\beta\in\Sigma^{*}$ we have $\alpha\prec\beta$ if and only if the reverse string $\alpha^{R}$ is lexicographically smaller than the reverse string $\beta^{R}$).
For $i\geq 0$, let $\Sigma^{i}\subseteq\Sigma^{*}$ be the set of all strings of length $i$ on the alphabet $\Sigma$. If $\alpha,\beta\in\Sigma^{*}$, we write $\alpha\dashv\beta$ if and only if $\alpha$ is a suffix of $\beta$ (equivalently, if and only if there exists $\beta^{\prime}\in\Sigma^{*}$ such that $\beta=\beta^{\prime}\alpha$). If $\alpha\in\Sigma^{*}$, let $\alpha[i]$ be the $i$-th character of $\alpha$ (for $1\leq i\leq|\alpha|$), let $\alpha[i,j]=\alpha[i]\alpha[i+1]\dots\alpha[j-1]\alpha[j]$ (for $1\leq i\leq j\leq|\alpha|$), and let $p_{i}(\alpha)$ and $s_{i}(\alpha)$ be the prefix and the suffix of $\alpha$ of length $i$, respectively (for $0\leq i\leq|\alpha|$). If $\mathcal{A}=(Q,E,s,F)$ is a GNFA and $u\in Q$, let $\lambda(u)$ be the set of all strings in $\Sigma^{*}$ labeling an edge reaching $u$ (note that $\lambda(u)\not$$\emptyset$ if $u\not=s$ because every state is reachable from the initial state).
Our data structure results hold in the word RAM model with words of size $w\in\Theta(\log N)$ bits, where $N$ is the input size. When we describe our data structures in detail, we assume to be working with integer alphabets of the form $\Sigma=\{0,1\dots,\sigma-1\}$, where $\preceq$ is the usual total order such that $0\prec 1\prec\dots\prec\sigma-1$. All logarithms are in base 2.
4.2. Wheeler GDFAs
Let us define Wheeler GDFAs. Let $\mathcal{A}=(Q,E,s,F)$ be a GDFA. Let $\preceq_{\mathcal{A}}$ be the reflexive relation on $Q$ such that, for every $u,v\in Q$ with $u\not=v$, we have $u\prec_{\mathcal{A}}v$ if and only if $(\forall\alpha\in I_{u})(\forall\beta\in I_{v})(\alpha\prec\beta)$. Since each $I_{u}$ is nonempty, it is immediate to realize that $\preceq_{\mathcal{A}}$ is a partial order, but in general it is not a total order. We can then give the following definition (see Figure 5 for an example).
{defi}
Let $\mathcal{A}=(Q,E,s,F)$ be a GDFA. We say that $\mathcal{A}$ is Wheeler if $\preceq_{\mathcal{A}}$ is a total order.
If $\mathcal{A}$ is a conventional DFA, it is not immediately clear that Definition 4.2 is equivalent to the local definition of Wheeler DFA commonly used in the literature [ADPP20, CCG+23, CGKP23]. According to the local definition, a DFA $\mathcal{A}=(Q,E,s,F)$ is Wheeler if there exists a total order $\leq$ on $Q$ such that (i) $s$ comes first in the total order, (ii) for every $(u^{\prime},u,a),(v^{\prime},v,b)\in E$, if $u<v$, then $a\preceq b$ and (iii) for every $(u^{\prime},u,a),(v^{\prime},v,a)\in E$, if $u<v$, then $u^{\prime}<v^{\prime}$. Alanko et al. [ADPP20, Corollary 3.1] proved that, if such a total order $\leq$ exists, then it is unique and it is equal to $\preceq_{\mathcal{A}}$, so we only have to prove that if $\preceq_{\mathcal{A}}$ is a total order, then it satisfies properties (i), (ii), (iii). This follows from the next lemma.
[…]
Figure 6. Left: A Wheeler GDFA $\mathcal{A}=(Q,E,s,F)$. The states are numbered following the total order $\preceq_{\mathcal{A}}$. Note that $(u_{1},u_{2},ba),(u_{4},u_{3},a)\in E$, $u_{2}\prec_{\mathcal{A}}u_{3}$, $a$ is a strict suffix of $ba$ and $a\prec ba$. Right: A GDFA $\mathcal{A}=(Q,E,s,F)$ for which the states are numbered following a total order $\leq$ such that (i) $s$ comes first in the total order $\leq$, (ii) for every $(u^{\prime},u,\rho),(v^{\prime},v,\rho^{\prime})\in E$, if $u<v$ and $\rho^{\prime}$ is not a strict suffix of $\rho$, then $\rho\preceq\rho^{\prime}$ and (iii) for every $(u^{\prime},u,\rho),(v^{\prime},v,\rho)\in E$, if $u<v$, then $u^{\prime}<v^{\prime}$. Note that $\mathcal{A}$ is not Wheeler because $u_{2}$ and $u_{3}$ are not $\preceq_{\mathcal{A}}$-comparable, since $b\prec c\prec ac$, $c\in I_{u_{3}}$ and $b,ac\in I_{u_{2}}$.
Lemma 13.
Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler GDFA. Then:
  1. (1)
    $s$ comes first in the total order $\preceq_{\mathcal{A}}$.
  2. (2)
    For every $(u^{\prime},u,\rho),(v^{\prime},v,\rho^{\prime})\in E$, if $u\prec_{\mathcal{A}}v$ and $\rho^{\prime}$ is not a strict suffix of $\rho$, then $\rho\preceq\rho^{\prime}$.
  3. (3)
    For every $(u^{\prime},u,\rho),(v^{\prime},v,\rho)\in E$, if $u\prec_{\mathcal{A}}v$, then $u^{\prime}\prec_{\mathcal{A}}v^{\prime}$.
Proof 4.1.
  1. (1)
    Let $u\in Q\setminus\{s\}$. We must prove that $s\prec_{\mathcal{A}}u$. Since $\preceq_{\mathcal{A}}$ is a total order, we only have to prove that if $s$ and $u$ are $\preceq_{\mathcal{A}}$-comparable, then it must be $s\prec_{\mathcal{A}}u$. This follows from the fact that for every $\alpha\in I_{u}$ we have $\epsilon\prec\alpha$, and $\epsilon\in I_{s}$.
  2. (2)
    If $\rho=\rho^{\prime}$ or $\rho$ is a strict suffix of $\rho^{\prime}$, then the conclusion is immediate, so we can assume that $\rho\not=\rho^{\prime}$, $\rho$ is not a strict suffix of $\rho^{\prime}$, and $\rho^{\prime}$ is not a strict suffix of $\rho$. Let $\alpha^{\prime}\in I_{u^{\prime}}$ and $\beta^{\prime}\in I_{v^{\prime}}$. Then, $\alpha^{\prime}\rho\in I_{u}$ and $\beta^{\prime}\rho^{\prime}\in I_{v}$. Since $u\prec_{\mathcal{A}}v$, then $\alpha^{\prime}\rho\prec\beta^{\prime}\rho^{\prime}$, Since $\rho\not=\rho^{\prime}$, $\rho$ is not a strict suffix of $\rho^{\prime}$, and $\rho^{\prime}$ is not a strict suffix of $\rho$, we conclude $\rho\prec\rho^{\prime}$.
  3. (3)
    Let $\alpha^{\prime}\in I_{u^{\prime}}$ and $\beta^{\prime}\in I_{v^{\prime}}$. We must prove that $\alpha^{\prime}\prec\beta^{\prime}$. From $(u^{\prime},u,\rho),(v^{\prime},v,\rho)\in E$ we obtain $\alpha^{\prime}\rho\in I_{u}$ and $\beta^{\prime}\rho\in I_{v}$, so from $u\prec_{\mathcal{A}}v$ we obtain $\alpha^{\prime}\rho\prec\beta^{\prime}\rho$ and thus $\alpha^{\prime}\prec\beta^{\prime}$.
In case 2 of Lemma 13 we cannot remove the assumption that $\rho^{\prime}$ is not a strict prefix of $\rho$ (see Figure 6); as a consequence, we cannot use Lemma 13 to provide an equivalent, local definition of Wheeler GDFA. The local definition of Wheeler DFA easily implies that the problem of deciding whether a given DFA is Wheeler can be solved in polynomial time [ADPP20], but since we do not have a local definition of Wheeler GDFA, it is not clear whether the corresponding problem on GDFAs is also solvable in polynomial time (and we saw in the introduction that computational problems on generalized automata are usually hard). However, in Lemma 14 below, we prove that the problem is still solvable in polynomial time by reducing it to the problem of computing the partial order $\preceq_{\mathcal{A^{*}}}$ on a conventional DFA $\mathcal{A^{*}}$ equivalent to a given GDFA $\mathcal{A}$.
To follow the proof of Lemma 14, let us recall tries and their properties. Let $C\subseteq\Sigma^{+}$ be a nonempty finite set, and assume that $C$ is prefix-free. The trie $\mathcal{T}_{C}$ of $C$ (see [Nav16]) is the unique rooted directed tree such that (i) every edge is labeled with a character, (ii) two edges leaving the same node have distinct label, (iii) $\mathcal{T}_{C}$ contains $|C|$ leaves and (iv) for every leaf $z$, the string $\rho$ that can be read following the unique path from the root to $z$ is in $C$ (see Figure 7 for an example). For every node $u$ of $\mathcal{T}_{C}$, let $\tau_{u}$ be the string that can be read by following the unique path from the root to $u$ (in particular, $\tau_{u}\in C$ if and only if $u$ is a leaf of $\mathcal{T}_{C}$). For example, in Figure 7 we have $\tau_{5}=ac$. If $C=\emptyset$, we assume that $\mathcal{T}_{C}$ consists of a single node and no edges.
1122334455aabbaacc
Figure 7. The trie of the prefix-free set $C=\{aa,ac,b\}$.
Consider a prefix-free set $C=\{\rho_{1},\rho_{2},\dots,\rho_{h}\}$ of size $h\geq 1$, and assume that we know that $\rho_{1}\prec\rho_{2}\prec\dots\prec\rho_{h}$. Let us show that in $O(\sum_{i=1}^{h}|\rho_{i}|)$ time we can (i) build (the adiacency-list representation of) $\mathcal{T}_{C}$ and (ii) associate to every $1\leq i\leq h$ the unique leaf $u$ of $\mathcal{T}_{C}$ such that $\tau_{u}=\rho_{i}$. We can identify each node $u$ of $\mathcal{T}_{C}$ with the triple $\phi_{u}=(k_{u},\ell_{u},r_{u})$, where $k_{u}$ is the distance of $u$ from the root and $1\leq\ell_{u}\leq r_{u}\leq h$ are the two integers such that for every $1\leq i\leq h$ we have that $\tau_{u}$ is a prefix of $\rho_{i}$ if and only if $\ell_{u}\leq i\leq r_{u}$. For example, in Figure 7 we have $\tau_{2}=a$ and $\phi_{2}=(1,1,2)$. Note that for every node of $u$ of $\mathcal{T}_{C}$ we have $\rho_{i}[1,k_{u}]=\tau_{u}$ if and only if $\ell_{u}\leq i\leq r_{u}$. Moreover, for every node $u$ of $\mathcal{T}_{C}$, we have that $u$ is a leaf if and only if (i) $\ell_{u}=r_{u}$ and (ii) $|\rho_{\ell_{u}}|=k_{u}$; if $u$ is not a leaf, then $|\rho_{i}|\geq k_{u}+1$ for every $\ell_{u}\leq i\leq r_{u}$ (because $C$ is prefix-free). We build $\mathcal{T}_{C}$ using a queue. At any time, each element in the queue is equal to $\phi_{u}$ for some node $u$. At the beginning of the algorithm, we add $(0,1,h)$ to $\mathcal{T}_{C}$ (which corresponds to the root of $\mathcal{T}_{C}$), and we enqueue $(0,1,h)$. We now process all elements of the queue until it becomes empty. Assume that we pop $(k,\ell,r)$ from the queue. If $(k,\ell,r)$ corresponds to a leaf (which can be checked through the characterization mentioned earlier), we record that $\ell$ is associated with the node $(k,\ell,r)$, and we pop the next element of the queue. Now, assume that $(k,\ell,r)$ is not a leaf. Then, we have $|\rho_{i}|\geq k+1$ for every $\ell\leq i\leq r$. Our goal is to determine the children of $(k,\ell,r)$. We scan $\rho_{\ell}[k+1]$, $\rho_{\ell+1}[k+1]$, $\dots$, $\rho_{r-1}[k+1]$, $\rho_{r}[k+1]$, and we compute the set $D=\{\ell\}\cup\{\ell+1\leq i\leq r\;|\;\rho_{i}[k+1]\not=\rho_{i-1}[k+1]\}$. Let $D=\{i_{1},i_{2},\dots,i_{q}\}$, with $q\geq 1$ and $i_{1}<i_{2}<\dots<i_{q}$. Then, for every $1\leq j\leq q$, we add the new node $(k+1,i_{j},i_{j+1}-1)$ to $\mathcal{T}_{C}$ (where $i_{q+1}=r+1$), we add an edge from $(k,\ell,r)$ to $(k+1,i_{j},i_{j+1}-1)$ labeled $\rho_{i_{j}}[k+1]$, and we enqueue $(k+1,i_{j},i_{j+1}-1)$. We can now pop the next element of the queue. When the queue is empty, we have correctly computed $\mathcal{T}_{C}$ in $O(\sum_{i=1}^{h}|\rho_{i}|)$ time, and we have associated to every $1\leq i\leq h$ the unique leaf $u$ of $\mathcal{T}_{C}$ such that $\tau_{u}=\rho_{i}$.
We are now ready to prove Lemma 14.
[…]
Figure 8. Top left: The GDFA $\mathcal{A}$ of Figure 5. Top right: The DFA $A^{*}$ built starting from $\mathcal{A}$ in the proof of Lemma 14. Bottom: An arborescence $\mathcal{A}^{\prime}$ obtained from $A^{*}$.
Lemma 14.
Let $\mathcal{A}=(Q,E,s,F)$ be a GDFA, and let $\mathfrak{e}$ be the total length of all edge labels. In $O(\mathfrak{e}\log\mathfrak{e})$ time we can decide whether $\mathcal{A}$ is Wheeler and, if so, we can compute $\preceq_{\mathcal{A}}$ (that is, we can sort the elements of $Q$ with respect to $\preceq_{\mathcal{A}}$). The bound $O(\mathfrak{e}\log\mathfrak{e})$ is also true in the comparison model.
Proof 4.2.
In the rest of the proof, we can assume without loss of generality that the alphabet $\Sigma$ is an integer alphabet (that is, $\Sigma=\{1,2,\dots,\sigma\}$, with $\sigma=|\Sigma|$) and is effective (that is, every character in $\Sigma$ occurs in at least one edge label, and in particular $\sigma\leq\mathfrak{e}$). Indeed, even in the more general case of the comparison model, we can sort all characters labeling some edges in $O(\mathfrak{e}\log\mathfrak{e})$ time via any comparison-based sorting algorithm (e.g., merge sort, see [CLRS22]) and we can replace each character with its rank in the sorted list of all characters, which does not affect the partial order $\preceq_{\mathcal{A}}$.
Next, we can assume without loss of generality that, for every $u\in V$, all the edges leaving $u$ are sorted with respect to the lexicographic order of their labels. Indeed, if the edges are not sorted, we can sort all edges of $\mathcal{A}$ by first comparing their start states and then comparing their labels. This can be achieved in $O(|Q|+\sigma+\mathfrak{e})=O(\mathfrak{e})$ time via radix sort [HUA83, PT87] (recall that $\sigma\leq\mathfrak{e}$ because the alphabet is effective and $|Q|\leq|E|+1\leq\mathfrak{e}+1$ because (i) every state is reachable from the initial state and (ii) every edge label is a nonempty string by the definition of GDFA).
Let us build a DFA $\mathcal{A}^{*}$ starting from $A$ (see Figure 8 for an example). For every $u\in Q$, let $C_{u}$ be the prefix-free set of all strings labeling some edge leaving $u$. In particular, if $u$ has no outgoing edges, then $C_{u}$ consists of a single node and no edges. Note that the number of edges in the trie $\mathcal{T}_{C_{u}}$ of $C_{u}$ is upper-bounded by the sum of the lengths of all edges leaving $u$.
We define $\mathcal{A}^{*}=(Q^{*},E^{*},s^{*},F^{*})$ as follows. We first consider the trie $\mathcal{T}_{C_{u}}$ for every $u\in Q$. Then, for every $(u^{\prime},u,\rho)\in E$, we pick the unique leaf $v$ of $\mathcal{T}_{C_{u^{\prime}}}$ such that $\tau_{v}=\rho$ (that is, the leaf associated to $\rho$), we remove the leaf $v$, and we redirect the unique edge of $\mathcal{T}_{C_{u}}$ reaching $v$ to the root of $\mathcal{T}_{C_{u}}$. The initial state $s^{*}$ is the root of $\mathcal{T}_{C_{s}}$ and a state is in $F^{*}$ if and only it is the root of a trie $\mathcal{T}_{C_{u}}$ for some $u\in F$. By construction, $\mathcal{A}^{*}$ is a DFA (in particular, every state of $\mathcal{A}^{*}$ is both reachable and co-reachable). If we identify each $u\in Q$ with the root of $\mathcal{T}_{C_{u}}$ (which is a state in $Q^{*}$), then we have $I_{u}^{\mathcal{A}}=I_{u}^{\mathcal{A^{*}}}$ for every $u\in Q$, so we conclude $\mathcal{L(A^{*})}=\mathcal{L(A)}$. Moreover, $|Q^{*}|-1\leq|E^{*}|\leq\mathfrak{e}$ because (i) every state of $Q^{*}$ is reachable from $s^{*}$ and (ii) for every $u\in V$ the number of edges in the trie $\mathcal{T}_{C_{u}}$ of $C_{u}$ is upper-bounded by the sum of the lengths of all edges leaving $u$.
Let us show that we can build $\mathcal{A}^{*}$ in $O(\mathfrak{e})$ time. For every $u\in Q$, all the edges leaving $u$ are sorted with respect to the lexicographic order of their labels, so we can build all the $\mathcal{T}_{C_{u}}$’s (including the $\mathcal{T}_{C_{u}}$’s for which $u$ has no outgoing edges) in $O(|Q|+\mathfrak{e})=O(\mathfrak{e})$ time, and for every $u\in V$ and for every $\rho\in C_{u}$ we record the corresponding leaf of $\mathcal{T}_{C_{u}}$. Through this information, we can correctly redirect the edge reaching each leaf in $O(\mathfrak{e})$ time, and within the same time bound we can store $F^{*}$.
Since $I_{u}^{\mathcal{A}}=I_{u}^{\mathcal{A^{*}}}$ for every $u\in Q$, we conclude that $\mathcal{A}$ is Wheeler if and only if the restriction of the partial order $\preceq_{\mathcal{A}^{*}}$ to $Q$ is a total order.
Consider the DFA $\mathcal{A}^{\prime}=(Q^{\prime},E^{\prime},s^{\prime},F^{\prime})$ obtained from $\mathcal{A}^{*}$ in $O(|E^{*}|)=O(\mathfrak{e})$ time as follows (see Figure 8). We define $Q^{\prime}=Q^{*}$, $s^{\prime}=s^{*}$ and $F^{\prime}=Q^{\prime}$. To define $E^{\prime}\subseteq E^{*}$, we navigate $\mathcal{A}^{*}$ starting from $s^{*}$ and we visit all states of $\mathcal{A}^{*}$, discarding all edges that reach a state that we have already visited (in other words, $\mathcal{A}^{\prime}$ is any arborescence of $\mathcal{A}^{*}$). Then, for every $u\in Q^{\prime}$ the set $I_{u}^{\mathcal{A}^{\prime}}$ contains exactly one string $\gamma_{u}$ (because $s^{\prime}$ has no incoming edges, all the remaining states of $Q^{\prime}$ have exactly one incoming edge, and every state is reachable from $s^{\prime}$), and $\gamma_{u}\in I_{u}^{\mathcal{A}^{*}}$. Notice that the $\gamma_{u}$’s are pairwise distinct because $\mathcal{A}^{\prime}$ is a DFA. Since $|I_{u}^{\mathcal{A}^{\prime}}|=1$ for every $u\in Q^{\prime}$, then $\mathcal{A}^{\prime}$ is Wheeler. Moreover, we can sort the elements of $Q^{\prime}$ with respect to the $\gamma_{u}$’s (which yields the total order $\preceq_{\mathcal{A}^{\prime}}$) in $O(|Q^{*}|)=O(\mathfrak{e})$ time, as shown by Ferragina et al. in their construction algorithm for the XBWT [FLMM09]. Then, in $O(|Q^{*}|)=O(\mathfrak{e})$ time we compute the restriction of $\preceq_{\mathcal{A}^{\prime}}$ to $Q$. In other words, we now know the enumeration $q_{1}$, $q_{2}$, $\dots$, $q_{|Q|}$ such that (i) $Q=\{q_{1},q_{2},\dots,q_{|Q|}\}$ and (ii) $q_{1}\prec_{\mathcal{A}^{\prime}}q_{2}\prec_{\mathcal{A}^{\prime}}q_{3}\prec\dots\prec q_{|Q|}$.
For every $u\in Q^{*}$ (and in particular, for every $u\in Q$) we have $\gamma_{u}\in I_{u}^{\mathcal{A}^{*}}$. Consequently, for every $u,v\in Q$, if $u\preceq_{\mathcal{A}^{*}}v$, then $u\preceq_{\mathcal{A}^{\prime}}v$. Hence, $\mathcal{A}$ is Wheeler if and only if the restriction of the partial order $\preceq_{\mathcal{A}^{*}}$ to $Q$ is a total order, if and only if $q_{i}\prec_{\mathcal{A}^{*}}q_{i+1}$ for every $1\leq i\leq|Q|-1$, and if $\mathcal{A}$ is Wheeler, then the enumeration $q_{1}$, $q_{2}$, $\dots$, $q_{|Q|}$ yields $\preceq_{\mathcal{A}}$. To conclude the proof, we only have to show how to check whether $q_{i}\prec_{\mathcal{A}^{*}}q_{i+1}$ for every $1\leq i\leq|Q|-1$.
We know that $\mathcal{A}^{*}$ is a conventional DFA, so we can compute the partial order $\preceq_{\mathcal{A}^{*}}$ in polynomial time [CP21, KOP23, Cot23, BCC+23]. More precisely, the algorithm in [BCC+23] runs in $O(|E^{*}|\log|Q^{*}|)$ time (and so $O(\mathfrak{e}\log\mathfrak{e})$ time), returning a data structure that in $O(1)$ time supports the following query: given two distinct states $u,v\in Q^{*}$, decide whether (i) $u\prec_{\mathcal{A}^{*}}v$, (ii) $v\prec_{\mathcal{A}^{*}}u$ or (iii) $u$ and $v$ are not $\preceq_{\mathcal{A}^{*}}$-comparable. Hence, in $O(|Q|)=O(\mathfrak{e})$ time we can check whether $q_{i}\prec_{\mathcal{A}^{*}}q_{i+1}$ for every $1\leq i\leq|Q|-1$.
Lemma 14 shows that, if $\mathcal{A}$ is Wheeler, then we can compute $\preceq_{\mathcal{A}}$ in $O(\mathfrak{e}\log\mathfrak{e})$ time. In the comparison model, this bound is optimal: we cannot compute $\preceq_{\mathcal{A}}$ is $o(\mathfrak{e}\log\mathfrak{e})$ time, otherwise we would break the well-known lower bound $\Omega(n\log n)$ to sort $n$ elements, which holds even if we know that the $n$ elements are pairwise distinct [Knu98, CLRS22]. Indeed, in $O(n)$ time we can reduce the problem of sorting $n$ distinct elements to the problem of computing $\preceq_{\mathcal{A}}$ for a DFA $\mathcal{A}$ consisting of a single path (see Figure 9). Another proof of the same lower bound can be obtained as follows. Observe that the problem of computing $\preceq_{\mathcal{A}}$ is a generalization of the problem of computing the suffix array of a string (that is, the problem of lexicographically sorting all the suffixes of a string), which has complexity $\Omega(n\log n)$ in the comparison model [FCFM00]. More precisely, given a string $\alpha$, the problem of computing the suffix array of $\alpha$ is equivalent to the problem of computing $\preceq_{\mathcal{A}}$ for a DFA $\mathcal{A}$ consisting of a single path obtained by considering the reverse string $\alpha^{R}$ (see Figure 9).
In the case of an integer alphabet in a polynomial range, the suffix array can be built in linear time [FCFM00], and we leave it as an open problem to determine whether in Lemma 14 we can achieve $O(\mathfrak{e})$ time. In particular, in the case of an integer alphabet in a polynomial range, we can sort in linear time [HUA83, PT87], so the proof of Lemma 14 implies that the bound $O(\mathfrak{e})$ would follow immediately if we proved that, given a DFA $\mathcal{A}=(Q,E,s,F)$, we can compute the partial order $\preceq_{\mathcal{A}}$ in $O(|E|)$ time (determining if this is possible is also an open problem).
startccddaabb
startddffddcc
Figure 9. Left: Reducing the problem of sorting $c$, $d$, $a$, $b$ to the problem of computing $\preceq_{\mathcal{A}}$. Right: Reducing the problem of computing the suffix array of $\alpha=cdfd$ to the problem of computing $\preceq_{\mathcal{A}}$ (where $\alpha^{R}=dfdc$).
4.3. The Burrows-Wheeler Transform of a Wheeler GDFA
Let $\mathcal{A}=(Q,E,s,F)$ be a GDFA with $n=|Q|$ states. We say that $\mathcal{A}$ is an $r$-GDFA if all edge labels have length at most $r$. Fix $1\leq i\leq r$. Let $E_{i}=\{(u^{\prime},u,\rho)\in E\;|\;\rho\in\Sigma^{i}\}$ be the set of all edges labeled with a string of length $i$, and let $\Sigma_{i}=\{\rho\in\Sigma^{i}\;|\;(\exists u^{\prime},u\in Q)((u^{\prime},u,\rho)\in E_{i})\}$ be the set of all strings of length $i$ labeling some edge. Let $e_{i}=|E_{i}|$ and $\sigma_{i}=|\Sigma_{i}|$; we have $\sigma_{i}\leq\min\{\sigma^{i},e_{i}\}$ (where $\sigma^{i}$ is $\sigma$ to the $i$-th power). The $i$-outdegree (ii-indegree, respectively) of a state is equal to the number of edges in $E_{i}$ leaving (reaching, respectively) the state. The sum of the $i$-outdegrees of all the states and the sum of the $i$-indegrees of all the states are both equal to $e_{i}$. Lastly, $\sum_{i=1}^{r}e_{i}=e$ and the total length of all edge labels is $\mathfrak{e}=\sum_{i=1}^{r}e_{i}i$.
If $\mathcal{A}=(Q,E,s,F)$ is a Wheeler GDFA, we write $Q=\{Q[1],Q[2],\dots,Q[n]\}$, where $Q[1]\prec_{\mathcal{A}}Q[2]\prec_{\mathcal{A}}\dots\prec_{\mathcal{A}}Q[n]$. If $1\leq j_{1}\leq j_{2}\leq n$, let $Q[j_{1},j_{2}]=\{Q[j_{1}],Q[j_{1}+1],\dots,Q[j_{2}-1],Q[j_{2}]\}$, and if $j_{1}>j_{2}$, let $Q[j_{1},j_{2}]=\emptyset$.
Let us define the Burrows-Wheeler Transform (BWT) of a Wheeler GDFA (see Figure 5 for an example).
{defi}
[BWT of a Wheeler GDFA] Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler GDFA. The Burrows-Wheeler Transform $\operatorname{BWT}(\mathcal{A})$ of $\mathcal{A}$ consists of the following strings.
  • For every $1\leq i\leq r$, the bit string $\mathtt{OUT}_{i}\in\{0,1\}^{e_{i}+n}$ that stores the $i$-outdegrees in unary. More precisely, (i) $\mathtt{OUT}_{i}$ contains exactly $n$ characters equal to $1$, (ii) $\mathtt{OUT}_{i}$ contains exactly $e_{i}$ characters equal to 0, and (iii) for every $1\leq j\leq n$, the number of zeros between the (j1)(j-1)-th character equal to one (or the beginning of the sequence if $j=1$) and the $j$-th character equal to $1$ yields the $i$-outdegree of $Q[j]$.
  • For every $1\leq i\leq r$, the bit string $\mathtt{IN}_{i}\in\{0,1\}^{e_{i}+n}$ that stores the $i$-indegrees in unary. More precisely, (i) $\mathtt{IN}_{i}$ contains exactly $n$ characters equal to $1$, (ii) $\mathtt{IN}_{i}$ contains exactly $e_{i}$ characters equal to 0, and (iii) for every $1\leq j\leq n$, the number of zeros between the (j1)(j-1)-th character equal to one (or the beginning of the sequence if $j=1$) and the $j$-th character equal to $1$ yields the $i$-indegree of $Q[j]$.
  • For every $1\leq i\leq r$, the string $\mathtt{LAB}_{i}\in(\Sigma_{i})^{e_{i}}$ that stores the edge labels of length $i$ (with their multiplicities). More precisely, we sort of all edges in $E_{i}$ by the index of the start states (w.r.t $\preceq_{\mathcal{A}}$). Edges with the same start state are sorted by label. Then, we obtain $\mathtt{LAB}_{i}$ by concatenating the labels of all edges following this edge order.
  • The bit string $\mathtt{FIN}\in\{0,1\}^{n}$ that marks the final states, that is, for every $1\leq j\leq n$, the $j$-th bit of $\mathtt{FIN}$ is equal to 1 if and only if $Q[j]\in F$.
We can now prove that the BWT of a Wheeler GDFA $\mathcal{A}$ is a valid encoding of $\mathcal{A}$ (recall that the BWT of a string is a valid encoding of the string [BW94]).
Theorem 15.
Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler GDFA. If we only know $\operatorname{BWT}(\mathcal{A})$, then we can retrieve $\mathcal{A}$ (up to isomorphism). In other words, $\operatorname{BWT}(\mathcal{A})$ is an encoding of $\mathcal{A}$.
Proof 4.3.
Note that $\operatorname{BWT}(\mathcal{A})$ consists of $3r+1$ strings, so from $\operatorname{BWT}(\mathcal{A})$ we can retrieve $r$. From $\operatorname{BWT}(\mathcal{A})$ we can also retrieve the value $n$ and, for every $1\leq i\leq r$, the value $e_{i}$.
Let us show that (i) for every $1\leq j\leq n$, we can determine whether $Q[j]\in F$, and (ii) for every $1\leq j^{\prime},j\leq n$ and for every $\rho\in\Sigma^{+}$ such that $|\rho|\leq r$, we can determine whether $(Q[j^{\prime}],Q[j],\rho)\in E$. This is sufficient to retrieve $\mathcal{A}$ up to isomorphism. Indeed, if we define the generalized automaton $\mathcal{A}^{\prime}=(Q^{\prime},E^{\prime},s^{\prime},F^{\prime})$ such that $Q^{\prime}=\{1,2,\dots,n\}$, $E^{\prime}=\{(j^{\prime},j,\rho)\in Q^{\prime}\times Q^{\prime}\times\Sigma^{*}\;|\;(Q[j^{\prime}],Q[j],\rho)\in E\}$, $s^{\prime}=1$, and $F^{\prime}=\{j\in Q\;|\;Q[j]\in F\}$, then $\mathcal{A}^{\prime}$ is isomorphic to $\mathcal{A}$, with isomorphism $\phi:Q^{\prime}\mapsto Q$ given by $\phi(j)=Q[j]$ for every $j\in Q^{\prime}$ (note that $\phi(s^{\prime})=s$ because $s=Q[1]$ by Lemma 13).
Proving (i) is immediate because we can use $F$. To prove (ii), we must show how to retrieve $E$. It will suffice to retrieve the set $E_{i}$ for every $1\leq i\leq r$, because $E$ is the (disjoint) union of the $E_{i}$’s. Fix $1\leq i\leq r$. From $\mathtt{LAB}_{i}$ we can retrieve the labels of all edges in $E_{i}$, with their multiplicities. From $\mathtt{IN}_{i}$ we can retrieve the $i$-indegree of each $Q[j]$. By Lemma 13, for every $\rho\in\Sigma^{i}$ labeling some edge reaching some state $Q[j]$ and for every $\rho^{\prime}\in\Sigma^{i}$ labeling some edge reaching $Q[j+1]$ it must be $\rho\preceq\rho^{\prime}$. Since we know the labels of all edges in $E_{i}$ with multiplicities and we know the $i$-indegrees, then we can retrieve the labels of all edges reaching each $Q[j]$, with multiplicities. From $\mathtt{OUT}_{i}$ we can retrieve the $i$-outdegrees of each $Q[j]$, and the order used in the definition of $\mathtt{LAB}_{i}$ implies that we can retrieve the labels of all edges leaving each $Q[j]$. Since we know the labels of all edges reaching each $Q[j]$ and we know the labels of all edges leaving each $Q[j]$, then for every $\rho\in\Sigma^{i}$ we know the set of all states reached by an edge labeled $\rho$, with multiplicities, and the set of all states having an outgoing edge labeled $\rho$. By Lemma 13, for every $1\leq j_{1}<j_{2}\leq n$, if $Q[j_{1}]$ is reached by an edge labeled $\rho$ leaving the state $Q[j^{\prime}_{1}]$ and $Q[j_{2}]$ is reached by an edge labeled $\rho$ leaving the state $Q[j^{\prime}_{2}]$, then it must be $j^{\prime}_{1}<j^{\prime}_{2}$. As a consequence, we can retrieve the set $E_{i,\rho}$ of all edges labeled $\rho$ for every $\rho\in\Sigma^{i}$, and so we can retrieve $E_{i}$, because $E_{i}$ is the (disjoint) union of the $E_{i,\rho}$’s.
4.4. The FM-index of a Wheeler GDFA
Let us give the following definition.
{defi}
Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler GDFA, and let $\alpha\in\Sigma^{*}$. Define:
  • $G^{\prec}(\alpha)=\{u\in Q\;|\;(\forall\beta\in I_{u})(\beta\prec\alpha)\}$;
  • $G_{\dashv}(\alpha)=\{u\in Q\;|\;(\exists\beta\in I_{u})(\alpha\dashv\beta)\}$;
  • $G^{\prec}_{\dashv}(\alpha)=G^{\prec}(\alpha)\cup G_{\dashv}(\alpha)=\{u\in Q\ \;|\;(\forall\beta\in I_{u})(\beta\prec\alpha)\vee(\exists\beta\in I_{u})(\alpha\dashv\beta)\}$.
Intuitively, the set $G_{\dashv}(\alpha)$ is the set of states that the SMLG problem must return on input $\alpha$, and $G^{\prec}(\alpha)$ is the set of all states reached only by strings smaller than $\alpha$. For example, in Figure 5 we have $G^{\prec}(ac)=\{u_{1},u_{2}\}$, $G_{\dashv}(ac)=\{u_{3}\}$ and $G^{\prec}_{\dashv}(ac)=\{u_{1},u_{2},u_{3}\}$.
Remark 16.
Note that $G^{\prec}(\epsilon)=\emptyset$ (because $\epsilon$ is the smallest string in $\Sigma^{*}$) and $G_{\dashv}(\epsilon)=Q$ (because $\epsilon$ is a suffix of every string in $\Sigma^{*}$), so $G^{\prec}_{\dashv}(\epsilon)=Q$. In particular, $|G^{\prec}(\epsilon)|=0$ and $|G_{\dashv}(\epsilon)|=|G^{\prec}_{\dashv}(\epsilon)|=n$, where $n=|Q|$.
The following lemma shows that, as in the case of conventional Wheeler automata [ADPP21], both $G^{\prec}(\alpha)$ and $G_{\dashv}(\alpha)$ are intervals with respect to the total order $\preceq_{\mathcal{A}}$, and there is no state between $G^{\prec}(\alpha)$ and $G_{\dashv}(\alpha)$.
Lemma 17.
Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler GDFA, and let $\alpha\in\Sigma^{*}$. Then:
  1. (1)
    $G^{\prec}(\alpha)\cap G_{\dashv}(\alpha)=\emptyset$.
  2. (2)
    $G_{\dashv}(\alpha)$ is $\preceq_{\mathcal{A}}$-convex.
  3. (3)
    If $u,v\in Q$ are such that $u\prec_{\mathcal{A}}v$ and $v\in G^{\prec}(\alpha)$, then $u\in G^{\prec}(\alpha)$. In other words, $G^{\prec}(\alpha)=Q[1,|G^{\prec}(\alpha)|]$.
  4. (4)
    If $u,v\in Q$ are such that $u\prec_{\mathcal{A}}v$ and $v\in G^{\prec}_{\dashv}(\alpha)$, then $u\in G^{\prec}_{\dashv}(\alpha)$. In other words, $G^{\prec}_{\dashv}(\alpha)=Q[1,|G^{\prec}_{\dashv}(\alpha)|]$.
  5. (5)
    $G_{\dashv}(\alpha)=Q[|G^{\prec}(\alpha)|+1,|G^{\prec}_{\dashv}(\alpha)|]$.
Proof 4.4.
  1. (1)
    If $u\in G_{\dashv}(\alpha)$, then there exists $\beta\in I_{u}$ such that $\alpha\dashv\beta$. In particular, $\alpha\preceq\beta$, so $u\not\in G^{\prec}(\alpha)$.
  2. (2)
    Assume that $u,v,z\in Q$ are such that $u\prec_{\mathcal{A}}v\prec_{\mathcal{A}}z$ and $u,z\in G_{\dashv}(\alpha)$. We must prove that $v\in G_{\dashv}(\alpha)$. Since $u,z\in G_{\dashv}(\alpha)$, then there exist $\beta\in I_{u}$ and $\delta\in I_{z}$ such that $\alpha\dashv\beta$ and $\alpha\dashv\delta$. Fix any $\gamma\in I_{v}$; we only have to prove that $\alpha\dashv\gamma$. From $u\prec_{\mathcal{A}}v\prec_{\mathcal{A}}z$ we obtain $\beta\prec\gamma\prec\delta$. As a consequence, from $\alpha\dashv\beta$ and $\alpha\dashv\delta$ we conclude $\alpha\dashv\gamma$.
  3. (3)
    Let $\beta\in I_{u}$. We must prove that $\beta\prec\alpha$. Fix any $\gamma\in I_{v}$. Since $u\prec_{\mathcal{A}}v$, we have $\beta\prec\gamma$. From $v\in G^{\prec}(\alpha)$ we obtain $\gamma\prec\alpha$, so we conclude $\beta\prec\alpha$.
  4. (4)
    Since $v\in G^{\prec}_{\dashv}(\alpha)$, we have either $v\in G^{\prec}(\alpha)$ or $v\in G_{\dashv}(\alpha)$. If $v\in G^{\prec}(\alpha)$, then $u\in G^{\prec}(\alpha)$ by the previous point and so $u\in G^{\prec}_{\dashv}(\alpha)$. Now assume that $v\in G_{\dashv}(\alpha)$. If $u\in G_{\dashv}(\alpha)$, then $u\in G^{\prec}_{\dashv}(\alpha)$ and we are done, so we can assume $u\not\in G_{\dashv}(\alpha)$. Let us prove that it must be $u\in G^{\prec}(\alpha)$, which again implies $u\in G^{\prec}_{\dashv}(\alpha)$. Fix $\beta\in I_{u}$; we must prove that $\beta\prec\alpha$. Since $v\in G_{\dashv}(\alpha)$, then there exists $\gamma\in\Sigma^{*}$ such that $\gamma\alpha\in I_{v}$. From $u\prec_{\mathcal{A}}v$ we obtain $\beta\prec\gamma\alpha$. Since $u\not\in G_{\dashv}(\alpha)$ implies $\lnot(\alpha\dashv\beta)$, from $\beta\prec\gamma\alpha$ we conclude $\beta\prec\alpha$.
  5. (5)
    By the definition of $G^{\prec}_{\dashv}(\alpha)$ and the first point we obtain that $G_{\dashv}(\alpha)$ is the disjoint union of $G^{\prec}(\alpha)$ and $G_{\dashv}(\alpha)$, so the conclusion follows from the third point and the fourth point.
Since $G_{\dashv}(\alpha)=Q[|G^{\prec}(\alpha)|+1,|G^{\prec}_{\dashv}(\alpha)|]$, to compute $G_{\dashv}(\alpha)$, it will suffice to compute $|G^{\prec}(\alpha)|$ and $|G^{\prec}_{\dashv}(\alpha)|$. To this end, we will repeatedly use Property 3 in Lemma 13, which is also crucial for conventional Wheeler automata (it is a generalization of the LF mapping used in the FM-index [FM05]). However, the data structures and the algorithm required for solving the SMLG problem on Wheeler GDFAs are significantly more complex than those required for conventional Wheeler automata.
First, let us show how to compute $|G^{\prec}(\alpha)|$. The next lemma formalizes the following intuition: to compute $|G^{\prec}(\alpha)|$, we have to consider all states whose incoming edges have a label smaller than $\alpha$; moreover, if the label is $s_{k}(\alpha)$ (for some $k$), then the start state must be in $G^{\prec}(p_{|\alpha|-k}(\alpha))$.
Lemma 18.
Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler GDFA, and let $\alpha\in\Sigma^{*}$, with $\alpha\not=\epsilon$. Then, $u\in G^{\prec}(\alpha)$ if and only if (i) $\rho\prec\alpha$ for every $\rho\in\lambda(u)$ and (ii) if $u^{\prime}\in Q$ is such that that $(u^{\prime},u,s_{k}(\alpha))\in E$ for some $1\leq k\leq|\alpha|-1$, then $u^{\prime}\in G^{\prec}(p_{|\alpha|-k}(\alpha))$.
Proof 4.5.
$(\Longrightarrow)$. Let us prove (i). Pick $\rho\in\lambda(u)$. Then, there exists $u^{\prime}\in Q$ such that $(u^{\prime},u,\rho)\in E$. We must prove that $\rho\prec\alpha$. Let $\beta^{\prime}\in I_{u^{\prime}}$; from $(u^{\prime},u,\rho)\in E$ we obtain $\beta^{\prime}\rho\in I_{u}$, so from $u\in G^{\prec}(\alpha)$ we obtain $\beta^{\prime}\rho\prec\alpha$, which implies $\rho\prec\alpha$. Let us prove (ii). Assume that $u^{\prime}\in Q$ is such that that $(u^{\prime},u,s_{k}(\alpha)\in E$ for some $1\leq k\leq|\alpha|-1$. We must prove that $u^{\prime}\in G^{\prec}(p_{|\alpha|-k}(\alpha))$. Pick $\gamma^{\prime}\in I_{u^{\prime}}$. We must prove that $\gamma^{\prime}\prec p_{|\alpha|-k}(\alpha)$. Since $(u^{\prime},u,s_{k}(\alpha))\in E$, we have $\gamma^{\prime}s_{k}(\alpha)\in I_{u}$, so from $u\in G^{\prec}(\alpha)$ we obtain $\gamma^{\prime}s_{k}(\alpha)\prec\alpha$, which implies $\gamma^{\prime}\prec p_{|\alpha|-k}(\alpha)$.
[…]
We now want to give a computational variant of Lemma 18. Let us give the following definition.
{defi}
Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler GDFA. Let $U\subseteq Q$ and $\rho\in\Sigma^{+}$. We denote by $\mathtt{out}(U,\rho)$ the number of edges labeled with $\rho$ that leave states in $U$, and we denote by $\mathtt{in}(U,\rho)$ the number of edges labeled with $\rho$ that reach states in $U$.
We are now ready to give a variant of Lemma 18. In particular, we will show that, if $\mathcal{A}$ is an $r$-GDFA, then in Lemma 18 we do not need to check each $1\leq k\leq|\alpha|-1$, but at most $r$ values of $k$.
Lemma 19.
Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler $r$-GDFA, and let $\alpha\in\Sigma^{*}$, with $\alpha\not=\epsilon$. For $1\leq i\leq\min\{r,|\alpha|-1\}$, let $f_{i}=\mathtt{out}(Q[1,|G^{\prec}(p_{|\alpha|-i}(\alpha))|],s_{i}(\alpha))$. Then, $|G^{\prec}(\alpha)|$ is the largest integer $0\leq j\leq|Q|$ such that (i) $\rho\prec\alpha$ for every $1\leq t\leq j$ and for every $\rho\in\lambda(Q[t])$, and (ii) $\mathtt{in}(Q[1,j],s_{i}(\alpha))\leq f_{i}$ for every $1\leq i\leq\min\{r,|\alpha|-1\}$.
Proof 4.6.
First, note that by Lemma 17 we have $G^{\prec}(\alpha)=Q[1,|G^{\prec}(\alpha)|]$ and, for every $1\leq i\leq\min\{r,|\alpha|-1\}$, we have $G^{\prec}(p_{|\alpha|-i}(\alpha))=Q[1,|G^{\prec}(p_{|\alpha|-i}(\alpha))|]$. Let $j^{*}$ be the largest integer $0\leq j\leq|Q|$ such that (i) $\rho\prec\alpha$ for every $1\leq t\leq j$ and for every $\rho\in\lambda(Q[t])$, and (ii) $\mathtt{in}(Q[1,j],s_{i}(\alpha))\leq f_{i}$ for every $1\leq i\leq\min\{r,|\alpha|-1\}$. We want to prove that $|G^{\prec}(\alpha)|=j^{*}$.
$(\leq)$ By the maximality of $j^{*}$, it will suffice to prove that (i) $\rho\prec\alpha$ for every $1\leq t\leq|G^{\prec}(\alpha)|$ and for every $\rho\in\lambda(Q[t])$, and (ii) $\mathtt{in}(Q[1,|G^{\prec}(\alpha)|],s_{i}(\alpha))\leq f_{i}$ for every $1\leq i\leq\min\{r,|\alpha|-1\}$.
Let us prove (i). Fix $1\leq t\leq|G^{\prec}(\alpha)|$ and $\rho\in\lambda(Q[t])$. We must prove that $\rho\prec\alpha$. From $1\leq t\leq|G^{\prec}(\alpha)|$ we obtain $Q[t]\in G^{\prec}(\alpha)$, so by Lemma 18 we conclude $\rho\prec\alpha$.
Let us prove (ii). Fix $1\leq i\leq\min\{r,|\alpha|-1\}$. By Lemma 18, we know that for every $(u^{\prime},u,s(\alpha,i))\in E$, if $u\in G^{\prec}(\alpha)=Q[1,|G^{\prec}(\alpha)|]$, then $u^{\prime}\in G^{\prec}(p_{|\alpha|-i}(\alpha))=Q[1,|G^{\prec}(p_{|\alpha|-i}(\alpha))|]$. The conclusion follows from the definition of $f_{i}$.
$(\geq)$ We only have to prove that if $|G^{\prec}(\alpha)|+1\leq j\leq|Q|$, then at least one of the following statements is not true:
  • (a) $\rho\prec\alpha$ for every $1\leq t\leq j$ and for every $\rho\in\lambda(Q[t])$.
  • (b) $\mathtt{in}(Q[1,j],s_{i}(\alpha))\leq f_{i}$ for every $1\leq i\leq\min\{r,|\alpha|-1\}$.
Fix $|G^{\prec}(\alpha)|+1\leq j\leq|Q|$, and assume that (a) is true. Then, we must prove that (b) is not true.
Since $|G^{\prec}(\alpha)|+1\leq j\leq|Q|$, we have $Q[j]\not\in G^{\prec}(\alpha)$. We also know that for every $\rho\in\lambda(Q[j])$ we have $\rho\prec\alpha$, so by Lemma 18 we conclude that for some $1\leq i\leq|\alpha|-1$ there exists $v^{\prime}\in Q$ such that $(v^{\prime},Q[j],s_{i}(\alpha))\in E$ and $v^{\prime}\not\in G^{\prec}(p_{|\alpha|-i}(\alpha))=Q[1,|G^{\prec}(p_{|\alpha|-i}(\alpha))|]$. Since $\mathcal{A}$ is an $r$-GDFA, we have $1\leq s_{i}(\alpha)\leq r$, so $1\leq i\leq\min\{r,|\alpha|-1\}$. Let us prove that $\mathtt{in}(Q[1,j],s_{i}(\alpha))>f_{i}$, which will imply that (b) is false. We know that $(v^{\prime},Q[j],s_{i}(\alpha))\in E$ and $v^{\prime}\not\in Q[1,|G^{\prec}(p_{|\alpha|-i}(\alpha))|]$, so by the definition of $f_{i}$ we only have to prove that, if $u^{\prime},u\in Q$ are such that $u^{\prime}\in Q[1,|G^{\prec}(p_{|\alpha|-i}(\alpha))|]$ and $(u^{\prime},u,s_{i}(\alpha))\in E$, then $u\in Q[1,j]$. Suppose for the sake of a contradiction that $u\not\in Q[1,j]$. This implies that $Q[j]\prec_{\mathcal{A}}u$, so from $(v^{\prime},Q[j],s_{i}(\alpha)),(u^{\prime},u,s_{i}(\alpha))\in E$ and Lemma 13 we obtain $v^{\prime}\prec_{\mathcal{A}}u^{\prime}$. Since $u^{\prime}\in Q[1,|G^{\prec}(p_{|\alpha|-i}(\alpha))|]$, we conclude $v^{\prime}\in Q[1,|G^{\prec}(p_{|\alpha|-i}(\alpha))|]$, a contradiction.
Let us show how to compute $G^{\prec}_{\dashv}(\alpha)$. To this end, let $G^{*}(\alpha)$ be the set of all states reached by an edge labeled with a string suffixed by $\alpha$. Formally:
$G^{*}(\alpha)=\{u\in Q\;|\;(\exists\rho\in\lambda(u))(\alpha\dashv\rho)\}.$
Notice that $G^{*}(\alpha)\subseteq G_{\dashv}(\alpha)$. Indeed, pick $u\in G^{*}(\alpha)$. Then, there exists $\rho\in\lambda(u)$ such that $\alpha\dashv\rho$. In particular, there exists $u^{\prime}\in Q$ such that $(u^{\prime},u,\rho)\in E$. Pick any $\beta\in I_{u^{\prime}}$. Then, $\beta\rho\in I_{u}$. From $\alpha\dashv\rho$ we conclude $\alpha\dashv\beta\rho$, so $u\in G_{\dashv}(\alpha)$.
The following crucial lemma shows that, to compute $|G^{\prec}_{\dashv}(\alpha)|$, we only have to consider $|G^{\prec}(\alpha)|$, the largest (w.r.t $\preceq_{\mathcal{A}}$) state in $G^{*}(\alpha)$ and the states in $G^{\prec}_{\dashv}(\alpha)\setminus G^{*}(\alpha)$.
Lemma 20.
Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler $r$-GDFA, and let $\alpha\in\Sigma^{*}$, with $\alpha\not=\epsilon$. For $1\leq i\leq\min\{r,|\alpha|-1\}$, let $f_{i}=\mathtt{out}(Q[1,|G^{\prec}(p_{|\alpha|-i}(\alpha))|],s_{i}(\alpha))$ and $g_{i}=\mathtt{out}(Q[1,|G^{\prec}_{\dashv}(p_{|\alpha|-i}(\alpha))|],s_{i}(\alpha))$. Then, $g_{i}\geq f_{i}$ for every $1\leq i\leq\min\{r,|\alpha|-1\}$. Moreover, $|G^{\prec}_{\dashv}(\alpha)|$ is equal to the maximum among:
  • $|G^{\prec}(\alpha)|$.
  • the largest integer $0\leq j\leq|Q|$ such that, if $j\geq 1$, then $Q[j]\in G^{*}(\alpha)$.
  • the smallest integer $0\leq j\leq|Q|$ such that, for every $1\leq i<\min\{r,|\alpha|-1\}$ for which $g_{i}>f_{i}$, we have $\mathtt{in}(Q[1,j],s_{i}(\alpha))\geq g_{i}$.
Proof 4.7.
First, note that by Lemma 17 we have $G^{\prec}(\alpha)=Q[1,|G^{\prec}(\alpha)|]$, $G^{\prec}_{\dashv}(\alpha)=Q[1,|G^{\prec}_{\dashv}(\alpha)|]$ and, for every $1\leq i\leq\min\{r,|\alpha|-1\}$, we have $G^{\prec}(p_{|\alpha|-i}(\alpha))=Q[1,|G^{\prec}(p_{|\alpha|-i}(\alpha))|]$ and $G^{\prec}_{\dashv}(p_{|\alpha|-i}(\alpha))=Q[1,|G^{\prec}_{\dashv}(p_{|\alpha|-i}(\alpha))|]$. For every $1\leq i\leq\min\{r,|\alpha|-1\}$, we have $g_{i}\geq f_{i}$ because $G^{\prec}(p_{|\alpha|-i}(\alpha))\subseteq G^{\prec}_{\dashv}(p_{|\alpha|-i}(\alpha))$. Let $j_{1}^{*}$ be the largest integer $0\leq j\leq|Q|$ such that, if $j\geq 1$, then $Q[j]\in G^{*}(\alpha)$, and let $j_{2}^{*}$ be the smallest integer $0\leq j\leq|Q|$ such that $\mathtt{in}(Q[1,j],s_{i}(\alpha))\geq g_{i}$ for every $1\leq i\leq\min\{r,|\alpha|-1\}$ for which $g_{i}>f_{i}$. Let $l=\max\{|G^{\prec}(\alpha)|,j_{1}^{*},j_{2}^{*}\}$. We want to prove that $|G^{\prec}_{\dashv}(\alpha)|=l$.
$(\geq)$ We have to prove that $|G^{\prec}(\alpha)|\leq|G^{\prec}_{\dashv}(\alpha)|$, $j_{1}^{*}\leq|G^{\prec}_{\dashv}(\alpha)|$ and $j_{2}^{*}\leq|G^{\prec}_{\dashv}(\alpha)|$. The inequality $|G^{\prec}(\alpha)|\leq|G^{\prec}_{\dashv}(\alpha)|$ follows from $G^{\prec}(\alpha)\subseteq G^{\prec}_{\dashv}(\alpha)$. Let us prove that $j_{1}^{*}\leq|G^{\prec}_{\dashv}(\alpha)|$. If $j_{1}^{*}=0$, we are done, so we can assume $j_{1}^{*}\geq 1$. We know that $Q[j_{1}^{*}]\in G^{*}(\alpha)$, so $Q[j_{1}^{*}]\in G_{\dashv}(\alpha)$ and $Q[j_{1}^{*}]\in G^{\prec}_{\dashv}(\alpha)$, which implies $j_{1}^{*}\leq|G^{\prec}_{\dashv}(\alpha)|$.
Let us prove that $j_{2}^{*}\leq|G^{\prec}_{\dashv}(\alpha)|$. If $j_{2}^{*}=0$, we are done, so we can assume $j_{2}^{*}\geq 1$. We know that there exists $1\leq i\leq\min\{r,|\alpha|-1\}$ such that $g_{i}>f_{i}$, $\mathtt{in}(Q[1,j_{2}^{*}-1],s_{i}(\alpha))<g_{i}$ and $\mathtt{in}(Q[1,j_{2}^{*}],s_{i}(\alpha))\geq g_{i}$. We are only left with showing that there exists $u^{\prime}\in G_{\dashv}(p_{|\alpha|-i}(\alpha))$ such that $(u^{\prime},Q[j_{2}^{*}],s_{i}(\alpha))\in E$, because this will imply $Q[j_{2}^{*}]\in G_{\dashv}(\alpha)$, so $Q[j_{2}^{*}]\in G^{\prec}_{\dashv}(\alpha)$ and $j_{2}^{*}\leq|G^{\prec}_{\dashv}(\alpha)|$. Suppose for the sake of a contradiction that for every $u^{\prime}\in Q$ such that $(u^{\prime},Q[j_{2}^{*}],s_{i}(\alpha))\in E$ we have $u^{\prime}\not\in G_{\dashv}(p_{|\alpha|-i}(\alpha))$.
First, let us prove that there cannot exist two states $u^{\prime}_{1}\in Q[1,|G^{\prec}(p_{|\alpha|-i}(\alpha))|]$ and $u^{\prime}_{2}\in Q[|G^{\prec}_{\dashv}(p_{|\alpha|-i}(\alpha))|+1,|Q|]$ such that $(u^{\prime}_{1},Q[j_{2}^{*}],s_{i}(\alpha))\in E$ and $(u^{\prime}_{2},Q[j_{2}^{*}],s_{i}(\alpha))\in E$. Suppose for the sake of a contradiction that such $u^{\prime}_{1}$ and $u^{\prime}_{2}$ exist. We will obtain a contradiction by showing that the existence of $u^{\prime}_{1}$ and $u^{\prime}_{2}$ implies that there cannot exist $u^{\prime}_{3}\in G_{\dashv}(p_{|\alpha|-i}(\alpha))$ and $u_{3}\in Q$ such that $(u^{\prime}_{3},u_{3},s_{i}(\alpha))\in E$, because this would imply $g_{i}=f_{i}$, which contradicts $g_{i}>f_{i}$.
Suppose for the sake of a contradiction that there exist $u^{\prime}_{3}\in G_{\dashv}(p_{|\alpha|-i}(\alpha))$ and $u_{3}\in Q$ such that $(u^{\prime}_{3},u_{3},s_{i}(\alpha))\in E$. We obtain a contradiction by distinguishing three cases:
  • we cannot have $u_{3}=Q[j_{2}^{*}]$ because for every $u^{\prime}\in Q$ such that $(u^{\prime},Q[j_{2}^{*}],s_{i}(\alpha))\in E$ we have $u^{\prime}\not\in G_{\dashv}(p_{|\alpha|-i}(\alpha))$.
  • we cannot have $u_{3}\prec_{\mathcal{A}}Q[j_{2}^{*}]$ because, if we consider $(u^{\prime}_{3},u_{3},s_{i}(\alpha)),(u^{\prime}_{1},Q[j_{2}^{*}],s_{i}(\alpha))\in E$, by Lemma 13 we would obtain $u^{\prime}_{3}\prec_{\mathcal{A}}u^{\prime}_{1}$, and since $u^{\prime}_{1}\in Q[1,|G^{\prec}(p_{|\alpha|-i}(\alpha))|]$, we would conclude $u^{\prime}_{3}\in Q[1,|G^{\prec}(p_{|\alpha|-i}(\alpha))|]$, which contradicts $u^{\prime}_{3}\in G_{\dashv}(p_{|\alpha|-i}(\alpha))$.
  • we cannot have $Q[j_{2}^{*}]\prec_{\mathcal{A}}u_{3}$ because, if we consider $(u^{\prime}_{2},Q[j_{2}^{*}],s_{i}(\alpha)),(u^{\prime}_{3},u_{3},s_{i}(\alpha))\in E$, by Lemma 13 we would obtain $u^{\prime}_{2}\prec_{\mathcal{A}}u^{\prime}_{3}$, and since $u^{\prime}_{2}\in Q[|G^{\prec}_{\dashv}(p_{|\alpha|-i}(\alpha))|+1,|Q|]$, we would conclude $u^{\prime}_{3}\in Q[|G^{\prec}_{\dashv}(p_{|\alpha|-i}(\alpha))|+1,|Q|]$, which contradicts $u^{\prime}_{3}\in G_{\dashv}(p_{|\alpha|-i}(\alpha))$.
We have proved that there cannot exist two states $u^{\prime}_{1}\in Q[1,|G^{\prec}(p_{|\alpha|-i}(\alpha))|]$ and $u^{\prime}_{2}\in Q[|G^{\prec}_{\dashv}(p_{|\alpha|-i}(\alpha))|+1,|Q|]$ such that $(u^{\prime}_{1},Q[j_{2}^{*}],s_{i}(\alpha))\in E$ and $(u^{\prime}_{2},Q[j_{2}^{*}],s_{i}(\alpha))\in E$. Moreover, we know that for every $u^{\prime}\in Q$ such that $(u^{\prime},Q[j_{2}^{*}],s_{i}(\alpha))\in E$ we have $u^{\prime}\not\in G_{\dashv}(p_{|\alpha|-i}(\alpha))$. As a consequence, one of the following two cases must occur.
  • For every $u^{\prime}_{1}\in Q$ such that $(u^{\prime}_{1},Q[j_{2}^{*}],s_{i}(\alpha))\in E$ we have $u^{\prime}_{1}\in Q[1,|G^{\prec}(p_{|\alpha|-i}(\alpha))|]$. Let us prove that if $v^{\prime},v\in Q$ are such that $v\in Q[1,j_{2}^{*}]$ and $(v^{\prime},v,s_{i}(\alpha))\in E$, then $v^{\prime}\in Q[1,|G^{\prec}(p_{|\alpha|-i}(\alpha))|]$. This will imply $\mathtt{in}(Q[1,j_{2}^{*}],s_{i}(\alpha))\leq f_{i}<g_{i}$, the desired contradiction. If $v=Q[j_{2}^{*}]$, the conclusion follows because we know that for every $u^{\prime}_{1}\in Q$ such that $(u^{\prime}_{1},Q[j_{2}^{*}],s_{i}(\alpha))\in E$ we have $u^{\prime}_{1}\in Q[1,|G^{\prec}(p_{|\alpha|-i}(\alpha))|]$. Now, assume that $v\prec_{\mathcal{A}}Q[j_{2}^{*}]$. Since $\mathtt{in}(Q[1,j_{2}^{*}-1],s_{i}(\alpha))<g_{i}$ but $\mathtt{in}(Q[1,j_{2}^{*}],s_{i}(\alpha))\geq g_{i}$, then there exists $u^{*}_{1}\in Q$ such that $(u^{*}_{1},Q[j_{2}^{*}],s_{i}(\alpha))\in E$, and we know that it must be $u^{*}_{1}\in Q[1,|G^{\prec}(p_{|\alpha|-i}(\alpha))|]$. If we consider $(v^{\prime},v,s_{i}(\alpha)),(u^{*}_{1},Q[j_{2}^{*}],s_{i}(\alpha))\in E$, from $v\prec_{\mathcal{A}}Q[j_{2}^{*}]$ and Lemma 13 we obtain $v^{\prime}\prec_{\mathcal{A}}u^{*}_{1}$, and since $u^{*}_{1}\in Q[1,|G^{\prec}(p_{|\alpha|-i}(\alpha))|]$, we conclude $v^{\prime}\in Q[1,|G^{\prec}(p_{|\alpha|-i}(\alpha))|]$.
  • For every $u^{\prime}_{2}\in Q$ such that $(u^{\prime}_{2},Q[j_{2}^{*}],s_{i}(\alpha))\in E$ we have $u^{\prime}_{2}\in Q[|G^{\prec}_{\dashv}(p_{|\alpha|-i}(\alpha))|+1,|Q|]$. Let us prove that if $v^{\prime},v\in Q$ are such that $v^{\prime}\in Q[1,|G^{\prec}_{\dashv}((p_{|\alpha|-i}(\alpha))|]$ and $(v^{\prime},v,s_{i}(\alpha))\in E$, then $v\in Q[1,j_{2}^{*}-1]$. This will imply $\mathtt{in}(Q[1,j_{2}^{*}-1],s_{i}(\alpha))\geq g_{k}$, the desired contradiction. It cannot be $v=Q[j_{2}^{*}]$, because this would imply $v^{\prime}\in Q[|G^{\prec}_{\dashv}(p_{|\alpha|-i}(\alpha))|+1,|Q|]$. Now, assume for the sake of contradiction that $Q[j_{2}^{*}]\prec_{\mathcal{A}}v$. Since $\mathtt{in}(Q[1,j_{2}^{*}-1],s_{i}(\alpha))<g_{i}$ but $\mathtt{in}(Q[1,j_{2}^{*}],s_{i}(\alpha))\geq g_{i}$, then there exists $u^{*}_{2}\in Q$ such that $(u^{*}_{2},Q[j_{2}^{*}],s_{i}(\alpha))\in E$, and we know that it must be $u^{*}_{2}\in Q[|G^{\prec}_{\dashv}(p_{|\alpha|-i}(\alpha))|+1,|Q|]$. If we consider $(u^{*}_{2},Q[j_{2}^{*}],s_{i}(\alpha)),(v^{\prime},v,s_{i}(\alpha))\in E$, from $Q[j_{2}^{*}]\prec_{\mathcal{A}}v$ and Lemma 13 we obtain $u^{*}_{2}\prec_{\mathcal{A}}v^{\prime}$, and since $u^{*}_{2}\in Q[|G^{\prec}_{\dashv}(p_{|\alpha|-i}(\alpha))|+1,|Q|]$ , we conclude $v^{\prime}\in Q[|G^{\prec}_{\dashv}(p_{|\alpha|-i}(\alpha))|+1,|Q|]$, a contradiction.
$(\leq)$ We only have to prove that if $l+1\leq j\leq|Q|$, then $Q[j]\not\in G^{\prec}_{\dashv}(\alpha)$. Since $j\geq l+1>|G^{\prec}(\alpha)|$, we have $Q[j]\not\in G^{\prec}(\alpha)$, hence we are left with showing that $Q[j]\not\in G_{\dashv}(\alpha)$. Assume for the sake of a contradiction that $Q[j]\in G_{\dashv}(\alpha)$. It cannot be $Q[j]\in G^{*}(\alpha)$, otherwise $j\leq j_{1}^{*}\leq l$, a contradiction. Hence, $Q[j]\in G_{\dashv}(\alpha)\setminus G^{*}(\alpha)$. Since $\mathcal{A}$ is an $r$-GDFA, this means that for some $1\leq i\leq\min\{r,|\alpha|-1\}$ there exists $u^{\prime}\in Q$ such that $(u^{\prime},Q[j],s_{i}(\alpha))\in E$ and $u^{\prime}\in G_{\dashv}(p_{|\alpha|-i}(\alpha))$, which implies $g_{i}>f_{i}$. Let us prove that $\mathtt{in}(Q[1,j-1],s_{i}(\alpha))<g_{i}$. Assume that $(v^{\prime},v,s_{i}(\alpha))\in E$ is such that $v\in Q[1,j-1]$. If we consider $(v^{\prime},v,s_{i}(\alpha)),(u^{\prime},Q[j],s_{i}(\alpha))\in E$, from $v\prec_{\mathcal{A}}Q[j]$ and Lemma 13 we obtain $v^{\prime}\prec_{\mathcal{A}}u^{\prime}$, and since $u^{\prime}\in G_{\dashv}(p_{|\alpha|-i}(\alpha))$, and so $u^{\prime}\in Q[1,|G^{\prec}_{\dashv}(p_{|\alpha|-i}(\alpha))|]$, we obtain $v^{\prime}\in Q[1,|G^{\prec}_{\dashv}(p_{|\alpha|-i}(\alpha))|]$. This implies that $\mathtt{in}(Q[1,j-1],s_{i}(\alpha))\leq g_{i}$, and it must be $\mathtt{in}(Q[1,j-1],s_{i}(\alpha))<g_{i}$ because $(u^{\prime},Q[j],s_{i}(\alpha))\in E$ and $u^{\prime}\in G_{\dashv}(p_{|\alpha|-i}(\alpha))$. We conclude $j\leq j_{2}^{*}\leq l$, a contradiction.
Remark 21.
Consider the statement of Lemma 20. If $|\alpha|>r$, then the largest integer $0\leq j\leq|Q|$ such that, if $j\geq 1$, then $Q[j]\in G^{*}(\alpha)$ is equal to 0.
Let us present some helpful results in the realm of compressed data structures. We start with a classical lemma: we can store a string of length $t$ over an alphabet of size $\sigma$ using only slightly more than $t\log\sigma$ bits in such a way that we can solve the crucial rank and select operations efficiently.
Lemma 22 ([Nav16], Chapter 6).
Let $\Sigma=\{0,1,\dots,\sigma-1\}$ be an integer alphabet and let $\alpha\in\Sigma^{*}$, with $|\alpha|=t$. If $\sigma\leq t$, then $\alpha$ can be encoded using a data structure of $t\log\sigma(1+o(1))+O(t)$ bits that supports the following operations in $O(\log\log\sigma)$ time:
  • $\alpha.access(i)$, for $1\leq i\leq t$: return $\alpha[i]$.
  • $\alpha.rank(i,c)$, for $c\in\Sigma$ and $0\leq i\leq t$: return $|\{1\leq j\leq i\ |\ \alpha[j]=c\}|$.
  • $\alpha.select(i,c)$, for $c\in\Sigma$ and $1\leq i\leq\alpha.rank(t,c)$: return the unique integer $1\leq j\leq t$ such that (i) $\alpha[j]=c$ and (ii) $\alpha.rank(j,c)=i$.
For example, if $\alpha=abaaaabaab$, then we have $\alpha.access(2)=b$, $\alpha.rank(5,a)=4$ and $\alpha.select(4,a)=5$. Rank and select are closely related: for every $c\in\Sigma$ and for every $1\leq i\leq\alpha.rank(t,c)$, we have $\alpha.rank(\alpha.select(i,c),c))=i$. Note that not only is the data structure of Lemma 22 an encoding of $\alpha$ (that is, from the data structure we can retrieve $\alpha$), but we can also retrieve $\alpha$ quickly through access operations.
Remark 23.
To handle some boundary cases easily, it is expedient to introduce some extensions of rank and select:
  • $\alpha.rank(i,c)$, for $c\in\Sigma$ and $i\geq 0$: return $|\{1\leq j\leq\min\{i,t\}\ |\ \alpha[j]=c\}|$.
  • $\alpha.select(i,c)$, for $c\in\Sigma$ and $i\geq 1$: if $i\leq\alpha.rank(t,c)$, return the unique integer $1\leq j\leq t$ such that (i) $\alpha[j]=c$ and (ii) $\alpha.rank(j,c)=i$, and if $i>\alpha.rank(t,c)$, return $t+1$.
The data structure of Lemma 22 can compute these extensions in $O(\log\log\sigma)$ time. Indeed, to solve the extended rank, we only need to check whether $i\leq t$, and to solve the extended select, we only need to check (in $O(\log\log\sigma)$ time) whether $i\leq\alpha.rank(t,c)$.
The next lemma shows that we can store a dictionary on $\Sigma$ (that is, a subset of $\Sigma$) within compressed space in such a way that we can solve a variant of rank and select efficiently. Dictionaries are needed for two reasons: (i) the alphabet $\Sigma$ need not be effective, that is, some characters in $\Sigma$ may label no edge of $\mathcal{A}$ and, most importantly, (ii) even when $\Sigma$ is effective, in general only some strings in $\Sigma^{i}$ label some edge (that is, $\Sigma_{i}$ is strictly contained in $\Sigma^{i}$), so will need a dictionary for each $\Sigma^{i}$.
Lemma 24 ([FPS16], Theorem 4.1).
Let $\Sigma=\{0,1,\dots,\sigma-1\}$ be an integer alphabet, and let $A\subseteq\Sigma$, with $|A|=t$. Then, $A$ can be encoded using a data structures of $t\log(\sigma/t)+O(t)$ bits that supports the following operations in $O(\log\log(\sigma/t))$ time:
  • $A.rank(i)$, for $0\leq i\leq\sigma-1$: return $|\{j\in A\ |\ j\leq i\}|$.
  • $A.select(i)$, for $1\leq i\leq t$: return the unique integer $0\leq j\leq\sigma-1$ such that (i) $j\in A$ and (ii) $A.rank(j)=i$.
Remark 25.
The data structure of Lemma 24 also supports the following operations in $O(\log\log(\sigma/t))$ time:
  • $A.memb(i)$, for $0\leq i\leq\sigma-1$: decide whether $i\in A$ (membership).
  • $A.prec(i)$, for $0\leq i\leq\sigma-1$: return the largest integer $0\leq j<i$ such that $j\in A$, if such a $j$ exists, otherwise return $\bot$ (predecessor).
  • $A.succ(i)$, for $0\leq i\leq\sigma-1$: return the smallest integer $i<j\leq\sigma-1$ such that $j\in A$, if such a $j$ exists, otherwise return $\bot$ (successor).
Indeed, we can compute these operations as follows.
  • Let us show how to compute $A.memb(i)$ in $O(\log\log(\sigma/t))$ time. We have $i\in A$ if and only if $A.select(A.rank(i))=i$.
  • Let us show how to compute $A.prec(i)$ in $O(\log\log(\sigma/t))$ time. If $i=0$, return $\bot$. Now assume that $i\geq 1$. Compute A.rank(i1)A.rank(i-1). If A.rank(i1)=0A.rank(i-1)=0, return $\bot$, and if A.rank(i1)>0A.rank(i-1)>0, return A.select(A.rank(i1))A.select(A.rank(i-1)).
  • Let us show how to compute $A.succ(i)$ in $O(\log\log(\sigma/t))$ time. Compute $A.rank(i)$ and $A.rank(\sigma-1)$. If $A.rank(i)=A.rank(\sigma-1)$, return $\bot$, and if $A.rank(i)<A.rank(\sigma-1)$, return $A.select(A.rank(i)+1)$.
Note that the membership operation confirms that the data structure of Lemma 24 is an encoding of $A$.
We are ready to extend the FM-index to Wheeler GDFAs. Recall that $e$ is the number of edges, $\mathfrak{e}$ is the total length of all edge labels and $\sigma=|\Sigma|$.
Theorem 26 (FM-index of Wheeler GDFAs).
Let $\mathcal{A}=(Q,E,s,F)$ be a Wheeler $r$-GDFA, with $\sigma\leq\mathfrak{e}^{O(1)}$ and $r=O(1)$. Then, we can encode $\mathcal{A}$ by using $\mathfrak{e}\log\sigma(1+o(1))+O(e)$ bits so that later on, given a pattern $\alpha\in\Sigma^{*}$ of length $m$, we can solve the SMLG problem on $\mathcal{A}$ in $O(m\log\log\sigma)$ time. Within the same time bound, we can also decide whether $\alpha$ is recognized by $\mathcal{A}$.
Proof 4.8.
Note that, since $\sigma\leq\mathfrak{e}^{O(1)}$ and $r=O(1)$, for every $1\leq i\leq r$ the elements of $\Sigma^{i}$ fit in a constant number of computer words and thus can be manipulated in constant time. Let $n=|Q|$. Recall that $e\geq n-1$ because every state is reachable from the initial state, so every state different from the initial state must have an incoming edge. Moreover, recall that $\Sigma=\{0,1,\dots,\sigma-1\}$, with $1\prec 2\prec\dots\prec\sigma-1$.
Let us describe our data structures. Here is an intuition. First, for every $1\leq i\leq r$, we will map each string in $\Sigma_{i}$ to an an integer, obtaining the set $\Sigma_{i}^{*}$. In this way, we can manipulate each $\Sigma_{i}^{*}$ through a dictionary (Lemma 24), which will be expedient to solve the SMLG problem quickly. Then, we store each component of the Burrows-Wheeler transform of $\mathcal{A}$ (Definition 4.3) through the data structure of Lemma 22, with two caveats: (i) we store a variant $\mathtt{LAB}^{*}_{i}$ of each $\mathtt{LAB}_{i}$, and (ii) we store some auxiliary strings $\mathtt{AUX}^{*}_{i}$’s. We store the variant $\mathtt{LAB}^{*}_{i}$ to reduce the space of our data structures: since we can quickly check if a string is in $\Sigma_{i}^{*}$ (and so in $\Sigma_{i}$) through a dictionary, we can map each element in $\Sigma_{i}$ to an element in $\{0,1,\dots,\sigma_{i}-1\}$, which saves space (because each element in $\Sigma_{i}$ is a string of length $i$ on $\Sigma$, and $\sigma_{i}$ may be much smaller than $\sigma^{i}$). The auxiliary strings $\mathtt{AUX}^{*}_{i}$’s only require $O(e)$ bits but are helpful to solve the SMLG problem quickly. We will now give a formal description of our data structures.
For every $1\leq i\leq r$, let $\psi_{i}$ be the bijection from $\Sigma^{i}$ to $\{0,1,\dots,\sigma^{i}-1\}$ that maps every element $a_{1}a_{2}\dots a_{i-1}a_{i}$ in $\Sigma^{i}$ to the representation in base $\sigma$ of the reverse string $a_{i}a_{i-1}\dots a_{2}a_{1}$. In other words, we have $\psi_{i}(a_{1}a_{2}\dots a_{i-1}a_{i})=a_{1}+a_{2}\sigma+a_{3}\sigma^{2}+\dots a_{i-1}\sigma^{i-2}+a_{i}\sigma^{i-1}$. In particular, for every $\rho\in\Sigma_{i}$ we have $\psi_{i}(\rho)\in\{0,1,\dots,\sigma^{i}-1\}$. Note that $\psi_{i}$ is monotone: for every $\rho,\rho^{\prime}\in\Sigma^{i}$ we have $\rho\prec\rho^{\prime}$ if and only if $\psi(\rho)<\psi(\rho^{\prime})$. For every $1\leq i\leq r$, define $\Sigma^{*}_{i}=\{\psi_{i}(\rho)\;|\;\rho\in\Sigma_{i}\}$. Note that $|\Sigma^{*}_{i}|=|\Sigma_{i}|=\sigma_{i}$ for every $1\leq i\leq r$.
We store the following data structures.
  • For every $1\leq i\leq r$, the data structure of Lemma 24 for the set $\Sigma^{*}_{i}\subseteq\{0,1,\dots,\sigma^{i}-1\}$. We know that $|\Sigma^{*}_{i}|=\sigma_{i}$, so the total number of required bits is $\sum_{i=1}^{r}(\sigma_{i}\log(\sigma^{i}/\sigma_{i})+O(\sigma_{i}))\leq\sum_{i=1}^{r}(e_{i}\log(\sigma^{i}/\sigma_{i})+O(e_{i}))=\sum_{i=1}^{r}(e_{i}\log\sigma^{i})-\sum_{i=1}^{r}(e_{i}\log\sigma_{i})+\sum_{i=1}^{r}O(e_{i})=(\sum_{i=1}^{r}e_{i}i)\log\sigma-(\sum_{i=1}^{r}e_{i}\log\sigma_{i})+O(e)=\mathfrak{e}\log\sigma-(\sum_{i=1}^{r}e_{i}\log\sigma_{i})+O(e)$. We can solve rank and select queries (and membership, predecessor and successor queries, see Remark 25) on each $\Sigma^{*}_{i}$ in $O(\log\log(\sigma^{i}/\sigma_{i}))\subseteq O(\log\log\sigma^{r})\subseteq(\log r+\log\log\sigma)\subseteq O(\log\log\sigma)$ time.
  • For every $1\leq i\leq r$, the data structure of Lemma 22 for the string $\mathtt{OUT}_{i}\in\{0,1\}^{e_{i}+n}$ of Definition 4.3. The total number of required bits is $\sum_{i=1}^{r}((e_{i}+n)(1+o(1))+O(e_{i}+n))\subseteq\sum_{i=1}^{r}O(e_{i}+n)=O(e+nr)\subseteq O(e)$. We can solve access, rank and select queries on each $\mathtt{OUT}_{i}$ in $O(1)$ time.
  • For every $1\leq i\leq r$, the data structure of Lemma 22 for the string $\mathtt{IN}_{i}\in\{0,1\}^{e_{i}+n}$ of Definition 4.3. The total number of required bits is again $O(e)$. We can solve access, rank and select queries on each $\mathtt{IN}_{i}$ in $O(1)$ time.
  • For every $1\leq i\leq r$, the data structure of Lemma 22 for the string $\mathtt{LAB}^{*}_{i}\in(\{0,1,\dots,\sigma_{i}-1\})^{e_{i}}$ defined as follows. Consider the string $\mathtt{LAB}_{i}\in(\Sigma_{i})^{e_{i}}$ of Definition 4.3, and let $\mathtt{LAB}^{*}_{i}$ be the string of length $e_{i}$ such that $\mathtt{LAB}^{*}_{i}[j]=\Sigma_{i}^{*}.rank(\psi_{i}(\mathtt{LAB}_{i}[j]))-1$ for every $1\leq j\leq e_{i}$. In other words, (i) we compute $\psi_{i}(\mathtt{LAB}_{i}[j])\in\{0,1,\dots,\sigma^{i}-1\}$ and then (ii) we compute the position of $\psi_{i}(\mathtt{LAB}_{i}[j])$ in the sorted list of all elements in $\Sigma^{*}_{i}$. Note that $\psi_{i}(\mathtt{LAB}_{i}[j])\in\Sigma^{*}_{i}$ because $\mathtt{LAB}_{i}[j]\in\Sigma_{i}$, hence $1\leq\Sigma_{i}^{*}.rank(\psi_{i}(\mathtt{LAB}_{i}[j]))\leq\sigma_{i}$, which implies $\mathtt{LAB}^{*}_{i}\in(\{0,1,\dots,\sigma_{i}-1\})^{e_{i}}$. Moreover, $\sigma_{i}\leq e_{i}$, so the assumption required in Lemma 22 is satisfied. Notice that for every $1\leq j,j^{\prime}\leq e_{i}$ we have $\mathtt{LAB}_{i}^{*}[j]=\mathtt{LAB}_{i}^{*}[j^{\prime}]$ if and only if $\mathtt{LAB}_{i}[j]=\mathtt{LAB}_{i}[j^{\prime}]$ (because $\psi_{i}$ is a bijection), and we have $\mathtt{LAB}_{i}^{*}[j]<\mathtt{LAB}_{i}^{*}[j^{\prime}]$ if and only if $\mathtt{LAB}_{i}[j]\prec\mathtt{LAB}_{i}[j^{\prime}]$ (because $\psi_{i}$ is monotone). The total number of required bits is $\sum_{i=1}^{r}(e_{i}\log\sigma_{i}(1+o(1))+O(e_{i}))\leq(\sum_{i=1}^{r}e_{i}\log\sigma_{i})+(\sum_{i=1}^{r}e_{i}\log\sigma^{i})\cdot o(1)+O(e)=(\sum_{i=1}^{r}e_{i}\log\sigma_{i})+(\sum_{i=1}^{r}e_{i}i)\log\sigma\cdot o(1)+O(e)=(\sum_{i=1}^{r}e_{i}\log\sigma_{i})+\mathfrak{e}\log\sigma\cdot o(1)+O(e)$. We can solve access, rank and select queries on each $\mathtt{LAB}^{*}_{i}$ in $O(\log\log\sigma_{i})\subseteq O(\log\log\sigma^{i})\subseteq O(\log\log\sigma^{r})\subseteq O(\log r+\log\log\sigma)=O(\log\log\sigma)$ time.
  • For every $1\leq i\leq r$, the data structure of Lemma 22 for the (auxiliary) string $\mathtt{AUX}^{*}_{i}\in\{0,1\}^{e_{i}}$ defined as follows. Sort all edges in $E_{i}$ by the index of the end states (w.r.t to $\preceq_{\mathcal{A}}$). Edges with the same end state are sorted by label. Edges with the same end state and the same label are sorted by the index of the start states (w.r.t to $\preceq_{\mathcal{A}}$). Then, we obtain $\mathtt{AUX}_{i}\in(\Sigma_{i})^{e_{i}}$ by concatenating the labels of all edges following this edge order. Note that by Lemma 13, if $1\leq k\leq e_{i}-1$, then $\mathtt{AUX}_{i}[k]\preceq\mathtt{AUX}_{i}[k+1]$ (all edge labels in $E_{i}$ have length $i$, so no edge label is a strict suffix of some other edge label). The string $\mathtt{AUX}^{*}_{i}\in\{0,1\}^{e_{i}}$ is the string on $\{0,1\}$ such that, for every $1\leq k\leq e_{i}$, we have $\mathtt{AUX}^{*}[k]=1$ if and only if $k=1$ or $(k\geq 2)\land(\mathtt{AUX}_{i}[k]\neq\mathtt{AUX}_{i}[k-1])$. The total number of required bits is $\sum_{i=1}^{r}(e_{i}(1+o(1))+O(e_{i}))\subseteq\sum_{i=1}^{r}O(e_{i})=O(e)$. We can solve access, rank and select queries on each $\mathtt{AUX}^{*}_{i}$ in $O(1)$ time.
  • The data structure of Lemma 22 for the string $\texttt{FIN}\in\{0,1\}^{n}$ of Definition 4.3. The number of required bits is $n(1+o(1))+O(n)\subseteq O(n)\subseteq O(e)$. We can solve access, rank and select queries on $\mathtt{FIN}$ in $O(1)$ time.
By adding up the space required of all data structures, we conclude that the total space is $\mathfrak{e}\log\sigma(1+o(1))+O(e)$ bits.
Let us show that our data structures are an encoding of $\mathcal{A}$. By Theorem 15, we only need to show that our data structures are an encoding of $\operatorname{BWT}(\mathcal{A})$. By Definition 4.3, we need to show that our data structures are an encoding of (i) $\mathtt{OUT}_{i}$, $\mathtt{IN}_{i}$ and $\mathtt{LAB}_{i}$, for every $1\leq i\leq r$, and (ii) $\mathtt{FIN}$. By Lemma 22 and Lemma 24, we know that our data structures are an encoding of (i) $\Sigma_{i}^{*}$, $\mathtt{OUT}_{i}$, $\mathtt{IN}_{i}$, $\mathtt{LAB}^{*}_{i}$ and $\mathtt{AUX}^{*}_{i}$, for every $1\leq i\leq r$, and (ii) $\mathtt{FIN}$. The conclusion will follow if we show that, for every $1\leq i\leq r$, $\Sigma_{i}^{*}$ and $\mathtt{LAB}^{*}_{i}$ are an encoding of $\mathtt{LAB}_{i}$. Fix $1\leq i\leq r$ and $1\leq j\leq e_{i}$; it will suffice to show that we can retrieve $\mathtt{LAB}_{i}[j]$ from $\Sigma_{i}^{*}$ and $\mathtt{LAB}^{*}_{i}[j]$. Notice that $\psi_{i}(\mathtt{LAB}_{i}[j])=\Sigma_{i}^{*}.select(\mathtt{LAB}^{*}_{i}[j]+1)$, so we can retrieve $\mathtt{LAB}_{i}[j]$ because $\psi_{i}$ is a bijection from $\Sigma^{i}$ to $\{0,1,\dots,\sigma^{i}-1\}$.
Let us show how to solve the SMLG problem. First, let us compute an auxiliary integer $\mathfrak{q}$. Recall that 0 is the smallest character in $\Sigma$ (w.r.t $\preceq$). Let $\mathfrak{q}$ be the largest integer $0\leq k\leq m$ such that $p_{k}(\alpha)=00\dots 0=0^{k}$. We can compute $\mathfrak{q}$ in $O(m)$ time by scanning $\alpha$ from left to right.
Consider a string $\rho\in\Sigma^{*}$. We define suffix-mapping $\rho$ as the process of computing $\psi_{i}(s_{i}(\rho))$ for every $1\leq i\leq r$. For example, if $\sigma=4$, $r=3$ and $\rho=1321$, then suffix-mapping $\rho$ means computing $\psi_{1}(1)=1$, $\psi_{2}(21)=6$ and $\psi_{3}(321)=27$. We can suffix-map $\rho$ in $O(r)\subseteq O(1)$ time because (i) $\psi(s_{1}(\rho))=s_{1}(\rho)$ and (ii) for $2\leq i\leq r$ we have $\psi_{i}(s_{i}(\rho))=\rho[|\rho|-i+1]+\psi_{i-1}(s_{i-1}(\rho))\cdot\sigma$. Indeed, we have $\psi_{i}(s_{i}(\rho))=\rho[|\rho|-i+1]+\rho[|\rho|-i+2]\sigma+\rho[|\rho|-i+3]\sigma^{2}+\dots+\rho[|\rho|]\sigma^{i-1}=\rho[|\rho|-i+1]+(\rho[|\rho|-i+2]+\rho[|\rho|-i+3]\sigma+\dots+\rho[|\rho|]\sigma^{i-2})\cdot\sigma=\rho[|\rho|-i+1]+\psi_{i-1}(s_{i-1}(\rho))\cdot\sigma$.
By Lemma 17, to solve the SMLG problem, we only need to compute $|G^{\prec}(\alpha)|$ and $|G^{\prec}_{\dashv}(\alpha)|$. To this end, in $m$ steps, we will recursively compute $|G^{\prec}(p_{k}(\alpha))|$ and $|G^{\prec}_{\dashv}(p_{k}(\alpha))|$ for every $0\leq k\leq m$, and we will obtain the conclusion by picking $k=m$. Note that the case $k=0$ is immediate because $p_{0}(\alpha)=\epsilon$, and $|G^{\prec}(\epsilon)|=0$ and $|G^{\prec}_{\dashv}(\epsilon)|=n$ (see Remark 16). To obtain the time bound $O(m\log\log\sigma)$, we only need to show the following. Fix $1\leq k\leq m$, and assume that we know $|G^{\prec}(p_{i}(\alpha))|$ and $|G^{\prec}_{\dashv}(p_{i}(\alpha))|$ for every $0\leq i\leq k-1$. Then, we must prove that in $O(\log\log\sigma)$ time we can compute $|G^{\prec}(p_{k}(\alpha))|$ and $|G^{\prec}_{\dashv}(p_{k}(\alpha))|$.
Let us define four operations.
  • $\mathcal{A}.op_{1}(i,x,j)$, for $1\leq i\leq r$, $0\leq x\leq\sigma^{i}-1$ and $1\leq j\leq n$: return $\mathtt{out}(Q[1,j],\psi_{i}^{-1}(x))$.
  • $\mathcal{A}.op_{2}(i,x,h)$, for $1\leq i\leq r$, $0\leq x\leq\sigma^{i}-1$ and $h\geq 0$: return the largest $0\leq j\leq n$ such that $\mathtt{in}(Q[1,j],\psi_{i}^{-1}(x))\leq h$.
  • $\mathcal{A}.op_{3}(i,x,h)$, for $1\leq i\leq r$, $0\leq x\leq\sigma^{i}-1$ and $h\geq 1$: return the smallest $0\leq j\leq n$ such that $\mathtt{in}(Q[1,j],\psi_{i}^{-1}(x))\geq h$, or report that such a $j$ does not exist.
  • $\mathcal{A}.op_{4}(i,x)$, for $1\leq i\leq r$ and $0\leq x\leq\sigma^{i}-1$: return the largest integer $0\leq j\leq n$ such that, for every $1\leq t\leq j$ and for every $\rho\in\lambda(Q[t])\cap\Sigma^{i}$, we have $\rho\preceq\psi_{i}^{-1}(x)$.
We will show that each operation can be solved in $O(\log\log\sigma)$ time. We now show that these operations are sufficient to compute $|G^{\prec}(p_{k}(\alpha))|$ and $|G_{\dashv}(p_{k}(\alpha))|$ in $O(\log\log\sigma)$ time.
  • Let us show how to compute $|G^{\prec}(p_{k}(\alpha))|$ in $O(\log\log\sigma)$ time. Let $j_{1}$ be the largest integer $0\leq j\leq n$ such that, for every $1\leq t\leq j$ and for every $\rho\in\lambda(Q[t])$, we have $\rho\prec p_{k}(\alpha)$. Let $j_{2}$ be the largest integer $0\leq j\leq n$ such that, for every $1\leq i\leq\min\{r,k-1\}$, we have $\mathtt{in}(Q[1,j],\alpha[k-i+1,k])\leq f_{i}$, where $f_{i}=\mathtt{out}(Q[1,|G^{\prec}(p_{k-i}(\alpha))|],\alpha[k-i+1,k])$. By Lemma 19 we have $|G^{\prec}(p_{k}(\alpha))|=\min\{j_{1},j_{2}\}$, so we only need to show how to compute $j_{1}$ and $j_{2}$ in $O(\log\log\sigma)$ time.
    • Let us show how to compute $j_{1}$ in $O(\log\log\sigma)$ time. For every $1\leq i\leq r$, let $j_{1,i}$ be the largest integer $0\leq j\leq n$ such that, for every $1\leq t\leq j$ and for every $\rho\in\lambda(Q[t])\cap\Sigma^{i}$, we have $\rho\prec p_{k}(\alpha)$. Then, we have $j_{1}=\min\{j_{1,1},j_{1,2},\dots,j_{1,r}\}$. Hence, we only have to show that, for every $1\leq i\leq r$, we can compute $j_{1,i}$ in $O(\log\log\sigma)$ time, because then we can compute $j_{1}$ in $O(r\log\log\sigma)\subseteq O(\log\log\sigma)$ time.
      We distinguish two cases: $\mathfrak{q}\geq k$ and $\mathfrak{q}<k$. First, assume that $\mathfrak{q}\geq k$. This means that $p_{k}(\alpha)=0^{k}$. Fix $1\leq i\leq r$. We distinguish two subcases.
        Assume that $1\leq i\leq\min\{r,k-1\}$. Then, $j_{1,i}$ is the largest integer $0\leq j\leq n$ such that, for every $1\leq t\leq j$ and for every $\rho\in\lambda(Q[t])\cap\Sigma^{i}$, we have $\rho\preceq 0^{i}$. Since $\psi_{i}(0^{i})=0$, we compute $j_{1,i}=\mathcal{A}.op_{4}(i,0)$ in $O(\log\log\sigma)$ time.
        Assume that $k\leq i\leq r$. Then, $j_{1,i}$ is the largest integer $0\leq j\leq n$ such that, for every $1\leq t\leq j$, there exists no $\rho\in\lambda(Q[t])\cap\Sigma^{i}$. By the definition of $\mathtt{IN}_{i}$, we compute $j_{1,i}=\mathtt{IN}_{i}.rank(\mathtt{IN}_{i}.select(1,0),1)$ in $O(\log\log\sigma)$ time (note that the formula for $j_{1,i}$ is correct even when $e_{i}=0$ because in this case $\mathtt{IN}=1^{n}$, $\mathtt{IN}_{i}.select(1,0)=n+1$ and $\mathtt{IN}_{i}.rank(\mathtt{IN}_{i}.select(1,0),1)=n$).
      Now, assume that $\mathfrak{q}<k$. This means that $p_{k}(\alpha)\not=0^{k}$, so $p_{k}(\alpha)$ contains at least one character distinct from 0. For every $1\leq i\leq r$, let $\beta_{i}$ the largest string (w.r.t $\preceq$) $\beta\in\Sigma^{i}$ such that $\beta\prec p_{k}(\alpha)$. Note that $\beta_{i}$ is well defined because (i) $\Sigma^{i}$ is a finite set and (ii) $p_{k}(\alpha)\not=0^{k}$ implies $0^{i}\prec p_{k}(\alpha)$.
      Fix $1\leq i\leq r$. Let us prove that, for every $\rho\in\Sigma^{i}$, we have $\rho\prec p_{k}(\alpha)$ if and only if $\rho\preceq\beta_{i}$. Indeed, (i) if $\rho\preceq\beta_{i}$, then $\rho\preceq\beta_{i}\prec p_{k}(\alpha)$, and (ii) if $\rho\prec p_{k}(\alpha)$, then $\rho\preceq\beta_{i}$ by the maximality of $\beta_{i}$. Consequently, $j_{1,i}$ is the largest integer $0\leq j\leq n$ such that, for every $1\leq t\leq j$ and for every $\rho\in\lambda(Q[t])\cap\Sigma^{i}$, we have $\rho\preceq\beta_{i}$. Hence, if we know $\psi_{i}(\beta_{i})$, then in $O(\log\log\sigma)$ time we can compute $j_{1,i}$ because $j_{1,i}=\mathcal{A}.op_{4}(i,\psi_{i}(\beta_{i}))$.
      We are only left with showing that in $O(1)$ time we can compute $\psi_{i}(\beta_{i})$ for every $1\leq i\leq r$. Let us first determine $\beta_{i}$ for every $1\leq i\leq r$. We consider two cases:
        Assume that $1\leq i\leq\min\{r,k-1\}$. Then, we have $\beta_{i}=\alpha[k-i+1,k]$ because (i) $\alpha[k-i+1,k]\prec p_{k}(\alpha)$ and (ii) if $\rho\in\Sigma^{i}$ satisfies $\rho\prec p_{k}(\alpha)$, then $\rho\preceq\alpha[k-i+1,k]$. For example, if $\sigma=10$ and $p_{k}(\alpha)=352$, then $\beta_{2}=52$.
        Assume that $k\leq i\leq r$. Then, we have $\beta_{i}=(\sigma-1)^{i-k+\mathfrak{q}}(\alpha[\mathfrak{q}+1]-1)\alpha[\mathfrak{q}+2,k]$, where $\alpha[\mathfrak{q}+1]-1$ is the character preceding $\alpha[\mathfrak{q}+1]$ (we have (i) $\mathfrak{q}<m$ because $\mathfrak{q}<k\leq m$ and (ii) $\alpha[\mathfrak{q}+1]\not=0$ by the definition of $\mathfrak{q}$) and $(\sigma-1)^{i-k+\mathfrak{q}}$ is the concatenation of $i-k+\mathfrak{q}$ occurrences of $\sigma-1$, which is the largest character in $\Sigma$. For example, if $\sigma=10$ and $p_{k}(\alpha)=000752$, then $\beta_{8}=99999652$.
      We conclude that in $O(1)$ time (i) we can compute $\psi_{i}(\beta_{i})$ for every $1\leq i\leq\min\{r,k-1\}$ by suffix-mapping the string $\alpha[k-\min\{r,k-1\}+1,k]$ and (ii) we can compute $\psi_{i}(\beta_{i})$ for every $k\leq i\leq r$ by suffix-mapping the string $(\sigma-1)^{r-k+\mathfrak{q}}(\alpha[\mathfrak{q}+1]-1)\alpha[\mathfrak{q}+2,k]$.
    • Let us show how to compute $j_{2}$ in $O(\log\log\sigma)$ time. We first prove that in $O(1)$ time we can compute $f_{i}$ for every $1\leq i\leq\min\{r,k-1\}$. We already know $|G^{\prec}(p_{i}(\alpha))|$ for every $1\leq i\leq\min\{r,k-1\}$. We suffix-map the string $\alpha[k-\min\{r,k-1\}+1,k]$ in $O(1)$ time, obtaining $\psi_{i}(\alpha[k-i+1,k])$ for every $1\leq i\leq\min\{r,k-1\}$. Hence in $O(r\log\log\sigma)\subseteq O(\log\log\sigma)$ time we compute $f_{i}=\mathtt{out}(Q[1,|G^{\prec}(p_{k-i}(\alpha))|],\alpha[k-i+1,k])=\mathcal{A}.op_{1}(i,\psi_{i}(\alpha[k-i+1,k]),|G^{\prec}(p_{k-i}(\alpha))|)$ for every $1\leq i\leq\min\{r,k-1\}$.
      For every $1\leq i\leq\min\{r,k-1\}$, let $j_{2,i}$ be the largest integer $0\leq j\leq n$ such that $\mathtt{in}(Q[1,j],\alpha[k-i+1,k])\leq f_{i}$. Then, we have $j_{2}=\min\{j_{2,1},j_{2,2},\dots,j_{2,\min\{r,k-1\}}\}$. Hence, we only have to show that, for every $1\leq i\leq\min\{r,k-1\}$, we can compute $j_{2,i}$ in $O(\log\log\sigma)$ time, because then we can compute $j_{2}$ in $O(r\log\log\sigma)\subseteq O(\log\log\sigma)$ time. To this end, we only need to observe that $j_{2,i}=\mathcal{A}.op_{2}(i,\psi_{i}(\alpha[k-i+1,k]),f_{i})$.
  • Let us show how to compute $|G^{\prec}_{\dashv}(p_{k}(\alpha))|$ in $O(\log\log\sigma)$ time. Let $j_{1}$ be the largest integer $0\leq j\leq n$ such that, if $j\geq 1$, then $Q[j]\in G^{*}(p_{k}(\alpha))$. Let $j_{2}$ be the smallest integer $0\leq j\leq n$ such that, for every $1\leq i\leq\min\{r,k-1\}$ for which $g_{i}>f_{i}$, we have $\mathtt{in}(Q[1,j],\alpha[k-i+1,k])\geq g_{i}$, where $f_{i}=\mathtt{out}(Q[1,|G^{\prec}(p_{k-i}(\alpha))|],\alpha[k-i+1,k])$ and $g_{i}=\mathtt{out}(Q[1,|G^{\prec}_{\dashv}(p_{k-i}(\alpha))|],\alpha[k-i+1,k])$. By Lemma 20 we have $|G^{\prec}_{\dashv}(p_{k}(\alpha))|=\max\{|G^{\prec}(p_{k}(\alpha))|,j_{1},j_{2}\}$, and we have already computed $|G^{\prec}(p_{k}(\alpha))|$, so we only need to show how to compute $j_{1}$ and $j_{2}$ in $O(\log\log\sigma)$ time.
    • Let us show how to compute $j_{1}$ in $O(\log\log\sigma)$ time. If $k>r$, we immediately conclude $j_{1}=0$ (see Remark 21), so we can assume $k\leq r$. For every $k\leq i\leq r$, let $j_{1,i}$ be the largest $0\leq j\leq n$ such that, if $j\geq 1$, then there exists $\rho\in Q[j]\cap\Sigma^{i}$ such that $p_{k}(\alpha)\dashv\rho$. Then, we have $j_{1}=\max\{j_{1,k},j_{1,k+1},\dots,j_{1,r}\}$. Hence, we only have to show that, for every $k\leq i\leq r$, we can compute $j_{1,i}$ in $O(\log\log\sigma)$ time, because then we can compute $j_{1}$ in $O(r\log\log\sigma)\subseteq O(\log\log\sigma)$ time.
      Fix $k\leq i\leq r$, and consider the strings $0^{i-k}p_{k}(\alpha)$ and $(\sigma-1)^{i-k}p_{k}(\alpha)$ (recall that 0 and $\sigma-1$ are the smallest and the largest character in $\Sigma$, respectively). In $O(1)$ time, we compute $\psi_{i}(0^{i-k}p_{k}(\alpha))$ and $\psi_{i}((\sigma-1)^{i-k}p_{k}(\alpha))$ for every $k\leq i\leq r$ by suffix-mapping the strings $0^{r-k}p_{k}(\alpha)$ and $(\sigma-1)^{r-k}p_{k}(\alpha)$. Then, compute $d_{1}=\Sigma_{i}^{*}.rank(\psi_{i}(0^{i-k}p_{k}(\alpha))-1)$ (if $p_{k}(\alpha)=0^{k}$, we assume $\Sigma_{i}^{*}.rank(\psi_{i}(0^{i-k}p_{k}(\alpha))-1)=0$) and $d_{2}=\Sigma_{i}^{*}.rank(\psi_{i}((\sigma-1)^{i-k}p_{k}(\alpha)))$ in $O(\log\log\sigma)$ time. Since $\psi_{i}$ is monotone, we have $d_{1}\leq d_{2}$. Moreover, we have $d_{1}=d_{2}$ if and only if $j_{1,i}=0$, so in the rest of the proof we can assume $d_{1}<d_{2}$ (and so $j_{1,i}\geq 1$).
      Since $j_{1,i}\geq 1$ and $\Sigma^{i}$ is finite, we can consider the largest string $\rho^{*}_{i}$ (w.r.t. $\preceq$) in $\Sigma_{i}$ suffixed by $p_{k}(\alpha)$. By Lemma 13, $j_{1,i}$ is the largest $0\leq j\leq n$ such that $\rho^{*}_{i}\in\lambda(Q[j])$. Since $\psi_{i}$ is monotone, then number $f$ of strings in $\Sigma_{i}$ smaller than or equal to $\rho^{*}_{i}$ can be computed in $O(\log\log\sigma)$ time because $f=\Sigma_{i}^{*}.rank(d_{2})$. Now consider the list of all edges of $\mathcal{A}$ sorted in the order used to define $\mathtt{AUX}_{i}$. Then, the largest edge labeled $\rho^{*}_{i}$ in the list is the $g$-th smallest edge of the list, where $g=\mathtt{AUX}.rank(f+1,1)-1$, and we can compute $g$ in $O(1)$ time. By the definition of $\mathtt{AUX}_{i}$, this edge reaches state $Q[j_{1,i}]$. By the definitions of $\mathtt{AUX}_{i}$ and $\mathtt{IN}_{i}$, we have $j_{1,i}=\mathtt{IN}_{i}.rank(\mathtt{IN}_{i}.select(g,0),1)+1$, and we can compute $j_{1,i}$ in $O(1)$ time.
    • Let us show how to compute $j_{2}$ in $O(\log\log\sigma)$ time. We first prove that in $O(1)$ time we can compute $f_{i}$ and $g_{i}$ for every $1\leq i\leq\min\{r,k-1\}$. We already know $|G^{\prec}(p_{i}(\alpha))|$ and $|G^{\prec}_{\dashv}(p_{i}(\alpha))|$ for every $1\leq i\leq\min\{r,k-1\}$. We suffix-map the string $\alpha[k-\min\{r,k-1\}+1,k]$ in $O(1)$ time, obtaining $\psi_{i}(\alpha[k-i+1,k])$ for every $1\leq i\leq\min\{r,k-1\}$. Hence in $O(r\log\log\sigma)\subseteq O(\log\log\sigma)$ time we compute $f_{i}=\mathtt{out}(Q[1,|G^{\prec}(p_{k-i}(\alpha))|],\alpha[k-i+1,k])=\mathcal{A}.op_{1}(i,\psi_{i}(\alpha[k-i+1,k]),|G^{\prec}(p_{k-i}(\alpha))|)$ and $g_{i}=\mathtt{out}(Q[1,|G^{\prec}_{\dashv}(p_{k-i}(\alpha))|],\alpha[k-i+1,k])=\mathcal{A}.op_{1}(i,\psi_{i}(\alpha[k-i+1,k]),|G^{\prec}_{\dashv}(p_{k-i}(\alpha))|)$ for every $1\leq i\leq\min\{r,k-1\}$.
      For every $1\leq i\leq\min\{r,k-1\}$, let $j_{2,i}$ be the smallest $0\leq j\leq n$ such that $\mathtt{in}(Q[1,j],\alpha[k-i+1,k])\geq g_{i}$. Then, we have $j_{2}=\max\{j_{2,i}\;|\;1\leq i\leq\min\{r,k-1\},g_{i}>f_{i}\}$. Hence, we only have to show that, for every $1\leq i\leq\min\{r,k-1\}$ for which $g_{1}>f_{i}$ (and in particular $g_{i}>0$), we can compute $j_{2,i}$ in $O(\log\log\sigma)$ time, because then we can compute $j_{2}$ in $O(r\log\log\sigma)\subseteq O(\log\log\sigma)$ time. To this end, we only need to observe that $j_{2,i}=\mathcal{A}.op_{3}(i,\psi_{i}(\alpha[k-i+1,k]),g_{i})$.
To conclude the description of our algorithm solving the SMLG problem, we only need to show that the four operations defined above can be solved in $O(\log\log\sigma)$ time. Here are the details.
  • $\mathcal{A}.op_{1}(i,x,j)$, for $1\leq i\leq r$, $0\leq x\leq\sigma^{i}-1$ and $1\leq j\leq n$: return $\mathtt{out}(Q[1,j],\psi_{i}^{-1}(x))$. We have $\psi_{i}^{-1}(x)\in\Sigma^{i}$. We fist check whether $\psi_{i}^{-1}(x)\in\Sigma_{i}$. We have $\psi_{i}^{-1}(x)\in\Sigma_{i}$ if and only if $x\in\Sigma_{i}^{*}$, so we only need to solve the query $\Sigma_{i}^{*}.memb(x)$ in $O(\log\log\sigma)$ time. If $\psi_{i}^{-1}(x)\not\in\Sigma_{i}$, then $\mathtt{out}(Q[1,j],\psi_{i}^{-1}(x))=0$. Now assume that $\psi_{i}^{-1}(x)\in\Sigma_{i}$. All occurrences of $\psi_{i}^{-1}(x)$ in $\mathtt{LAB}_{i}$ have been replaced with $x^{\prime}=\Sigma_{i}^{*}.rank(\psi_{i}(\psi_{i}^{-1}(x)))-1=\Sigma_{i}^{*}.rank(x)-1$ in $\mathtt{LAB}_{i}^{*}$, and we can compute $x^{\prime}$ in $O(\log\log\sigma)$ time. By the definition of $\mathtt{OUT}_{i}$, the number of edges in $E_{i}$ leaving a state in $Q[1,k]$ is given by $d=\mathtt{OUT}_{i}.rank(\mathtt{OUT}_{i}.select(k,1),0)$, which can be computed in $O(1)$ time. As a consequence, in $O(\log\log\sigma)$ time we can compute $\mathtt{out}(Q[1,j],\psi_{i}^{-1}(x))$ because by the definitions of $\mathtt{LAB}$ and $\mathtt{LAB}^{*}$ we have $\mathtt{out}(Q[1,j],\psi_{i}^{-1}(x))=\mathtt{LAB}_{i}^{*}.rank(d,x^{\prime})$.
  • $\mathcal{A}.op_{2}(i,x,h)$, for $1\leq i\leq r$, $0\leq x\leq\sigma^{i}-1$ and $h\geq 0$: return the largest $0\leq j\leq n$ such that $\mathtt{in}(Q[1,j],\psi_{i}^{-1}(x))\leq h$. We have $\psi_{i}^{-1}(x)\in\Sigma^{i}$. In $O(\log\log\sigma)$ time, we check whether $\psi_{i}^{-1}(x)\in\Sigma_{i}$ by proceedings as in the previous point. If $\psi_{i}^{-1}(x)\not\in\Sigma_{i}$, then we conclude $j=n$. Now assume that $\psi_{i}^{-1}(x)\in\Sigma_{i}$. All occurrences of $\psi_{i}^{-1}(x)$ in $\mathtt{LAB}_{i}$ have been replaced with $x^{\prime}=\Sigma_{i}^{*}.rank(\psi_{i}(\psi_{i}^{-1}(x)))-1=\Sigma_{i}^{*}.rank(x)-1$ in $\mathtt{LAB}_{i}^{*}$, and we can compute $x^{\prime}$ in $O(\log\log\sigma)$ time. Since $\psi_{i}$ is a bijection, the total number of edges in $\mathcal{A}$ labeled $\psi_{i}^{-1}(x)$ is $d=\mathtt{LAB}^{*}_{i}.rank(e_{i},x^{\prime})$. If $d\leq h$, we conclude $j=n$. Now assume that $d>h$. Since $\psi_{i}$ is monotone, the number $f$ of strings in $\Sigma_{i}$ smaller than or equal to $\psi_{i}^{-1}(x)$ can be computed in $O(\log\log\sigma)$ time because $f=\Sigma_{i}^{*}.rank(d)$. Now consider the list of all edges of $\mathcal{A}$ sorted in the order used to define $\mathtt{AUX}_{i}$. Then, the $h+1$-th smallest edge labeled $\psi_{i}^{-1}(x)$ in the list is the $g$-th smallest edge of the list, where $g=\mathtt{AUX}.rank(f,1)+h$, and we can compute $g$ in $O(1)$ time. By the definitions of $\mathtt{AUX}_{i}$ and $\mathtt{IN}_{i}$, this edge reaches state $Q[y]$, where $y=\mathtt{IN}_{i}.rank(\mathtt{IN}_{i}.select(g,0),1)+1$, and we can compute $y$ in $O(1)$ time. Hence, we conclude that the largest $0\leq j\leq n$ such that $\mathtt{in}(Q[1,j],\psi_{i}^{-1}(x))\leq h$ is y1y-1.
  • $\mathcal{A}.op_{3}(i,x,h)$, for $1\leq i\leq r$, $0\leq x\leq\sigma^{i}-1$ and $h\geq 1$: return the smallest $0\leq j\leq n$ such that $\mathtt{in}(Q[1,j],\psi_{i}^{-1}(x))\geq h$, or report that such a $j$ does not exist. Let $j^{\prime}$ be the largest integer for which $0\leq j^{\prime}\leq n$ and $\mathtt{in}(Q[1,j^{\prime}],\psi_{i}^{-1}(x))\leq h-1$. We can compute $j^{\prime}$ in $O(\log\log\sigma)$ time by using $\mathcal{A}.op_{2}(i,x,h-1)$. If $j^{\prime}<n$, then the smallest $0\leq j\leq n$ such that $\mathtt{in}(Q[1,j],\psi_{i}^{-1}(x))\geq h$ is $j^{\prime}+1$ by the maximality of $j^{\prime}$, and if $j^{\prime}=n$, then such a $j$ does not exist.
  • $\mathcal{A}.op_{4}(i,x)$, for $1\leq i\leq r$ and $0\leq x\leq\sigma^{i}-1$: return the largest integer $0\leq j\leq n$ such that, for every $1\leq t\leq j$ and for every $\rho\in\lambda(Q[t])\cap\Sigma^{i}$, we have $\rho\preceq\psi_{i}^{-1}(x)$. We compute $\Sigma_{i}^{*}.succ(x)$ in $O(\log\log\sigma)$ time. If $\Sigma_{i}^{*}.succ(x)=\bot$, then for every $y\in\Sigma^{*}_{i}$ we have $y\leq x$. Hence, for every $\rho\in\Sigma_{i}$ we have $\rho\preceq\psi_{i}^{-1}(x)$ (because $\psi_{i}$ is monotone and $\psi_{i}(\rho)\leq x$). This implies that $j=n$. Now assume that $\Sigma_{i}^{*}.succ(x)\not$$\bot$, which implies that $\psi_{i}^{-1}(\Sigma_{i}^{*}.succ(x))\in\Sigma_{i}$. Note that for every $\rho\in\Sigma_{i}$ we have $\rho\prec\psi_{i}^{-1}(\Sigma_{i}^{*}.succ(x))$ if and only if $\psi_{i}(\rho)<\Sigma_{i}^{*}.succ(x)$, if and only if $\psi_{i}(\rho)\leq x$, if and only if $\rho\preceq\psi_{i}^{-1}(x)$. This means that we only have to compute the largest integer $0\leq j\leq n$ such that, for every $1\leq t\leq j$ and for every $\rho\in\lambda(Q[t])\cap\Sigma^{i}$, we have $\rho\prec\psi_{i}^{-1}(\Sigma_{i}^{*}.succ(x))$. Since $\psi_{i}^{-1}(\Sigma_{i}^{*}.succ(x))\in\Sigma_{i}$, by Lemma 13 we conclude that that we only have to compute the largest integer $0\leq j\leq n$ such that $\mathtt{in}(Q[1,j],\psi_{i}^{-1}(\Sigma_{i}^{*}.succ(x)))\leq 0$, so we can compute $\mathcal{A}.op_{2}(i,\Sigma_{i}^{*}.succ(x),0)$ in $O(\log\log\sigma)$ time.
To conclude the proof of the theorem, we need to show that in $O(m\log\log\sigma)$ time we can also decide whether $\alpha$ is recognized by $\mathcal{A}$. Let $\mathcal{A}^{\prime}$ be the GDFA obtained from $\mathcal{A}$ as follows: (i) we add a new state $s^{\prime}$, (ii) we add the edge $(s^{\prime},s,\#)$, where $\not\in\Sigma$ is a new character smaller than all characters in $\Sigma$ (w.r.t. $\preceq$) and $s$ is the initial state of $\mathcal{A}$, (iii) we let $s$ (and not $s^{\prime}$) be the initial state of $\mathcal{A}$. Note that $s^{\prime}$ has no incoming edges, so if we navigate $\mathcal{A}^{\prime}$ starting from $s^{\prime}$, after leaving $s^{\prime}$ we only move within $\mathcal{A}$. For every $\alpha\in\Sigma^{*}$ and for every state $u$ of $\mathcal{A}$, we have $\alpha\in I^{\mathcal{A}}_{u}$ if and only if $\alpha\in I^{\mathcal{A}^{\prime}}_{u}$. Consequently, $\mathcal{A}^{\prime}$ is also a Wheeler GDFA (because for every $\alpha,\beta\in\Sigma^{*}$ we have $\alpha\prec\beta$ if and only if $\alpha\prec\#\beta$), and $\preceq_{\mathcal{A}^{\prime}}$ is obtained from $\preceq_{\mathcal{A}}$ by letting $s^{\prime}$ be the smallest state, without changing the mutual order of the remaining states. Checking whether $\alpha$ is recognized by $\mathcal{A}$ is equivalent to checking whether $\alpha$ is recognized by $\mathcal{A}^{\prime}$. Since $\mathcal{A}^{\prime}$ contains exactly one edge labeled #\#, we can proceed as follows. We first solve the SMLG problem on $\mathcal{A}^{\prime}$ (with input $\alpha$). The SMLG problem returns at most one state. If the SMLG problem returns no state, we conclude that $\alpha$ is not recognized by $\mathcal{A}$. If the SMLG problem returns exactly one state $j$, then $\alpha$ is recognized by $\mathcal{A}$ if and only if $j$ is a final state of $\mathcal{A}$, so we only need to solve $\mathtt{FIN}.access(j)$ in $O(1)$ time. To solve the SMLG problem on $\mathcal{A}^{\prime}$ (with input $\alpha$), we first process the character #\#. Note that $G^{\prec}(\#)=\{s^{\prime}\}$ and $G^{\prec}_{\dashv}(\#)=\{s^{\prime},s\}$. After processing #\#, to process $\alpha$ we only move within $\mathcal{A}$. This means that, to solve the SMLG problem on $\mathcal{A}^{\prime}$ (with input $\alpha$), we can simply solve the SMLG problem of $\mathcal{A}$ (with input $\alpha$), as long as we artificially replace $|G^{\prec}(\epsilon)|=0$ and $|G^{\prec}_{\dashv}(\epsilon)|=n$ with $|G^{\prec}(\epsilon)|=0$ and $|G^{\prec}_{\dashv}(\epsilon)|=1$. The time bound $O(m\log\log\sigma)$ follows from the bound for the SMLG problem.
Remark 27.
The proof of Theorem 26 shows that our algorithm for solving the SMLG problem and deciding whether a string is recognized by a Wheeler GDFA is online: we iteratively solve the same problems for every prefix of the pattern $\alpha$.
5. Conclusions and Future Work
In this paper, we considered the model of generalized automata, and we introduced the set $\mathcal{W(A)}$. We showed that $\mathcal{W(A)}$ plays the same role played by $\operatorname{Pref}(\mathcal{L(A)})$ in conventional NFAs: the set $\mathcal{W(A)}$ can be used to derive a Myhill-Nerode theorem, and it represents the starting point for extending the FM-index to generalized automata.
Further lines of research include extending the Burrows-Wheeler Transform and the FM-index to arbitrary GNFAs. Indeed, the Burrows-Wheeler Transform and the FM-index were recently generalized from Wheeler NFAs to arbitrary NFAs through the so-called co-lex orders [CDPP23, CP21] and co-lex relations [Cot22]. However, we remark that the efficient time bounds for the SMLG problem that we derived in this paper cannot hold for arbitrary GNFAs due to the (conditional) lower bounds by Equi et al. that we recalled in the introduction.
Giammaresi and Montalbano described an effective procedure for computing a state-minimal GDFA equivalent to a given GDFA [GM99, GM95], but we do not know if there exists an efficient algorithm for minimizing a GDFA. On the one hand, the Myhill-Nerode theorem for generalized automata implies that for every state-minimal GDFA $\mathcal{A}$ there exists a set $\mathcal{W}\subseteq\Sigma^{*}$ such that $\mathcal{A}$ is isomorphic to the minimal $\mathcal{W}$-GDFA recognizing $\mathcal{L(A)}$. On the other hand, given a $\mathcal{W}$-GDFA recognizing $\mathcal{L}$, it should be possible to build the minimal $\mathcal{W}$-GDFA recognizing $\mathcal{L}$ by extending Hopcroft’s algorithm [Hop71] to GDFAs. If we could prove that, for every admissible $\mathcal{W}\subseteq\Sigma^{*}$, the number of states of the minimal $\mathcal{W}$-GDFA recognizing $\mathcal{L}$ is comparable to the number of states of a minimal GDFA recognizing $\mathcal{L}$, then we would obtain a fast algorithm that significantly reduces the number of states of a GDFA without changing the recognized language.
As discussed in Section 4, we do not know whether in Lemma 14 we can achieve $O(\mathfrak{e})$ time in the case of an integer alphabet in a polynomial range. Moreover, the paper leaves many questions of theoretical interest open. The class of Wheeler languages is the class of all regular languages that are recognized by some Wheeler NFA [ADPP21]. Wheeler languages enjoy several properties: for example, they admit a characterization in terms of convex equivalence relations [ADPP21]. In addition, every Wheeler language is also recognized by some DFA, and, in particular, there exists a unique state-minimal DFA recognizing a given Wheeler language [ACP22]. The main limitation of Wheeler languages is that they capture only a small subclass of regular languages: for example, a unary language (that is, a language over an alphabet of size one) is Wheeler if and only if it is either finite or co-finite [ADPP21]. The intuitive reason why most regular languages are not Wheeler is that, if $\mathcal{A}$ is a Wheeler NFA, then Wheelerness induces strong constraints on the set $\operatorname{Pref}(\mathcal{L(A)})$. However, when we switch to GNFAs, the role of $\operatorname{Pref}(\mathcal{L(A)})$ is played by $\mathcal{W(A)}$, and it may hold $\mathcal{W(A)}\subsetneqq\operatorname{Pref}(\mathcal{L(A)})$, which means that now the same constraints only apply to a smaller subset. The natural question is whether Wheeler GNFAs extend the class of Wheeler languages. The answer is affirmative: there exists a regular language $\mathcal{L}$ such that $\mathcal{L}$ is not Wheeler (that is, no Wheeler NFA recognizes $\mathcal{L}$), but $\mathcal{L}$ is recognized by a Wheeler GDFA. Define $\mathcal{L}=\{a^{2n}\;|\;n\geq 0\}$. Then, $\mathcal{L}$ is not Wheeler [ADPP21], but $\mathcal{L}$ is recognized by the GDFA consisting of a single state, both initial and final, with a self-loop labeled $aa$. As a consequence, the class of all languages recognized by some Wheeler GNFA is strictly larger than the class of Wheeler languages. We can call the languages in this new class generalized Wheeler languages: the next step is to understand which properties of Wheeler languages are still true and how it is possible to characterize this new class.
Appendix A The FM-index: from GDFAs to GNFAs without $\epsilon$-transitions
This appendix aims to show that the results of Section 4 can be extended to generalized nondeterministic automata. We will consider the case of GNFAs without $\epsilon$-transitions (that is, GNFAs where no edge is labeled with $\epsilon$), and we will focus on the differences between the case of GDFAs and the case of GNFAs without $\epsilon$-transitions. Extending our results to GNFA with $\epsilon$-transitions (within the same space and time bounds) requires additional technical machinery that goes beyond the scope of this paper. The case of GNFAs with $\epsilon$-transitions is discussed in detail in a separate article [Cot25].
The first step is to extend the notion of Wheelerness to GNFAs without $\epsilon$-transitions. Let us recall the definition of Wheeler NFA (see [ADPP20, CCG+23, CGKP23]). An NFA $\mathcal{A}=(Q,E,s,F)$ is Wheeler if there exists a total order $\leq$ on $Q$ such that (i) $s$ comes first in the total order, (ii) for every $(u^{\prime},u,a),(v^{\prime},v,b)\in E$, if $u<v$, then $a\preceq b$ and (iii) for every $(u^{\prime},u,a),(v^{\prime},v,a)\in E$, if $u<v$, then $u^{\prime}\leq v^{\prime}$. Alanko et al. [ADPP20, Lemma 2.3] proved that, if $u<v$ in the total order, then $(\forall\alpha\in I_{u})(\forall\beta\in I_{v})((\{\alpha,\beta\}\not\subseteq I_{u}\cap I_{v})\Longrightarrow(\alpha\prec\beta))$. Let us see how to extend Definition 4.2 to GNFAs without $\epsilon$-transitions. We will first generalize the definition of $\preceq_{\mathcal{A}}$ from GDFAs to GNFAs without $\epsilon$-transitions drawing inspiration from Alanko et al.’s result.
Let $\mathcal{A}=(Q,E,s,F)$ be a GNFA without $\epsilon$-transitions. Let $\preceq_{\mathcal{A}}$ be the relation on $Q$ such that, for every $u,v\in Q$, we have $u\preceq_{\mathcal{A}}v$ if and only if $(\forall\alpha\in I_{u})(\forall\beta\in I_{v})((\{\alpha,\beta\}\not\subseteq I_{u}\cap I_{v})\Longrightarrow(\alpha\prec\beta))$. If $\mathcal{A}$ is a GDFA, then $\preceq_{\mathcal{A}}$ reduces to the definition given in Section 4, because for every $u,v\in Q$ such that $u\not=v$ we have $I_{u}\cap I_{v}=\emptyset$ (see Remark 4). We have seen that, if $\mathcal{A}$ is a GDFA, then $\preceq_{\mathcal{A}}$ is a partial order. If $\mathcal{A}$ is an arbitrary GNFA without $\epsilon$-transitions, in general $\preceq_{\mathcal{A}}$ is only a preorder, that is, it is a reflexive and transitive relation, but it need not be antisymmetric, see Figure 10 (this is already true for NFAs).
We can now give the following definition.
{defi}
Let $\mathcal{A}=(Q,E,s,F)$ be a GNFA without $\epsilon$-transitions. We say that $\mathcal{A}$ is Wheeler if there exists a total order $\leq$ on $Q$ such that:
  • (Property 1) For every $u,v\in Q$, if $u\leq v$, then $u\preceq_{\mathcal{A}}v$.
  • (Property 2) For every $(u^{\prime},u,\rho),(v^{\prime},v,\rho^{\prime})\in E$, if $u<v$ and $\rho^{\prime}$ is not a strict suffix of $\rho$, then $\rho\preceq\rho^{\prime}$.
  • (Property 3) For every $(u^{\prime},u,\rho),(v^{\prime},v,\rho)\in E$, if $u<v$, then $u^{\prime}\leq v^{\prime}$.
We say that $\leq$ is a Wheeler order on $\mathcal{A}$.
Note that if $\leq$ is a Wheeler order, then $s$ comes first by Property 1, because $\epsilon\in I_{s}$ and for every $u\in Q\setminus\{s\}$ we have $\epsilon\not\in I_{u}$. As a consequence, Lemma 13 is true also for GNFAs without $\epsilon$-transitions (if in the statement of Lemma 13 we replace $\preceq_{\mathcal{A}}$ with a Wheeler order $\leq$).
If $\mathcal{A}$ is an NFA, then Definition A reduces to the definition of Wheeler NFA, because by Alanko et al.’s result Property 1 follows from Property 2, Property 3, and the fact that $s$ comes first in a Wheeler order. If $\mathcal{A}$ is a GDFA, then Definition A reduces to the definition of Wheeler GDFA (Definition 4.2) by Lemma 13.
Let us present some preliminary remarks (see Figure 10). In general, a Wheeler order on a GNFA without $\epsilon$-transitions is not unique (this is already true for NFAs). Moreover, Property 2 and Property 3 in Definition A do not follow from Property 1 (this is already true for NFAs). Lastly, Property 1 does not follow from Properties 2, 3 and the requirement that $s$ must come first (while we have just seen that, if $\mathcal{A}$ is an NFA, then Property 1 follows from Properties 2, 3 and the requirement that $s$ must come first).
The problem of deciding whether a GDFA is Wheeler can be solved in polynomial time (see Lemma 14), but it becomes NP-hard on GNFAs without $\epsilon$-transitions, because it is already NP-hard on NFAs [GT22]. There are natural ways of defining a Wheeler order on a GNFA without $\epsilon$-transitions. For example, it is easy to check that a Wheeler order on a GNFA without $\epsilon$-transitions is induced by any Wheeler order on the equivalent NFA (without $\epsilon$-transitions) defined in Section 2.
[…]
Figure 10. Left: An NFA such that $\preceq_{\mathcal{A}}$ is not antisymmetric and both total orders in which $u_{1}$ comes first are Wheeler (in particular, a Wheeler order need not be unique). Center: The total order $\leq$ given by $u_{1}<u_{2}<u_{3}<u_{4}<u_{5}$ is such that $\leq$ satisfies Property 1 of Definition A, but it does not satisfy Property 2 and Property 3. Right: The total order $\leq$ given by $u_{1}<u_{2}<u_{3}<u_{4}$ is such that $\leq$ satisfies Property 2 and Property 3 of Definition A and the initial state comes first, but it does not satisfy Property 1.
We will now outline how our results can be generalized from Wheeler GDFAs to Wheeler GNFAs without $\epsilon$-transitions (see [Cot25] for more details). Lemma 17 is still true if we replace $\preceq_{\mathcal{A}}$ with any Wheeler order $\leq$ on the GNFA without $\epsilon$-transitions. This follows from how $\preceq_{\mathcal{A}}$ is defined on GNFAs without $\epsilon$-transitions. For example, let us prove Property 3 in Lemma 17. We know that $u,v\in Q$ are such that $u<v$ and $v\in G^{\prec}(\alpha)$, and we must prove that $u\in G^{\prec}(\alpha)$. Since $u<v$ and $\leq$ is a Wheeler order, then $u\preceq_{\mathcal{A}}v$. Let $\beta\in I_{u}$; we must prove that $\beta\prec\alpha$. If $\beta\in I_{v}$, then from $v\in G^{\prec}(\alpha)$ we immediately conclude $\beta\prec\alpha$. Now assume that $\beta\not\in I_{v}$, and pick any $\gamma\in I_{v}$. From $u\preceq_{\mathcal{A}}v$, $\beta\in I_{u}\setminus I_{v}$ and $\gamma\in I_{v}$ we obtain $\beta\prec\gamma$. From $\gamma\in I_{v}$ and $v\in G^{\prec}(\alpha)$ we obtain $\gamma\prec\alpha$, so we conclude $\beta\prec\alpha$ and we are done. Next, one can readily check that the proofs of all remaining results in Section 4 still hold true (if everywhere we replace $\preceq_{\mathcal{A}}$ with a Wheeler order $\leq$) because, as we have seen, Lemma 13 is also true for GNFAs without $\epsilon$-transitions. A minor tweak is needed in Theorem 26 to decide whether a string $\alpha$ is recognized by a GNFA without $\epsilon$-transitions. In the case of GDFAs, the algorithm returns at most one state $j$, and we only need to check whether $j$ is final by solving $\mathtt{FIN}.access(j)$ in $O(1)$ time. In the case of GNFAs without $\epsilon$-transitions, the algorithm returns an interval $Q[d_{1},d_{2}]$, and we need to determine whether at least one state is final. To this end, we only need to check in $O(1)$ time whether $\mathtt{FIN}.rank(d_{2},1)>\mathtt{FIN}.rank(d_{1}-1,1)$.